### v2

# Whitehead’s problem and condensed mathematics

Abstract.
One of the better-known independence results in general mathematics is Shelah’s solution to Whitehead’s problem of whether $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$ implies that an abelian group $A$ is free. The point of departure for the present work is Clausen and Scholze’s proof that, in contrast, one natural interpretation of Whitehead’s problem within their recently-developed framework of condensed mathematics has an affirmative answer in $\mathsf{ZFC}$. We record two alternative proofs of this result, as well as several original variations on it, both for their intrinsic interest and as a springboard for a broader study of the relations between condensed mathematics and set theoretic forcing. We show more particularly how the condensation $\underline{X}$ of any locally compact Hausdorff space $X$ may be viewed as an organized presentation of the forcing names for the points of canonical interpretations of $X$ in all possible set-forcing extensions of the universe, and we argue our main result by way of this fact. We also show that when interpreted within the category of light condensed abelian groups, Whitehead’s problem is again independent of the $\mathsf{ZFC}$ axioms. In fact we show that it is consistent that Whitehead’s problem has a negative solution within the category of $\kappa$-condensed abelian groups for every uncountable cardinal $\kappa$, but that this scenario, in turn, is inconsistent with the existence of a strongly compact cardinal.
Key words and phrases:
Whitehead’s problem, condensed mathematics, forcing
2020 Mathematics Subject Classification:
03E35, 03E40, 03E05, 13C10, 20J05
The first author was supported by a María Zambrano Fellowship at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S. A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023. We thank the Fields Institute for their support and hospitality. The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023. We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks.
1. Introduction
Recall that an abelian group $A$ is Whitehead if $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$. Around 1950, J.H.C. Whitehead asked whether every such group is free. The converse is easily seen to be true (every free abelian group $A$ satisfies $\mathrm{Ext}^{1}(A,G)=0$ for every abelian group $G$, and hence for $G=\mathbb{Z}$), and Stein showed soon after the question was posed that every countable Whitehead group is free [25]. Progress on the general question was slow, though, and the latter soon assumed the status of a major open problem in homological algebra. Shelah’s solution to the question 20 years later was completely unexpected, for Whitehead’s problem thereby became one of the first originating outside of logic or set theory to be proven independent of the $\sf ZFC$ axioms.
Theorem 1.1 (Shelah, [21]).
Whitehead’s problem is independent of $\sf ZFC$. In particular:
  1. (1)
    If $V=L$, then every Whitehead group is free.
  2. (2)
    If $\sf MA_{\omega_{1}}$ holds, then there is a nonfree Whitehead group of size $\aleph_{1}$.
Shelah’s solution of Whitehead’s problem initiated a program of research applying set theoretic methods to module theory that continues to bear fruit today. This program has uncovered deep connections between Whitehead’s problem and its variations and purely combinatorial set theoretic phenomena, and has touched on additional topics including the study of almost free modules and of the approximation theory of modules. For more on such developments in the intervening years, see, e.g., [6, 8, 12].
Recently, Clausen and Scholze have developed a framework, known as condensed mathematics, for applying algebraic tools to algebraic structures endowed with nontrivial topologies (cf. [20], [19], [3]). Here, we will be particularly interested in the category $\mathsf{CondAb}$ of condensed abelian groups, where the objects are, roughly speaking, certain contravariant functors from the category of Stone spaces to the category of abelian groups. Every T1 topological abelian group $A$ has a corresponding condensed abelian group $\underline{A}$ (if an abelian group is given without specifying its topology, then we interpret it as a topological abelian group with the discrete topology). In addition, $\mathsf{CondAb}$ has an internal $\mathrm{Hom}$ functor $\underline{\mathrm{Hom}}:\mathsf{CondAb}^{\mathrm{op}}\times\mathsf{CondAb}\rightarrow\mathsf{CondAb}$, and corresponding internal $\mathrm{Ext}$ functors $\underline{\mathrm{Ext}}^{n}$ for $n>0$. It turns out that, when interpreted appropriately in $\mathsf{CondAb}$, Whitehead’s problem is no longer independent of $\sf ZFC$. In particular, Clausen and Scholze proved the following theorem (cf. [2]).
Theorem 1.2 (Clausen–Scholze, 2020).
It follows from the $\sf ZFC$ axioms that, for every abelian group $A$, if
$\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0,$
then $A$ is free.
Clausen and Scholze’s proof of Theorem 1.2 is quite slick and relies on a structural analysis of the abelian subcategory of $\mathsf{CondAb}$ spanned by what are known as solid abelian groups. In this paper, we present a more elementary, combinatorial proof of Theorem 1.2 that is also more constructive in the sense that, given a nonfree abelian group $A$, we specify a highly natural associated Stone space $S$ and construct a nonzero element of $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)$. Our motivation for doing this is twofold. First, we seek to excavate some of the combinatorial and set theoretic ideas underlying Theorem 1.2 and thereby to improve our understanding of the mathematics surrounding Whitehead’s problem, in both general and condensed contexts. To this end, we prove the following theorem.
Theorem 1.3.
Suppose that $A$ is a nonfree abelian group and $\kappa$ is the least cardinality of a nonfree subgroup of $A$. Let $\mathbb{B}$ be the Boolean completion of the forcing to add $\kappa$-many Cohen reals, and let $S$ be the Stone space of $\mathbb{B}$. Then:
  1. (1)
    $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(2^{\kappa})\neq 0$;
  2. (2)
    in $V^{\mathbb{B}}$, $A$ is not Whitehead;
  3. (3)
    $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$.
Second, using Theorem 1.3, particularly clauses (2) and (3) thereof, as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing. We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions. We then exhibit a correspondence between satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$. This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (3) of Theorem 1.3 from clause (2) of the same theorem.
It is natural, in light of the preceding paragraph, to wonder whether there are provable implications between instances of clauses (2) and (3) of Theorem 1.3 for more general complete Boolean algebras $\mathbb{B}$. We record several partial answers to this question. In particular, in Theorem 9.2 we show that if $\mathbb{B}$ is a complete Boolean algebra, $S$ is its Stone space, and $V^{\mathbb{B}}$ satisfies a certain technical strengthening of the failure of $A$ to be Whitehead, then $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$. In the ensuing discussion, we show that the converse fails in general. The question of whether clause (2) more generally implies clause (3) (i.e., of whether the aforementioned technical strengthening can be dropped) remains open and is recorded as Conjecture 13.1.
We also investigate some variations of Theorem 1.2. First, we show that, if we restrict our attention to the subcategory $\mathsf{CondAb}_{\kappa}$ of $\mathsf{CondAb}$ for certain uncountable cardinals $\kappa$, then the classical independence of Whitehead’s problem remains, while it again vanishes when $\kappa$ is a sufficiently large cardinal. More precisely, we prove the following theorem. (In what follows, given an uncountable cardinal $\kappa$, $\mathsf{CondAb}_{\kappa}$ denotes the category of $\kappa$-condensed sets, with the particular case of $\mathsf{CondAb}_{\omega_{1}}$ denoting the category of light condensed abelian groups, the ambient framework for the lecture series “Analytic Stacks” delivered by Clausen and Scholze in 2023–24 [4]).
Theorem 1.4.
  1. (1)
    If $\sf MA_{\omega_{1}}$ holds, then there is a nonfree abelian group of size $\aleph_{1}$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{1}}}(\underline{A},\underline{\mathbb{Z}})=0$.
  2. (2)
    It is consistent with the axioms of $\mathsf{ZFC}$ that for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0$.
  3. (3)
    If $\kappa$ is a strongly compact cardinal, $A$ is an abelian group, and
    $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0,$
    then $A$ is free.
Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups.
The structure of the paper is as follows. In Section 2, we introduce the basic definitions and facts regarding condensed mathematics that we will need. In Section 3, we recall some classical algebraic facts pertinent to Whitehead’s problem and introduce their analogues in the condensed setting. In Section 4, we prove a slight strengthening of Stein’s aforementioned classical result [25] that all countable Whitehead groups are free. In Section 5, we prove clause (1) of Theorem 1.3, and, in Section 6, we adapt that proof to establish clause (2) of Theorem 1.3. From here we present two paths to clause (3) of Theorem 1.3. The first, recorded in Section 7, is by far the shorter and more direct of the two. The second, comprising Sections 8 and 9, passes through a more general investigation into the interplay between condensed mathematics and forcing. In Section 8, a robust correspondence is established between
  • $\mathbb{B}$-names for points in a certain extension of a topological space $X$; and
  • continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$.
Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas. We expect the results of this section to be applicable in a wide variety of settings, and derive clause (3) of Theorem 1.3 as an immediate corollary. In Section 10, we prove Theorem 1.4(1). In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form. This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4. In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups. We close out our work in Section 13 with a few concluding remarks and main open questions.
We have pursued wherever possible an account of our results accessible to readers with no more than a basic knowledge of either set theory or homological algebra. With this in mind, let us record here the places in the paper where more background knowledge may be required. First, some of the arguments in Section 3 presume some knowledge of derived functors and derived categories. That said, at the end of Section 3 we isolate an elementary combinatorial statement equivalent to the assertion that $\underline{A}$ is not Whitehead for a given abelian group $A$. It is this combinatorial statement that we work with for most of the remainder of the paper, so the reader may safely take the calculations of Section 3 as a black box. Sections 69 require some basic knowledge of forcing. Both [10] and [11] contain more than is needed here; we also direct the reader to [14], which provides a concise introduction to forcing aimed at non-specialists. The final sections are slightly more demanding, with Section 10 requiring a somewhat more sophisticated understanding of forcing and Sections 11 and 12 requiring some knowledge of homological algebra and condensed mathematics.
1.1. Notation and conventions
Our standard references for undefined notions and notations are [10] for set theory and [28] for homological algebra.
Given a group $G$, we will sometimes slightly abuse notation and write $G$ to refer to its underlying set. We let $|G|$ denote the cardinality of the underlying set of $G$. If $f$ is a function and $X\subseteq\mathrm{dom}(f)$, then $f[X]$ denotes the image of $X$ under $f$, i.e., $f[X]=\{f(x)\mid x\in X\}$.
Given a Boolean algebra $\mathbb{B}$, we let $S(\mathbb{B})$ denote the Stone space of $\mathbb{B}$. Concretely, $S(\mathbb{B})$ is the space of ultrafilters on $\mathbb{B}$, with basic clopen sets given by
$N_{b}:=\{U\in S(\mathbb{B})\mid b\in U\}$
for $b\in\mathbb{B}$.
2. Condensed abelian groups
[…]
Definition 2.1.
Let $\kappa$ be an uncountable cardinal. A $\kappa$-condensed set (resp. $\kappa$-condensed abelian group) is a contravariant functor $T:\mathsf{Prof}_{\kappa}\rightarrow\mathsf{Set}$ (resp. $T:\mathsf{Prof}_{\kappa}\rightarrow\mathsf{Ab}$) such that the following conditions hold.
  1. (1)
    $T(\emptyset)=\ast$.
  2. (2)
    For all $S_{0},S_{1}\in\mathsf{Prof}_{\kappa}$, the natural map
    $T(S_{0}\sqcup S_{1})\rightarrow T(S_{0})\times T(S_{1})$
    is a bijection.
  3. (3)
    Suppose that $S,S^{\prime}\in\mathsf{Prof}_{\kappa}$ and $f:S^{\prime}\rightarrow S$ is a surjective continuous map. Let $p_{0}$ and $p_{1}$ denote the two projections from $S^{\prime}\times_{S}S^{\prime}$ to $S^{\prime}$, and $p_{0}^{*}$ and $p_{1}^{*}$ their images under $T$. Then the natural map
    $T(S)\rightarrow\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$
    is a bijection.
[…]
Three sorts of choices of $\kappa$ in Definition 2.1 have predominated so far:
  • strong limit cardinals $\kappa$, valued for their closure properties in [20, 19, 3], with the associated categories $\mathsf{Cond}_{\kappa}$ assembling to form $\mathsf{Cond}$ in the fashion described just below;
  • inaccessible cardinals $\kappa$, valued for their even stronger closure properties in the closely related and contemporaneous pyknotic framework of Barwick and Haine [1];
  • $\kappa=\omega_{1}$, valued for its relative minimality in [4], wherein the associated condensed sets are termed light. We will return to this setting in Section 10. Note that $\mathsf{Prof}_{\omega_{1}}$ is precisely the category of totally disconnected metrizable compact Hausdorff spaces.
[…]
Given a Hausdorff space $X$, one can define a condensed set $\underline{X}$ by letting $\underline{X}(S)=C(S,X)$, i.e., $X(S)$ is the set of continuous functions from $S$ to $X$ for all $S\in\mathsf{Prof}$; we will sometimes term this $\underline{X}$ the condensation of $X$. This operation defines a functor from the category $\mathsf{Haus}$ of Hausdorff topological spaces to $\mathsf{Cond}$. When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf. [20, Proposition 1.7]).
Similarly, given a topological abelian group $A$, one can define a condensed abelian group $\underline{A}$ by letting $\underline{A}(S)=C(S,A)$ (considered as an abelian group with pointwise addition) for all $S\in\mathsf{Prof}$. When restricted to the class $\mathsf{LCA}$ of locally compact abelian groups, this again describes a fully faithful embedding of $\mathsf{LCA}$ into $\mathsf{CondAb}$.
For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$. Concretely, as noted in [20, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$. Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$. These facts straightforwardly propagate upward to the category $\mathsf{CondAb}$ of all condensed abelian groups.
3. Algebraic preliminaries
Given any abelian group $A$, a free resolution of $A$ is a short exact sequence of the form
[…]
in which both $F$ and $K$ are free abelian groups. In such resolutions, we will often implicitly assume that $K$ is a subgroup of $F$ and $\iota$ is the inclusion map. We note that every infinite abelian group $A$ has a canonical free resolution, in which $F$ is the free abelian group generated by the underlying set of $A$ and $K$ is the kernel of the induced surjection from $F$ onto $A$. The elements of $K$ are therefore the formal linear combinations of elements of $F$ that evaluate to 00 in $A$.
By applying the functor $\mathrm{Hom}(\cdot,\mathbb{Z})$ to a free resolution of $A$ and using the fact that $F$ and $K$ are free, and therefore Whitehead, we obtain an exact sequence of the form
$0\rightarrow\mathrm{Hom}(A,\mathbb{Z})\rightarrow\mathrm{Hom}(F,\mathbb{Z})\rightarrow\mathrm{Hom}(K,\mathbb{Z})\rightarrow\mathrm{Ext}^{1}(A,\mathbb{Z})\rightarrow 0$
We thus see that $A$ is Whitehead if and only if the map $\mathrm{Hom}(F,\mathbb{Z})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ is surjective, i.e., if and only if every group homomorphism $\varphi:K\rightarrow\mathbb{Z}$ extends to a homomorphism $\psi:F\rightarrow\mathbb{Z}$.
We now turn our attention to the situation in the category of condensed abelian groups. $\mathsf{CondAb}$ is a symmetric monoidal category with an internal Hom-functor, which we will denote by $\underline{\mathrm{Hom}}:\mathsf{CondAb}^{\mathrm{op}}\times\mathsf{CondAb}\rightarrow\mathsf{CondAb}$. If $A$ and $B$ are sufficiently nice topological groups, then $\underline{\mathrm{Hom}}(\underline{A},\underline{B})$ has a particularly nice description:
Proposition 3.1 (Clausen–Scholze, [20, Proposition 4.2]).
Suppose that $A$ and $B$ are Hausdorff topological abelian groups and $A$ is compactly generated. Then there is a natural isomorphism of condensed abelian groups
$\underline{\mathrm{Hom}}(\underline{A},\underline{B})\cong\underline{\mathrm{Hom}(A,B)},$
where, on the right-hand side, $\mathrm{Hom}(A,B)$ is computed in the category of topological abelian groups and endowed with the compact-open topology.
$\underline{\mathrm{Hom}}$ has corresponding derived functors, the internal Ext-functors
$\underline{\mathrm{Ext}}^{n}:\mathsf{CondAb}^{\mathrm{op}}\times\mathsf{CondAb}\rightarrow\mathsf{CondAb}$
for $n\geq 1$ and, at the level of the derived category, an internal $R\mathrm{Hom}$-functor
$R\underline{\mathrm{Hom}}:D(\mathsf{CondAb})^{\mathrm{op}}\times D(\mathsf{CondAb})\rightarrow D(\mathsf{CondAb});$
The former arise, of course, as the $n^{\mathrm{th}}$ cohomology groups of the latter. The following lemma shows that these functors behave as expected in relation to their classical analogues in $\mathsf{Ab}$; namely, the classical functors can be recovered by evaluating their condensed analogues at a point. When we need to be careful about the category in which certain functors are being considered, we will write the name of the category as a subscript on the functor, i.e., $\mathrm{Ext}^{1}_{\mathsf{Ab}}$ is the $\mathrm{Ext}$-functor in the category of abelian groups.
Lemma 3.2.
For any abelian groups $A$ and $B$ (with the discrete topology),
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{B})(\ast)=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B).$
More generally, for any condensed abelian groups $T$ and $M$, we have
$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(T,M)(\ast)=R\mathrm{Hom}_{\mathsf{CondAb}}(T,M),$
and, for discrete abelian groups $A$ and $B$, we in turn have
$R\mathrm{Hom}_{\mathsf{CondAb}}(\underline{A},\underline{B})=R\mathrm{Hom}_{\mathsf{Ab}}(A,B).$
Proof.
As noted in [20, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(T,M)(S)=R\mathrm{Hom}_{\mathsf{CondAb}}(T\otimes\mathbb{Z}[S],M).$
for any condensed abelian groups $T,M$ and any extremally disconnected profinite set $S$. When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$. To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [13, Props. 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$. Writing $\mathcal{P}$ and $\underline{\mathcal{P}}$ for $\cdots\to P^{1}\to P^{0}\to 0$ and $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to 0$, respectively, we then have
$R\mathrm{Hom}_{\mathsf{CondAb}}(\underline{A},\underline{B})=\mathrm{Hom}_{\mathsf{CondAb}}(\underline{\mathcal{P}},\underline{B})=\mathrm{Hom}_{\mathsf{Ab}}(\mathcal{P},B)=R\mathrm{Hom}_{\mathsf{Ab}}(A,B),$
as desired. ∎
It is natural at this point in the discussion to declare a condensed abelian group $T$ to be Whitehead if $\underline{\mathrm{Ext}}^{1}(T,\underline{\mathbb{Z}})=0$. Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$. To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [13, Proposition 2.19]. By [13, Remark 3.10 and Lemma 4.2], we then have
$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=\prod_{\kappa}R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})=\prod_{\kappa}\underline{\mathbb{Z}}[0],$
where ¯$\underline{\mathbb{Z}}[0]$ denotes the chain complex that is $\underline{\mathbb{Z}}$ in degree 00 and 00 elsewhere. In particular, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=0$.
One reasonable translation (though, as we will discuss later, not the only possible one) of Whitehead’s problem to the condensed setting is then the following:
Suppose that $A$ is a discrete abelian group and $\underline{A}$ is Whitehead in $\mathsf{CondAb}$. Must $A$ be free?
Note that, by Lemma 3.2, the assertion that $\underline{A}$ is Whitehead in $\mathsf{CondAb}$ is a strengthening of the assertion that $A$ is Whitehead in $\mathsf{Ab}$; the latter is equivalent to asserting that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(\ast)=0$, whereas the former is equivalent to asserting that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for every profinite set $S$. And, indeed, Theorem 1.2 implies that, unlike the classical Whitehead problem, this translation to the condensed setting has a positive answer in $\sf ZFC$.
In the coming sections, we will give a more concrete, combinatorial proof of Theorem 1.2. In particular, given a nonfree discrete abelian group $A$, we will identify a natural profinite set $S$ such that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$. The process by which this will be done is analogous to that described in the classical setting at the beginning of this section, so in the interest of symmetry, we end this section by giving an outline of the structure of the argument.
Suppose that $A$ is a nonfree discrete abelian group, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Applying $\underline{\mathrm{Hom}}(\cdot,\underline{\mathbb{Z}})$ to the image of this short exact sequence in $\mathsf{CondAb}$ yields the exact sequence
$0\rightarrow\underline{\mathrm{Hom}}(\underline{A},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{F},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{K},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})\rightarrow 0$
Now, evaluating this sequence at a profinite set $S$ and recalling Proposition 3.1, we obtain
$0\rightarrow\underline{\mathrm{Hom}(A,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\rightarrow 0$
It follows that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)=0$ if and only if the map
$\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$
is a surjection. Recall that $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)$ consists of all continuous functions $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ (and similarly for $K$), where $\mathrm{Hom}(F,\mathbb{Z})$ is given the compact-open topology (which, since $F$ and $\mathbb{Z}$ are discrete, is simply the topology $\mathrm{Hom}(F,\mathbb{Z})$ inherits as a subspace of $\prod_{F}\mathbb{Z}$ endowed with the product topology). Moreover, the map $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$ sends a continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ to the map $\phi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ defined by letting $\phi(s)=\psi(s)\restriction K$ for all $s\in S$.
Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$). The construction of such $S$ and $\varphi$ will be the subject of the next sections.
4. Countable groups
Recall that, if $A$ is a torsion-free abelian group and $X$ is a subgroup of $A$, we say that $X$ is a pure subgroup of $A$ if $A/X$ is torsion-free. Given any subgroup $X$ of $A$, the pure closure of $X$ in $A$ is the smallest pure subgroup of $A$ containing $X$; concretely, this can be seen to be equal to $\{a\in A\mid\exists n>0~{}[na\in X]\}$. The following theorem gives a useful characterization of countable free abelian groups (recall that a finitely generated abelian group is free if and only if it is torsion-free).
Theorem 4.1 ([17] (cf. [5, Theorem 4.2])).
Suppose that $A$ is a countable torsion-free abelian group. Then $A$ is free if and only if every finitely generated subgroup of $A$ is contained in a finitely generated pure subgroup of $A$.
As mentioned above, Stein [25] proved that every countable Whitehead group is free. The following theorem gives one way of seeing this (and slightly more).
Theorem 4.2.
Suppose that $A$ is a countable, torsion-free, nonfree abelian group,
[…]
is a free resolution of $A$, and $B_{K}$ is a basis for $K$. Then there is a homomorphism $\varphi:K\rightarrow\mathbb{Z}$ such that
  1. (1)
    $\varphi$ does not extend to a homomorphism $\psi:F\rightarrow\mathbb{Z}$; and
  2. (2)
    $\varphi[B_{K}]\subseteq\{0,1\}$.
Proof.
First, fix any basis $B_{F}$ for the free group $F$ and notice that $B_{K}$ is infinite: if not, then $K$ would be included in a subgroup $G$ of $F$ generated by a finite $S\subset B_{F}$; then $F=G\oplus H$ where $H$ is generated by $B_{F}\setminus S$, and we get $A\cong F/K\cong(G/K)\oplus H$ where $H$ is free and $G/K$ is finitely generated torsion-free, hence free, in contradiction with the nonfreeness of $A$.
Next, let us argue that, without loss of generality, we may assume that $F$ (and hence also $K$) is countable. To see this, suppose that $K$ is uncountable. Let $\chi$ be a sufficiently large regular cardinal, and let $N$ be a countable elementary submodel of $(H(\chi),\in,A,B_{F},B_{K})$. Let $F_{0}:=F\cap N$ and $K_{0}:=K\cap N$. By elementarity, $F_{0}$ and $K_{0}$ are free, with bases $B_{F}\cap N$ and $B_{K}\cap N$, respectively, and
$0\rightarrow K_{0}\xrightarrow{\subseteq}F_{0}\xrightarrow{\pi\restriction F_{0}}A\rightarrow 0$
is a free resolution of $A$ by countable free groups. Suppose that we can find a homomorphism $\varphi_{0}:K_{0}\rightarrow\mathbb{Z}$ such that $\varphi_{0}[B_{K}\cap N]\subseteq\{0,1\}$ and $\varphi_{0}$ does not extend to a homomorphism $\psi_{0}:F_{0}\rightarrow\mathbb{Z}$. Then we can extend $\varphi_{0}$ to a homomorphism $\varphi:K\rightarrow\mathbb{Z}$ by setting $\varphi(t)=0$ for all $t\in B_{K}\setminus N$. The fact that $\varphi_{0}$ does not extend to a homomorphism from $F_{0}$ to $\mathbb{Z}$ implies that $\varphi$ does not extend to a homomorphism from $F$ to $\mathbb{Z}$. Therefore, we may assume that $F$ is countable.
We can thus injectively enumerate $B_{K}$ as $\langle t_{n}\mid n<\omega\rangle$ and, for each $n<\omega$, let $K_{n}$ denote the subgroup of $K$ generated by $\{t_{k}\mid k<n\}$. We will define our homomorphism $\varphi:K\rightarrow\mathbb{Z}$ by recursion, requiring that $\varphi(t_{n})\in\{0,1\}$ for all $n<\omega$. We will do this by constructing a strictly increasing sequence $\langle n_{i}\mid i<\omega\rangle$ of natural numbers and, for each $i<\omega$, a homomorphism $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$ in such a way that $\varphi_{i}[\{t_{k}\mid k<n_{i}\}]\subseteq\{0,1\}$ and, for all $i<j<\omega$, $\varphi_{j}\restriction K_{n_{i}}=\varphi_{i}$.
By Theorem 4.1, since $A$ is countable, torsion-free, and nonfree, there is a finitely generated subgroup $X_{0}$ of $A$ such that its pure closure $X$ is not finitely generated. Let $Y_{0}$ be a finite subset of $B_{F}$ such that $X_{0}\subseteq\pi[F_{0}]$ where $F_{0}$ denotes the subgroup of $F$ generated by $Y_{0}$. Put $n_{0}=0$ and enumerate $\mathrm{Hom}(F_{0},\mathbb{Z})=\{\psi_{i}\mid i<\omega\}$. Finally, let $\varphi_{0}$ be the zero homomorphism. Assume that we have already defined $Y_{i}$, $F_{i}$, $n_{i}$ and $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$, for some $i<\omega$, in such a way that $K_{n_{i}}\subseteq F_{i}$. Consider any $z_{i}\in F$ such that $x_{i}:=\pi(z_{i})\in X\setminus\pi[F_{i}]$. Since $X$ is the pure closure of $X_{0}$, there is a positive integer $j_{i}>1$ such that $j_{i}x_{i}=\pi(j_{i}z_{i})\in X_{0}\subseteq\pi[F_{0}]$. It follows that there exists $y_{i}\in F_{0}$ such that $\pi(y_{i})=\pi(j_{i}z_{i})$, whence $s_{i}:=j_{i}z_{i}-y_{i}\in K=\mathrm{Ker}(\pi)$. We can thus express $s_{i}$ as $a_{i}+\sum_{n_{i}\leq k<n_{i+1}}b_{k}t_{k}$ where $a_{i}\in K_{n_{i}}$, $n_{i}<n_{i+1}<\omega$ and $b_{k}\in\mathbb{Z}$ for each $k\in[n_{i},n_{i+1})$. Let $Y_{i+1}$ be any finite subset of $B_{F}$ such that $Y_{i}\subseteq Y_{i+1}$ and $K_{n_{i+1}}\subseteq F_{i+1}$ where $F_{i+1}$ denotes the subgroup of $F$ generated by $Y_{i+1}$.
Claim 4.3.
There is $k\in[n_{i},n_{i+1})$ such that $j_{i}\nmid b_{k}$.
Proof.
Suppose not, and set
$s_{i}^{*}:=s_{i}-a_{i}=\sum_{k=n_{i}}^{n_{i+1}-1}b_{k}t_{k}.$
Then $s_{i}^{*}\in K$ is of the form $j_{i}z_{i}-r$, where $r\in F_{i}$. Moreover, since all of the coefficients $b_{k}$ for $k\in[n_{i},n_{i+1})$ are divisible by $j_{i}$, it follows that $s_{i}^{*}$ is divisible by $j_{i}$ (in $K$). Since $r=j_{i}z_{i}-s_{i}^{*}$, this means that there exists $r^{*}\in F_{i}$ such that $r=j_{i}r^{*}$. We get $s_{i}^{*}=j_{i}(z_{i}-r^{*})\in K$ which implies $z_{i}-r^{*}\in K$ using the fact that $A$ is torsion-free. In particular, $\pi(z_{i})=\pi(r^{*})$, so $\pi(z_{i})\in\pi[F_{i}]$, contradicting the fact that $\pi(z_{i})=x_{i}\notin\pi[F_{i}]$. ∎
Fix a $k_{i}\in[n_{i},n_{i+1})$ as given by the claim. To construct $\varphi_{i+1}$, it suffices to define $\varphi_{i+1}(t_{k})$ for all $k\in[n_{i},n_{i+1})$. For $k\in[n_{i},n_{i+1})\setminus\{k_{i}\}$, simply let $\varphi_{i+1}(t_{k})=0$. To define $\varphi_{i+1}(t_{k_{i}})$, consider the homomorphisms $\psi_{i}:F_{0}\rightarrow\mathbb{Z}$ and $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$ and the number
$d_{i}:=\varphi_{i}(a_{i})+\psi_{i}(y_{i}).$
If $d_{i}$ is divisible by $j_{i}$, then let $\varphi_{i+1}(t_{k_{i}})=1$. Otherwise, let $\varphi_{i+1}(t_{k_{i}})=0$.
Now let $\varphi:=\bigcup_{i<\omega}\varphi_{i}$. We claim that this is as desired. Clearly, $\varphi:K\rightarrow\mathbb{Z}$ is a homomorphism and $\varphi[B_{K}]\subseteq\{0,1\}$. We must therefore show that every homomorphism $\psi:F\rightarrow\mathbb{Z}$ fails to extend $\varphi$.
To this end, fix a homomorphism $\psi:F\rightarrow\mathbb{Z}$. Find $i<\omega$ such that $\psi\restriction F_{0}=\psi_{i}$. We claim that $\psi(s_{i})\neq\varphi(s_{i})$. Note that
$\psi(s_{i})=j_{i}\psi(z_{i})-\psi_{i}(y_{i}),$
and
$\varphi(s_{i})=b_{k_{i}}\varphi_{i+1}(t_{k_{i}})+\varphi_{i}(a_{i}).$
Therefore, if it were the case that $\psi(s_{i})=\varphi(s_{i})$, we would have
$j_{i}\psi(z_{i})$ $=b_{k_{i}}\varphi_{i+1}(t_{k_{i}})+\varphi_{i}(a_{i})+\psi_{i}(y_{i})$
$=b_{k_{i}}\varphi_{i+1}(t_{k_{i}})+d_{i}.$
Since $j_{i}\psi(z_{i})$ is clearly divisible by $j_{i}$, it follows that the right hand side of this equation must also be divisible by $j_{i}$. Suppose first that $d_{i}$ is divisible by $j_{i}$. Then $\varphi_{i+1}(t_{k_{i}})=1$, so, since $k_{i}$ was chosen so that $b_{k_{i}}$ is not divisible by $j_{i}$, it follows that the right hand side of the above equation is not divisible by $j_{i}$, which is a contradiction. If, on the other hand, $d_{i}$ is not divisible by $j_{i}$, then $\varphi_{i+1}(t_{k_{i}})=0$, so the right hand side of the above equation equals $d_{i}$ and is hence not divisible by $j_{i}$, which is again a contradiction. Therefore, it must be the case that $\psi(s_{i})\neq\varphi(s_{i})$. ∎
It is also well-known that all Whitehead groups are torsion-free. We now record a more general result, analogous to Theorem 4.2, implying this fact.
Proposition 4.4.
Let
[…]
be a short exact sequence of groups where $F$ is torsion-free and $K$ is free with a basis $B_{K}$. Assume that $A$ is not torsion-free (i.e. that $K$ is not a pure subgroup of $F$). Then there is a homomorphism $\varphi:K\rightarrow\mathbb{Z}$ such that
  1. (1)
    $\varphi$ does not extend to a homomorphism $\psi:F\rightarrow\mathbb{Z}$; and
  2. (2)
    there is a unique $t^{*}\in B_{K}$ such that
    • $\varphi(t^{*})=1$;
    • $\varphi(t)=0$ for all $t\in B_{K}\setminus\{t^{*}\}$.
Proof.
Being free, $K$ is isomorphic to the direct sum of copies of $\mathbb{Z}$ indexed by $B_{K}$. We consider the canonical embedding $\nu:K\to\mathbb{Z}^{B_{K}}$, i.e. for every $t,s\in B_{K}$, $\nu(t)(t)=1$ and $\nu(t)(s)=0$ if $s\neq t$. Then $\nu$ is pure, i.e. $\nu[K]$ is pure in $\mathbb{Z}^{B_{K}}$ (in fact, $\nu$ is even an elementary embedding). Since $K$ is not pure in $F$, it follows that there does not exist any homomorphism $\mu:F\to\mathbb{Z}^{B_{K}}$ such that $\mu\restriction K=\nu$.
For each $t\in B_{K}$, let us denote by $p_{t}:\mathbb{Z}^{B_{K}}\to\mathbb{Z}$ the $t$-th canonical projection. Since any homomorphism into direct product is uniquely determined by specifying its compositions with all the canonical projections, there has to be some $t^{*}\in B_{K}$ such that $\varphi:=p_{t^{*}}\nu$ cannot be extended to a homomorphism from $F$ into $\mathbb{Z}$. One readily checks that $t^{*}$ and $\varphi$ possess the desired properties. ∎
Remark 4.5.
In the formulation of Theorem 4.2, we could not have hoped to find $\varphi$ with $(\varphi\restriction B_{K})^{-1}\{1\}$ finite, let alone a $\varphi$ satisfying the stronger condition (2) from Proposition 4.4. The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf. [27, Proposition 3.1] or [18, Theorem 3.3.2]. Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
Let $B=\{e_{n}\mid n<\omega\}$ denote the canonical basis of the free group $\mathbb{Z}^{(\omega)}$. Then for $A$ as above, there exists a subgroup $F$ of $\mathbb{Z}^{\omega}$ such that the following diagram with exact rows, where $\mu$ and $\nu$ denote the respective identity embeddings, commutes.
[…]
It follows that $F$ is countable and thus free since all countable subgroups of $\mathbb{Z}^{\omega}$ are such, cf. [6, Theorems IV.2.3 and IV.2.8]. Since $\mu$ extends the map $\nu$, we see, in particular, that $p\mu$ extends the map $p\nu$ for any $p\in\mathrm{Hom}(\mathbb{Z}^{\omega},\mathbb{Z})$. This applies specifically for any $p$ which is a $\mathbb{Z}$-linear combination of the canonical projections $p_{n}:\mathbb{Z}^{\omega}\to\mathbb{Z}$, $n<\omega$. It remains to notice that any $\varphi\in\mathrm{Hom}(\mathbb{Z}^{(\omega)},\mathbb{Z})$ which is nonzero only on finitely many elements from $B$ is of the form $p\nu$ for $p$ as above.
On the other hand, since $\mathbb{Z}$ is a slender group ([6, Corollary III.2.4]), no homomorphism $\varphi:\mathbb{Z}^{(\omega)}\to\mathbb{Z}$ which is nonzero on infinitely many elements from $B$ can be extended to a homomorphism from $\mathbb{Z}^{\omega}$ into $\mathbb{Z}$. Of course, it can happen that a particular such $\varphi$ can be extended to a homomorphism from $F$ into $\mathbb{Z}$.
5. Uncountable groups
In what follows, given a cardinal $\kappa$, $2^{\kappa}$ denotes the topological space consisting of all functions $f:\kappa\rightarrow 2$, equipped with the product topology. In particular, basic clopen subsets of $2^{\kappa}$ are of the form $U_{\sigma}:=\{f\in 2^{\kappa}\mid\sigma\subseteq f\}$, where $\sigma$ is a finite partial function from $\kappa$ to $2$. Note that $2^{\kappa}$ is a profinite set.
Theorem 5.1.
Suppose that $A$ is a nonfree abelian group, and let $\kappa$ be the minimal cardinality of a nonfree subgroup of $A$. Suppose moreover that $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$. Then there is a continuous map $\varphi:2^{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that cannot be lifted pointwise to a continuous map $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$.
Moreover, given a basis $B_{K}$ for $K$, we can require that, for all $f\in 2^{\kappa}$ and all $t\in B_{K}$, we have $\varphi(f)(t)\in\{0,1\}$.
Proof.
If $A$ has nonzero torsion elements, then Proposition 4.4 provides us with a homomorphism $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ such that $\tau[B_{K}]\subseteq\{0,1\}$ and $\tau$ cannot be lifted to an element of $\mathrm{Hom}(F,\mathbb{Z})$. We can in this case let $\varphi:2^{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be the constant map taking value $\tau$. Thus, assume for the rest of the proof that $A$ is torsion-free.
[…]
We first claim that we may assume both that $A$ is almost free, i.e., that $\kappa=|A|$, and that $|F|=\kappa$. To see this, let $N$ be an elementary submodel of $(H(\chi),\in,\vartriangleleft,$ $A,B_{F},B_{K})$ with $|N|=\kappa\subseteq N$. Let $A_{0}:=A\cap N$. By elementarity, $N$ contains as an element a nonfree subgroup of $A$ of cardinality $\kappa$. Since $\kappa\subseteq N$, this subgroup is a subset of $A_{0}$; therefore, $A_{0}$ is nonfree, and, by the minimality of $\kappa$, almost free. Let $F_{0}:=F\cap N$ and $K_{0}:=K\cap N$; by elementarity, $B_{F,0}:=B_{F}\cap N$ and $B_{K,0}:=B_{K}\cap N$ are bases for $F_{0}$ and $K_{0}$, respectively, and $0\rightarrow K_{0}\rightarrow F_{0}\rightarrow A_{0}\rightarrow 0$ is a free resolution of $A_{0}$. Moreover, we clearly have $|F_{0}|=\kappa$.
Suppose that we can find a continuous $\varphi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(K_{0},\mathbb{Z})$ that does not lift pointwise to a continuous map $\psi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(F_{0},\mathbb{Z})$, and assume moreover that, for all $f\in 2^{\kappa}$ and all $t\in B_{K,0}$, we have $\varphi_{0}(f)(t)\in\{0,1\}$. Then $\varphi_{0}$ can be extended pointwise to a continuous map $\varphi:2^{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ by setting
$\varphi(f)(t)=\begin{cases}\varphi_{0}(f)(t)&\text{if }t\in B_{K,0}\\ 0&\text{otherwise}\end{cases}$
for all $t\in B_{K}$ and extending linearly. The continuity of $\varphi$ follows from the continuity of $\varphi_{0}$, and our construction ensures that $\varphi(f)(t)\in\{0,1\}$ for all $f\in 2^{\kappa}$ and all $t\in B_{K}$. Moreover, $\varphi$ does not lift pointwise to a continuous $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$, as pointwise restriction of such a $\psi$ to $F_{0}$ would induce a continuous pointwise extension $\psi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(F_{0},\mathbb{Z})$ of $\varphi_{0}$. We can therefore assume without loss of generality that $A$ is almost free.
Thus, for the rest of the proof we may assume that $A$ is almost free and $|F|=\kappa$. The case of $\kappa=\aleph_{0}$ is covered by Theorem 4.2, so we may assume that $\kappa>\aleph_{0}$. Let $\vec{M}=\langle M_{\alpha}\mid\alpha<\kappa\rangle$ be a $\in$-increasing, continuous sequence of elementary submodels of $(H(\chi),\in,\vartriangleleft,A,F,K)$ such that $|M_{\alpha}|<\kappa$ and $\delta_{\alpha}:=M_{\alpha}\cap\kappa\in\kappa$ for all $\alpha<\kappa$. For each $\alpha<\kappa$, let $A_{\alpha}:=A\cap M_{\alpha}$, $F_{\alpha}:=F\cap M_{\alpha}$, and $K_{\alpha}:=K\cap M_{\alpha}$. Since $A$, $F$, and $K$ each have size at most $\kappa$ and $\vec{M}$ is $\in$-increasing, we have $A=\bigcup_{\alpha<\kappa}A_{\alpha}$, $F=\bigcup_{\alpha<\kappa}F_{\alpha}$, and $K=\bigcup_{\alpha<\kappa}K_{\alpha}$.
Let $E$ be the set of $\alpha<\kappa$ such that there exists $\beta$ such that $\alpha<\beta<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree. Note that, if $\alpha<\beta<\beta^{\prime}<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree, then also $A_{\beta^{\prime}}/A_{\alpha}$ is nonfree, since it contains $A_{\beta}/A_{\alpha}$ as a subgroup. We claim that $E$ is stationary. If not, then there exists a club $D\subseteq\kappa$ such that, for all $\alpha<\beta$, both in $D$, the group $A_{\beta}/A_{\alpha}$ is free. Moreover, since $A$ is almost free, we know that $A_{\min(D)}$ is free. Therefore, it is straightforward to build a basis for $A$ by recursion along $D$ (cf. [5, Theorem 5.3]), contradicting the assumption that $A$ is not free.
For all $\alpha\in E$, let $\alpha<\beta_{\alpha}<\kappa$ be such that $A_{\beta_{\alpha}}/A_{\alpha}$ is not free. Let $D:=\{\gamma<\kappa\mid\forall\alpha<\gamma\;[\beta_{\alpha}<\gamma]\}$. Then $D$ is a club in $\kappa$. Letting $\langle\gamma_{\eta}\mid\eta<\kappa\rangle$ be the increasing enumeration of $D$, the following two facts are immediate:
  1. (1)
    $\{\eta<\kappa\mid\gamma_{\eta}\in E\}$ is stationary in $\kappa$;
  2. (2)
    for all $\eta<\kappa$ for which $\gamma_{\eta}\in E$, the set $A_{\gamma_{\eta+1}}/A_{\gamma}$ is not free.
Therefore, by thinning out $\vec{M}$ to only include those models indexed by elements of $D$ and then reindexing, we may assume that the set
$S:=\{\alpha<\kappa\mid A_{\alpha+1}/A_{\alpha}\text{ is nonfree}\}$
is stationary in $\kappa$.
For each $\alpha\in S$, let $A_{\alpha}^{*}:=A_{\alpha+1}/A_{\alpha}$, $F_{\alpha}^{*}:=F_{\alpha+1}/F_{\alpha}$, and $K_{\alpha}^{*}:=K_{\alpha+1}/K_{\alpha}$. Similarly, let $B_{F,\alpha}^{*}:=B_{F}\cap(M_{\alpha+1}\setminus M_{\alpha})$ and $B_{K,\alpha}^{*}:=B_{K}\cap(M_{\alpha+1}\setminus M_{\alpha})$. Note that each of $F_{\alpha}^{*}$ and $K_{\alpha}^{*}$ is free, with bases given by $\{z+F_{\alpha}\mid z\in B_{F,\alpha}^{*}\}$ and $\{t+K_{\alpha}\mid t\in B_{K,\alpha}^{*}\}$, respectively. Also, our free resolution of $A$ induces a free resolution
$0\rightarrow K_{\alpha}^{*}\rightarrow F_{\alpha}^{*}\rightarrow A_{\alpha}^{*}\rightarrow 0$
of $A_{\alpha}^{*}$. Since $\alpha\in S$, we know that $A_{\alpha}^{*}$ is nonfree. Let $\kappa_{\alpha}:=|A_{\alpha}^{*}|<\kappa$, and apply the inductive hypothesis to find a continuous $\varphi_{\alpha}:2^{\kappa_{\alpha}}\rightarrow\mathrm{Hom}(K_{\alpha}^{*},\mathbb{Z})$ that does not lift pointwise to a continuous $\psi_{\alpha}:2^{\kappa_{\alpha}}\rightarrow\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$. Note that we are slightly abusing terminology here, since $K_{\alpha}^{*}$ is not formally a subgroup of $F_{\alpha}^{*}$. Nonetheless, we identify $K_{\alpha}^{*}$ with the subgroup $\{r+F_{\alpha}\mid r\in K_{\alpha+1}\}$, so our assumption is formally that there is no continuous $\psi_{\alpha}:2^{\kappa_{\alpha}}\rightarrow\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ such that, for all $f\in 2^{\kappa_{\alpha}}$ and all $r\in K_{\alpha+1}$, we have $\psi_{\alpha}(f)(r+F_{\alpha})=\varphi_{\alpha}(f)(r+K_{\alpha})$.
We are now ready to describe the construction of our desired continuous map $\varphi:2^{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$. We first need a bit of notation: given $f\in 2^{\kappa}$ and $\alpha\in S$, define a map $f_{\alpha}\in 2^{\kappa_{\alpha}}$ by letting $f_{\alpha}(\eta)=f(\alpha+1+\eta)$ for all $\eta<\kappa_{\alpha}$. Now, to define $\varphi$, it suffices to specify $\varphi(f)(t)$ for all $f\in 2^{\kappa}$ and $t\in B_{K}$, and then extend each $\varphi(f)$ linearly to all of $K$. To do this, suppose we are given $f\in 2^{\kappa}$ and $t\in B_{K}$. If there is $\alpha\in S$ such that $f(\alpha)=1$ and $t\in B^{*}_{K,\alpha}$, [^2] Note that such an $\alpha$ must be unique, if it exists. then let
$\varphi(f)(t)=\varphi_{\alpha}(f_{\alpha})(t+K_{\alpha}).$
In all other cases, let $\varphi(f)(t)=0$.
The continuity of $\varphi:2^{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ follows immediately from the continuity of $\varphi_{\alpha}$ for each $\alpha\in S$. Moreover, our construction, and the analogous fact about each $\varphi_{\alpha}$, ensures that $\varphi(f)(t)\in\{0,1\}$ for each $f\in 2^{\kappa}$ and $t\in B_{K}$. It thus remains to show that $\varphi$ cannot be lifted pointwise to a continuous $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$.
[…]
By the stationarity of $S$ and the fact that $\kappa$ is a regular, uncountable cardinal, we can fix $\alpha\in S$ such that $\delta_{\alpha}=\alpha$ and $v(\gamma)\subseteq\alpha$ for all $\gamma<\alpha$. We will complete the proof by constructing a continuous map $\psi_{\alpha}:2^{\kappa_{\alpha}}\rightarrow\mathrm{Hom}(F^{*}_{\alpha},\mathbb{Z})$ that lifts $\varphi_{\alpha}$ pointwise, contradicting our choice of $\varphi_{\alpha}$.
We first need a bit more notation. Given a function $g\in 2^{\kappa_{\alpha}}$ and an $i<2$, define a function $g^{+}_{i}\in 2^{\kappa}$ as follows. First, let $g^{+}_{i}(\alpha)=i$. Then, for all $\eta<\kappa_{\alpha}$, let $g^{+}_{i}(\alpha+1+\eta)=g(\eta)$. Finally, let $g^{+}_{i}(\beta)=0$ for all $\beta\in\kappa\setminus[\alpha,\alpha+\kappa_{\alpha})$. Note that, for $i<2$, we have $(g^{+}_{i})_{\alpha}=g$.
As usual, to define $\psi_{\alpha}$, it suffices to specify $\psi_{\alpha}(g)(z+F_{\alpha})$ for all $g\in 2^{\kappa_{\alpha}}$ and $z\in B^{*}_{F,\alpha}$. To this end, given such $g$ and $z$, let
$\psi_{\alpha}(g)(z+F_{\alpha})=\psi(g^{+}_{1})(z)-\psi(g^{+}_{0})(z).$
We will be done if we show that $\psi_{\alpha}(g)(t+F_{\alpha})=\varphi_{\alpha}(g)(t+K_{\alpha})$ for all $g\in 2^{\kappa_{\alpha}}$ and $t\in B^{*}_{K,\alpha}$. To this end, fix such $g$ and $t$. Let $B_{F,\alpha}:=\{z_{\gamma}\mid\gamma<\alpha\}=B_{F}\cap M_{\alpha}$, and note that $B_{F,\alpha}\cup B^{*}_{F,\alpha}$ is a basis for $F_{\alpha+1}$. Since $t\in K_{\alpha+1}\subseteq F_{\alpha+1}$, we can find finite subsets $I^{-}\subseteq B_{F,\alpha}$ and $I^{+}\subseteq B^{*}_{F,\alpha}$ together with integer coefficients $\{a_{z}\mid z\in I^{-}\cup I^{+}\}$ such that, in $F_{\alpha+1}$, we have
$t=\sum_{z\in I^{-}\cup I^{+}}a_{z}z.$
Let
$t^{-}:=\sum_{z\in I^{-}}a_{z}z\text{\ \ and \ }t^{+}:=\sum_{z\in I^{+}}a_{z}z.$
The following observations are immediate, but crucial:
  • $t=t^{-}+t^{+}$;
  • $t^{-}\in F_{\alpha}$;
  • $t^{+}$ is a linear combination of elements of $B^{*}_{F,\alpha}$; in consequence, we have $\psi_{\alpha}(g)(t^{+}+F_{\alpha})=\psi(g_{1}^{+})(t^{+})-\psi(g_{0}^{+})(t^{+})$.
Since $g_{1}^{+}\restriction\alpha=g_{0}^{+}\restriction\alpha$, our choice of $\alpha$ ensures that $\psi(g_{1}^{+})\restriction F_{\alpha}=\psi(g_{0}^{+})\restriction F_{\alpha}$, and hence in particular $\psi(g_{1}^{+})(t^{-})=\psi(g_{0}^{+})(t^{-})$. But now, putting all of this together we obtain
$\psi_{\alpha}(g)(t+F_{\alpha})$ $=\psi_{\alpha}(g)(t^{+}+F_{\alpha})$
$=\psi(g_{1}^{+})(t^{+})-\psi(g_{0}^{+})(t^{+})$
$=\psi(g_{1}^{+})(t^{-})+\psi(g_{1}^{+})(t^{+})-(\psi(g_{0}^{+})(t^{-})+\psi(g_{0}^{+})(t^{+}))$
$=\psi(g_{1}^{+})(t)-\psi(g_{0}^{+})(t)$
$=\varphi(g_{1}^{+})(t)-\varphi(g_{0}^{+})(t)$
$=\varphi_{\alpha}(g)(t+K_{\alpha})-0=\varphi_{\alpha}(g)(t+K_{\alpha}).$
In the above equation, the first equality holds because $t=t^{-}+t^{+}$ and $t^{-}\in F_{\alpha}$, and hence $t+F_{\alpha}=t^{+}+F_{\alpha}$. The second follows from the definition of $\psi$, the third from the observation above that $\psi(g_{1}^{+})(t^{-})=\psi(g_{0}^{+})(t^{-})$, the fourth by linearity, the fifth from the assumption that $\psi$ lifts $\varphi$ pointwise, and the sixth from the definition of $\varphi$, noting that $g^{+}_{i}(\alpha)=i$ for $i<2$. This shows that $\psi_{\alpha}$ lifts $\varphi_{\alpha}$ pointwise, furnishing the desired contradiction and completing the proof. ∎
6. Cohen forcing
Given an infinite set $X$, we let $\mathrm{Add}(\omega,X)$ denote the forcing to add $|X|$-many Cohen reals. Concretely, $\mathrm{Add}(\omega,X)$ consists of all finite partial functions from $X$ to $\{0,1\}$, ordered by reverse inclusion.
In this section, we will modify the arguments of the previous section to prove the following theorem:
Theorem 6.1.
Suppose that $A$ is a nonfree abelian group and $\kappa=|A|$. Then, in $V^{\mathrm{Add}(\omega,\kappa)}$, $A$ is not Whitehead.
Before proving Theorem 6.1, we make the following observation.
Lemma 6.2.
Suppose that $\mathbb{P}$ is a ccc forcing notion and $A$ is an abelian group. Then $A$ is free in $V$ if and only if $A$ is free in $V^{\mathbb{P}}$.
Proof.
If $A$ is free in $V$, then it clearly remains free in $V^{\mathbb{P}}$, and if $A$ is not torsion-free, then it is nonfree in both models. Therefore, we may assume that $A$ is torsion-free and nonfree in $V$, and we will show that $A$ remains nonfree in $V^{\mathbb{P}}$.
The proof is by induction on $|A|$. Suppose first that $A$ is countable. Then, by Theorem 4.1, $A$ has a finitely generated subgroup $X$ whose pure closure is not finitely generated. Since $A$ has the same finitely generated subgroups in $V$ and in $V^{\mathbb{P}}$, and since a subgroup of $A$ is pure in $V$ if and only if it is pure in $V^{\mathbb{P}}$, $X$ retains these properties in $V^{\mathbb{P}}$, so, again by Theorem 4.1, $A$ is nonfree in $V^{\mathbb{P}}$.
Now suppose that $|A|=\kappa>\aleph_{0}$. First, note that we may assume that $A$ is almost free. Indeed, if $A_{0}$ is a nonfree subgroup of $A$ and $|A_{0}|<\kappa$, then we can apply the inductive hypothesis to conclude that $A_{0}$, and hence also $A$, remains nonfree in $V^{\mathbb{P}}$.
Thus, assume that $A$ is almost free, and hence $\kappa$ is regular. Since $A$ is nonfree, we can fix a $\subseteq$-increasing, continuous sequence $\vec{A}:=\langle A_{\alpha}\mid\alpha<\kappa\rangle$ such that
  1. (1)
    each $A_{\alpha}$ is a subgroup of $A$ of cardinality less than $\kappa$;
  2. (2)
    $A=\bigcup_{\alpha<\kappa}A_{\alpha}$;
  3. (3)
    the set $S:=\{\alpha<\kappa\mid A_{\alpha+1}/A_{\alpha}\text{ is nonfree}\}$ is stationary in $\kappa$.
Items (1) and (2) are clearly upward absolute to $V^{\mathbb{P}}$. By the inductive hypothesis, for all $\alpha\in S$, the group $A_{\alpha+1}/A_{\alpha}$ remains nonfree in $V^{\mathbb{P}}$. Moreover, since $\mathbb{P}$ has the ccc, $S$ remains stationary in $\kappa$, so (3) is also upward absolute. Thus, $\vec{A}$ continues to witness that $A$ nonfree in $V^{\mathbb{P}}$. ∎
Given infinite cardinals $\kappa\leq\lambda$, note that $\mathrm{Add}(\omega,\kappa)$ is a regular suborder of $\mathrm{Add}(\omega,\lambda)$, and $\mathrm{Add}(\omega,\lambda)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,[\kappa,\lambda))$. To establish Theorem 6.1, we will actually prove the following formally stronger statement.
Theorem 6.3.
Suppose that $A$ is a nonfree abelian group, $\kappa$ is the least cardinality of a nonfree subgroup of $A$,
$0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$
is a free resolution of $A$, and $B_{K}$ is a basis for $K$. Then, in $V^{\mathrm{Add}(\omega,\kappa)}$, there is a $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ such that
  1. (1)
    $\varphi[B_{K}]\subseteq\{0,1\}$;
  2. (2)
    in every ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$, there is no $\psi\in\mathrm{Hom}(F,\mathbb{Z})$ extending $\varphi$.
Proof.
The proof is by induction on $|A|$; since the proof resembles that of Theorem 5.1, we will omit some details and frequently refer the reader back to that earlier proof. If $A$ has nonzero torsion elements or is torsion-free and countable, then the homomorphism $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ constructed (in $V$) in the proof of Proposition 4.4 or Theorem 4.2, respectively, satisfies the conclusion of the theorem, since, as the reader may readily verify in either case, the argument that $\varphi$ cannot be extended to an element of $\mathrm{Hom}(F,\mathbb{Z})$ can be carried out equally well in any outer model of $V$.
We may thus assume that $\kappa>\aleph_{0}$ and $A$ is torsion-free. As in the proof of Theorem 5.1, we may assume that $A$ is almost free and $|F|=\kappa$. To see this, let $N$, $A_{0}$, $F_{0}$, $K_{0}$, $B_{F,0}$, and $B_{K,0}$ be as in the analogous argument in the earlier proof. Suppose that we are able to obtain, in $V^{\mathrm{Add}(\omega,\kappa)}$, a $\varphi_{0}\in\mathrm{Hom}(K_{0},\mathbb{Z})$ such that $\varphi_{0}[B_{K,0}]\subseteq\{0,1\}$ and $\varphi_{0}$ cannot be extended to a map $\psi_{0}\in\mathrm{Hom}(F_{0},\mathbb{Z})$ in any ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$. Extend $\varphi_{0}$ to a map $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ by letting $\varphi(t)=0$ for all $t\in B_{K}\setminus B_{K,0}$. Then $\varphi\in V^{\mathrm{Add}(\omega,\kappa)}$. Moreover, if $\dot{\mathbb{P}}$ is an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, then $\varphi$ does not extend in $V^{\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}}$ to an element of $\mathrm{Hom}(F,\mathbb{Z})$, as the restriction of such an extension to $F_{0}$ would extend $\varphi_{0}$, contradicting our choice of $\varphi_{0}$.
Assume now that $A$ is almost free and $|F|=\kappa$. For $\alpha<\kappa$, let $M_{\alpha}$, $\delta_{\alpha}$, $A_{\alpha}$, $F_{\alpha}$, and $K_{\alpha}$ be as in the proof of Theorem 5.1. Let
$S:=\{\alpha<\kappa\mid A_{\alpha+1}/A_{\alpha}\text{ is nonfree}\};$
again, since $A$ is almost free and nonfree, we may assume that $S$ is stationary in $\kappa$. For each $\alpha\in S$, let $\kappa_{\alpha}$, $A_{\alpha}^{*}$, $F_{\alpha}^{*}$, $K_{\alpha}^{*}$, $B^{*}_{F,\alpha}$, and $B^{*}_{K,\alpha}$ be as in the proof of Theorem 5.1.
Note that $\mathrm{Add}(\omega,\kappa)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,\kappa)$, and hence we can assume that we are in fact working in an extension $V_{1}$ of the ground model $V$ by (the first copy of) $\mathrm{Add}(\omega,\kappa)$. In particular, by the inductive hypothesis, we can assume that, for every $\alpha\in S$, we have a $\varphi_{\alpha}\in\mathrm{Hom}(K_{\alpha}^{*},\mathbb{Z})$ such that
  • $\varphi_{\alpha}[\{t+K_{\alpha}\mid t\in B^{*}_{K,\alpha}\}]\subseteq\{0,1\}$;
  • in every ccc forcing extension of $V_{1}$, there is no $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$.
We now describe a method for producing elements of $\mathrm{Hom}(K,\mathbb{Z})$. Given a function $f:\kappa\rightarrow\{0,1\}$, define $\varphi(f)\in\mathrm{Hom}(K,\mathbb{Z})$ as follows. For every $t\in B_{K}$, if there is $\alpha\in S$ such that $t\in B^{*}_{K,\alpha}$ and $f(\alpha)=1$, then let $\varphi(f)(t)=\varphi_{\alpha}(t+K_{\alpha})$. In all other cases, let $\varphi(f)(t)=0$. Then extend linearly to all of $K$. By construction, we have $\varphi(f)[B_{K}]\subseteq\{0,1\}$ for all $f:\kappa\rightarrow\{0,1\}$.
Note that this construction continues to make sense in any outer model of $V_{1}$. Let $\dot{g}$ be the canonical $\mathrm{Add}(\omega,\kappa)$-name for the union of the generic filter, and note that $\dot{g}$ is a name for a function from $\kappa$ to $\{0,1\}$. Then let $\dot{\varphi}$ be an $\mathrm{Add}(\omega,\kappa)$-name for $\varphi(\dot{g})$. We claim that $\dot{\varphi}$ is forced to be as desired. To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
[…]
  • $\delta_{\alpha}=\alpha$;
  • for all $\gamma<\alpha$, we have $\mathrm{dom}(p_{\gamma},q_{\gamma})\subseteq\alpha$.
For each $p\in\mathrm{Add}(\omega,\kappa)$, define $\hat{p}\in\mathrm{Add}(\omega,\kappa)$ by letting $\mathrm{dom}(\hat{p})=\mathrm{dom}(p)$ and, for all $\beta\in\mathrm{dom}(p)$, letting
$\hat{p}(\beta)=\begin{cases}p(\beta)&\text{if }\beta\neq\alpha\\ 1-p(\beta)&\text{if }\beta=\alpha.\end{cases}$
Since the map $p\mapsto\hat{p}$ is an automorphism of $\mathrm{Add}(\omega,\kappa)$, the filter $\hat{G}=\{\hat{p}\mid p\in G\}$ is generic over $V_{1}$, and the function $\hat{g}=\bigcup\hat{G}$ agrees with $g$ everywhere except $\alpha$, where it takes the opposite value. Without loss of generality, assume that $g(\alpha)=0$ and $\hat{g}(\alpha)=1$. For ease of comprehension, we then denote $(G,g)$ by $(G_{0},g_{0})$ and $(\hat{G},\hat{g})$ by $(G_{1},g_{1})$. Note also that $V[G_{0}]=V[G_{1}]$, and hence $H$ is $\mathbb{P}$-generic over both. For $i<2$, let $\varphi_{i}:=\varphi(g_{i})=\dot{\varphi}^{G_{i}}$, and let $\psi_{i}:=\dot{\psi}^{G_{i}\ast H}$.
By our assumptions about $\dot{\psi}$, we know that, for each $i<2$, $\psi_{i}\in\mathrm{Hom}(F,\mathbb{Z})$ extends $\varphi_{i}$. Moreover, by our choice of $\alpha$ and the fact that $g_{0}\restriction\alpha=g_{1}\restriction\alpha$, we know that $\psi_{0}\restriction F_{\alpha}=\psi_{1}\restriction F_{\alpha}$. Now define a homomorphism $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ as follows. For all $z\in B^{*}_{F,\alpha}$, let $\psi_{\alpha}(z+F_{\alpha})=\psi_{1}(z)-\psi_{0}(z)$, and then extend linearly to the rest of $F_{\alpha}^{*}$. We claim that $\psi_{\alpha}$ extends $\varphi_{\alpha}$. The proof of this is exactly the same as that of the analogous fact in the proof of Theorem 5.1, so we leave it to the reader. But then, in $V_{1}[G\ast H]$, we have an element of $\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$, contradicting our choice of $\varphi_{\alpha}$ and the fact that $V_{1}[G\ast H]$ is a ccc extension of $V_{1}$. ∎
7. Stone spaces
In this section, we take a direct route to the proof of clause (3) of Theorem 1.3. A more indirect route, involving a broader investigation into more general connections between condensed mathematics and forcing, occupies Sections 8 and 9, which can be read independently of this one.
Given an infinite cardinal $\kappa$, let $\mathbb{B}_{\kappa}$ denote the Boolean completion of $\mathrm{Add}(\omega,\kappa)$, and let $S_{\kappa}$ denote the Stone space $S(\mathbb{B}_{\kappa})$. We will sometimes slightly abuse notation and think of $\mathrm{Add}(\omega,\kappa)$ as a subset of $\mathbb{B}_{\kappa}$. Given sets $X$ and $Y$, we now highlight a translation between $\mathbb{B}_{\kappa}$-names for functions from $X$ to $Y$ and continuous functions from $S_{\kappa}$ to the product space $Y^{X}$. Recall that a clopen basis for $S_{\kappa}$ is given by all sets of the form
$N_{b}:=\{U\in S_{\kappa}\mid b\in U\}$
for $b\in\mathbb{B}_{\kappa}$.
Definition 7.1.
Suppose that $\kappa$ is an infinite cardinal and $X$ and $Y$ are nonempty sets.
  1. (1)
    Suppose that $\varphi:S_{\kappa}\rightarrow Y^{X}$ is continuous, where $Y$ is discrete and $Y^{X}$ is given the product topology. Define a $\mathbb{B}_{\kappa}$-name $\dot{f}_{\varphi}$ for a function from $X$ to $Y$ as follows. Given $x\in X$ and $y\in Y$, let $U_{x,y}:=\{h\in Y^{X}\mid h(x)=y\}$, and set
    $\llbracket\dot{f}_{\varphi}(x)=y\rrbracket_{\mathbb{B}_{x}}:=\bigvee\{b\in\mathbb{B}_{\kappa}\mid N_{b}\subseteq\varphi^{-1}[U_{x,y}]\}.$
  2. (2)
    Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$. Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$. Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
It is routine to verify that $\dot{f}_{\varphi}$ as defined in Definition 7.1(1) is indeed a $\mathbb{B}_{\kappa}$ name for a function from $X$ to $Y$ and that $\varphi_{\dot{f}}$ as defined in Definition 7.1(2) is a continuous function from $S_{\kappa}$ to $Y^{X}$. (These facts will also follow from more general arguments appearing below in Section 8.) This translation, together with Theorem 6.3, allows us to replace the space $2^{\kappa}$ in Theorem 5.1 with the Stone space $S_{\kappa}$.
Theorem 7.2.
Suppose that $A$ is a nonfree abelian group and $\kappa=|A|$. Suppose moreover that $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ is a free resolution of $A$. Then there is a continuous map $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that cannot be lifted pointwise to a continuous map $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$.
Moreover, given a basis $B_{K}$ for $K$, we can require that, for all $U\in S_{\kappa}$ and all $t\in B_{K}$, we have $\varphi(U)(t)\in\{0,1\}$.
Proof.
By Theorem 6.3, we can fix an $\mathrm{Add}(\omega,\kappa)$-name $\dot{\tau}$ for an element of $\mathrm{Hom}(K,\mathbb{Z})$ for which it is forced to be the case that
  • $\dot{\tau}[B_{K}]\subseteq\{0,1\}$; and
  • there is no $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ that extends $\dot{\tau}$.
We think of $\dot{\tau}$ as a $\mathbb{B}_{\kappa}$-name. Let $\dot{f}$ be the $\mathbb{B}_{\kappa}$-name for the restriction of $\tau$ to $B_{K}$; $\dot{f}$ is thus a name for a function from $B_{K}$ to $2$. Let $\varphi_{\dot{f}}$ be the continuous function from $S_{\kappa}$ to $2^{B_{K}}$ given Definition 7.1(2), and let $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be defined by letting $\varphi(U)$ be the linear extension of $\varphi_{\dot{f}}(U)$ to all of $K$. Then $\varphi$ is also a continuous function.
We claim that $\varphi$ is as desired. We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$. Suppose for sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise. Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
Claim 7.3.
$\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}\in\mathrm{Hom}(F,\mathbb{Z})$.
Proof.
Suppose for sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$. By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(y)=k_{y}$, $\psi(U)(z)=k_{z}$, and $\psi(U)(y+z)=k_{yz}$, contradicting the fact that $\psi(U)\in\mathrm{Hom}(F,\mathbb{Z})$. ∎
Claim 7.4.
$\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}\restriction K=\dot{\tau}$.
Proof.
Suppose for sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$. By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(U)(t)=\varphi(U)(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise. ∎
But now we have shown that there is forced by $\mathbb{B}_{\kappa}$ to be a $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ (namely, $\dot{f}_{\psi}$) that extends $\dot{\tau}$, which is a contradiction, thus completing the proof. ∎
8. Topological spaces in forcing extensions
In this section and the next, we begin a general investigation into connections between condensed mathematics and forcing, the results of which will yield clause (3) of Theorem 1.3 as a corollary.
Suppose that $(X,\tau)$ is a compact Hausdorff space and $W$ is an outer model of $V$. Then there is a canonical way to interpret $(X,\tau)$ as a compact Hausdorff space $(\hat{X},\hat{\tau})$ in $W$. First, let $\tau_{c}$ denote the collection of all closed subsets of $X$ (in $V$). In $W$, let $\hat{X}$ denote the collection of all maximal filters on $\tau_{c}\setminus\{\emptyset\}$. In other words, $\hat{X}$ is the collection of all maximal collections of elements of $\tau_{c}$ with the finite intersection property.
There is a natural map $\pi:X\rightarrow\hat{X}$ defined by letting $\pi(x):=\{D\in\tau_{c}\mid x\in D\}$ for all $x\in X$. We also define $\pi:\tau\rightarrow\mathscr{P}(\hat{X})$ by letting $\pi(U):=\{F\in\hat{X}\mid(X\setminus U)\notin F\}$. Then let $\hat{\tau}$ be the topology on $\hat{X}$ generated by $\pi[\tau]$.
Remark 8.1.
The space $(\hat{X},\hat{\tau})$ identified above, together with the associated maps $\pi:X\rightarrow\hat{X}$ and $\pi:\tau\rightarrow\hat{\tau}$ is precisely the interpretation of $(X,\tau)$ in $W$, in the sense of [29]. In particular, this implies the following properties, which can also readily be directly verified:
  • $(\hat{X},\hat{\tau})$ is compact in $W$;
  • $\pi[\tau]$ is a basis for $\hat{\tau}$;
  • for all $x\in X$ and $U\in\tau$, we have $x\in U$ if and only if $\pi(x)\in\pi(U)$.
Now suppose that $\mathbb{B}$ is a complete Boolean algebra and $S=S(\mathbb{B})$ is its Stone space. We now show how to identify the space $(\hat{X},\hat{\tau})$, as computed in the forcing extension by $\mathbb{B}$, with an appropriate quotient of the space $C(S,X)^{V}$, equipped with the compact-open topology. In fact, we will prove something more general.
We first establish a correspondence between elements of $C(S,X)$ and $\mathbb{B}$-names for points in $\hat{X}$. [^3] Recent work in this direction, focusing especially on the case in which $X$ is a (compact) Polish space, can be found in [26]; many of the basic results we present here are implicit in that work. In the forward direction, fix a function $f\in C(S,X)$, and define a $\mathbb{B}$-name $\dot{F}_{f}$ for a subset of $\tau_{c}$ as follows. For every $D\in\tau_{c}$ and every $b\in\mathbb{B}$, put $(\check{D},b)$ in $\dot{F}_{f}$ if and only if $N_{b}\subseteq f^{-1}[D]$.
Lemma 8.2.
$\dot{F}_{f}$ is a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$.
Proof.
The fact that $\dot{F}_{f}$ is forced to be a filter is immediate from the definition. To show that it is forced to be maximal, fix an arbitrary $C\in\tau_{c}$. It suffices to show that the following set is dense in $\mathbb{B}$:
$E_{C}:=\{b\in\mathbb{B}\mid N_{b}\subseteq f^{-1}[C]\text{ {or} }\exists D\in\tau_{c}[D\cap C=\emptyset\wedge N_{b}\subseteq f^{-1}[D]]\}.$
To this end, fix an arbitrary $b_{0}\in\mathbb{B}$, and assume that $N_{b_{0}}\not\subseteq f^{-1}[C]$. Fix $\mathcal{U}\in N_{b_{0}}$ such that $f(\mathcal{U})\in X\setminus C$. Since $X$ is compact, it is regular, so we can find an open set $O\in\tau$ such that $\mathcal{U}\in O\subseteq\mathrm{cl}(O)\subseteq X\setminus C$. By the continuity of $f$, we can find $b\leq b_{0}$ such that $\mathcal{U}\in N_{b}$ and $N_{b}\subseteq f^{-1}[O]$. Then $b\in E_{C}$, as desired. ∎
In the other direction, fix a $\mathbb{B}$-name $\dot{F}$ for a maximal filter on $\tau_{c}$. For each $D\in\tau_{c}$, let $b_{D}:=\llbracket D\in\dot{F}\rrbracket_{\mathbb{B}}$. Given $\mathcal{U}\in S$, let $\dot{F}_{\mathcal{U}}:=\{D\in\tau_{c}\mid b_{D}\in\mathcal{U}\}$.
Lemma 8.3.
For all $\mathcal{U}\in S$, $\bigcap\dot{F}_{\mathcal{U}}$ contains exactly one point.
Proof.
Fix $\mathcal{U}\in S$. We first show that $\dot{F}_{\mathcal{U}}$ is a filter. To this end, fix $D_{0},D_{1}\in\dot{F}_{\mathcal{U}}$. Since $\dot{F}$ is forced to be a filter, we have $\llbracket D_{0}\cap D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}=\llbracket D_{0}\in\dot{F}\rrbracket_{\mathbb{B}}\wedge\llbracket D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}$. Then, since $\mathcal{U}$ is a filter and $\llbracket D_{i}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$ for $i<2$, it follows that $\llbracket D_{0}\cap D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$, so $D_{0}\cap D_{1}\in\dot{F}_{\mathcal{U}}$.
Thus, $\dot{F}_{\mathcal{U}}$ has the finite intersection property; since $X$ is compact, it follows that $\bigcap\dot{F}_{\mathcal{U}}$ is nonempty. Now suppose for sake of contradiction that $x_{0}$ and $x_{1}$ are distinct elements of $\dot{F}_{\mathcal{U}}$. Since $X$ is Hausdorff, we can find disjoint open neighborhoods $U_{0}$ and $U_{1}$ of $x_{0}$ and $x_{1}$, respectively. For $i<2$, let $D_{i}:=X\setminus U_{i}$. Since $D_{0}\cup D_{1}=X$ and $\dot{F}$ is forced to be a maximal filter on $\tau_{c}\setminus\{\emptyset\}$, a routine argument shows that $\llbracket D_{0}\in\dot{F}\rrbracket_{\mathbb{B}}\vee\llbracket D_{1}\in\dot{F}\rrbracket_{\mathbb{B}}=1_{\mathbb{B}}$. Therefore, since $\mathcal{U}$ is an ultrafilter, there must be $i<2$ such that $\llbracket D_{i}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$, and hence $D_{i}\in\dot{F}_{\mathcal{U}}$. But then $x_{1-i}\notin D_{i}$ and $\bigcap\dot{F}_{\mathcal{U}}\subseteq D_{i}$, so $x_{1-i}\notin\bigcap\dot{F}_{\mathcal{U}}$, which is a contradiction. It follows that $\bigcap\dot{F}_{\mathcal{U}}$ contains precisely one point. ∎
Using the claim, define a function $f_{\dot{F}}:S\rightarrow X$ by letting, for all $\mathcal{U}\in S$, $f_{\dot{F}}(\mathcal{U})$ be the unique $x\in\bigcap\dot{F}_{\mathcal{U}}$.
Lemma 8.4.
$f_{\dot{F}}$ is continuous.
Proof.
Fix $U\in\tau$ and $\mathcal{U}\in f_{\dot{F}}^{-1}[U]$. For every $y\in X\setminus U$, we can find $D_{y}\in\tau_{c}$ such that $y\notin D_{y}$ and $b_{D_{y}}\in\mathcal{U}$. Then $\{X\setminus D_{y}\mid y\in X\setminus U\}$ is an open cover of $X\setminus U$, so we can find a finite subset $a\subseteq X\setminus U$ such that $D:=\bigcap\{D_{y}\mid y\in a\}\subseteq U$. Then $b_{D}=\bigwedge\{b_{d_{y}}\mid y\in a\}\in\mathcal{U}$ and $N_{b_{D}}\subseteq f_{\dot{F}}^{-1}[U]$. ∎
It is routine to verify that this is a reflexive duality: Given $g\in C(S,X)$, we have $f_{\dot{F}_{g}}=g$ and, given a $\mathbb{B}$-name $\dot{E}$ for a maximal filter on $\tau_{c}$, we have $\Vdash_{\mathbb{B}}\dot{E}=\dot{F}_{f_{\dot{E}}}$. Also, this story can clearly be relativized below a given condition $b\in\mathbb{B}$: functions in $C(N_{b},X)$ correspond to $\mathbb{B}$-names $\dot{F}$ such that $b\Vdash_{\mathbb{B}}``\dot{F}\text{ is a maximal filter on }\tau_{c}"$.
Now suppose that $X$ is an arbitrary Hausdorff space, not necessarily compact. Now there may or may not be a well-behaved interpretation $\hat{X}$ of $X$ in the extension by $\mathbb{B}$. However, there is still a natural extension of $X$ in $V^{\mathbb{B}}$ that will correspond to a quotient of the space $C(S,X)^{V}$ in $V^{\mathbb{B}}$. We will denote this extension by $\hat{X}_{\mathcal{K}}$; it can be thought of as the extension of $X$ spanned by ground model compact sets. We now give a definition of this space.
Let $\mathcal{K}(X)$ denote the collection of compact subsets of $X$ (in $V$). Then $(\mathcal{K}(X),\subseteq)$ is a directed partial order. For each $K\in\mathcal{K}(X)$, let $\tau_{K}$ denote the subspace topology on $K$. For each $K\in\mathcal{K}(X)$, there is a canonical interpretation $(\hat{K},\hat{\tau}_{K})$ of $(K,\tau_{K})$ in $V^{\mathbb{B}}$, as above. Namely, $\hat{K}$ is the collection of all maximal filters on $(\tau_{K})_{c}$. Moreover, if $K\subseteq K^{\prime}$ are both in $\mathcal{K}(X)$, then the inclusion map $\iota_{KK^{\prime}}:K\rightarrow K^{\prime}$ induces an inclusion map $\hat{\iota}_{KK^{\prime}}:\hat{K}\rightarrow\hat{K^{\prime}}$. We now define $(X_{\mathcal{K}},\tau_{\mathcal{K}})$ in $V^{\mathbb{B}}$.
First, we let the underlying set, $X_{\mathcal{K}}$, be the direct limit (i.e., colimit) of the system $\langle\hat{K},\hat{\iota}_{KK^{\prime}}\mid K,K^{\prime}\in\mathcal{K}(X),\ K\subseteq K^{\prime}\rangle$. Formally, $X_{\mathcal{K}}$ consists of equivalence classes of the form $[(x,K)]$, where $K\in\mathcal{K}(X)$ and $x\in\hat{K}$. Given an open set $U\in\tau$ and $K\in\mathcal{K}(X)$, let $\hat{U}_{K}$ denote its interpretation in $\hat{\tau}_{K}$. We then define $\hat{U}_{\mathcal{K}}\subseteq\hat{X}_{\mathcal{K}}$ as follows: given $K\in\mathcal{K}(X)$ and $x\in\hat{K}$, put $[(x,K)]\in\hat{U}_{\mathcal{K}}$ if and only if $x\in\hat{U}_{K}$. It is routine to check that this is well-defined. Now let $\hat{\tau}_{\mathcal{K}}$ be the topology generated by $\{\hat{U}_{\mathcal{K}}\mid U\in\tau\}$.
We take a slight detour to note that this is not the only, nor arguably the most natural, topology we could have placed on $\hat{X}_{\mathcal{K}}$. We alternatively could have taken the direct limit topology, which we will denote by $\hat{\tau}^{\lim}_{\mathcal{K}}$, i.e., the finest topology that makes all of the limit maps $\hat{\iota}_{K}:\hat{K}\rightarrow\hat{X}_{\mathcal{K}}$ continuous. We always have $\hat{\tau}_{\mathcal{K}}\subseteq\hat{\tau}^{\lim}_{\mathcal{K}}$, and if $X$ is locally compact, then the two topologies coincide (moreover, if $X$ is locally compact, then $\hat{X}_{\mathcal{K}}$ is precisely the interpretation of $X$ in $V^{\mathbb{B}}$ in the sense of [29]). However, in general the direct limit topology can be strictly finer, even for such relatively nice spaces as ${{}^{\omega}}\omega$, as the following example shows.
Proposition 8.5.
Suppose that $X={{}^{\omega}}\omega$ and $\mathbb{B}$ adds a dominating real. Then $\hat{\tau}^{\lim}_{\mathcal{K}}\supsetneq\hat{\tau}_{\mathcal{K}}$ in $V^{\mathbb{B}}$.
Proof.
Let $W$ denote the extension by $\mathbb{B}$. In $V$, for every $x\in{{}^{\omega}}\omega$ let $K_{x}:=\{y\in{{}^{\omega}\omega}\mid y<x\}$. Then $\{K_{x}\mid x\in{{}^{\omega}}\omega\}$ is a $\subseteq$-cofinal subset of $\mathcal{K}(X)$. Moreover, it is routine to show that, for every $x\in{{}^{\omega}}\omega$, we have $\hat{K}_{x}=(\prod_{n<\omega}x(n))^{W}$. Therefore, the underlying set of $\hat{X}_{\mathcal{K}}$ can be identified with $\{y\in({{}^{\omega}}\omega)^{W}\mid\exists x\in({{}^{\omega}}\omega)^{V}~{}y<x\}$.
Let $d\in({{}^{\omega}}\omega)^{W}$ dominate every real in $V$, and let
$O=\{y\in\hat{X}_{\mathcal{K}}\mid\forall n<\omega~{}y(n)\neq d(n)\}.$
Claim 8.6.
$O\in\hat{\tau}^{\lim}_{\mathcal{K}}$.
Proof.
It suffices to show that, for every $x\in({{}^{\omega}}\omega)^{V}$, $O\cap\hat{K}_{x}\in\hat{\tau}_{K_{x}}$. Fix such an $x$, and let $A:=\{n<\omega\mid d(n)<x(n)\}$. By choice of $d$, $A$ is finite. Then, in $V$, the set $U=\{y\in{{}^{\omega}}\omega\mid\forall n\in A~{}y(n)\neq d(n)\}$ is in $\tau$, and we have $O\cap\hat{K}_{x}=\hat{U}_{K_{x}}$. ∎
Claim 8.7.
There is no $U\in\tau$ such that $\hat{U}_{\mathcal{K}}\subseteq O$.
Proof.
Fix $U\in\tau$. By shrinking $U$ if necessary, we may assume that $U$ is a basic open set, i.e., there is a finite set $A\subseteq\omega$ and a function $\sigma:A\rightarrow\omega$ such that $U=\{y\in{{}^{\omega}}\omega\mid y\restriction A=\sigma\}$. Choose $n^{*}\in\omega\setminus A$ and define $y\in{{}^{\omega}}\omega$ by
$y(n)=\begin{cases}\sigma(n)&\text{if }n\in A\\ d(n^{*})&\text{if }n=n^{*}\\ 0&\text{otherwise.}\end{cases}$
Then, in $W$, we have $y\in O\setminus\hat{U}_{\mathcal{K}}$, so $\hat{U}_{\mathcal{K}}\not\subseteq O$. ∎
Since $\{\hat{U}_{\mathcal{K}}\mid U\in\tau\}$ is a base for $\hat{\tau}_{\mathcal{K}}$, this completes the proof. ∎
We now show that $\hat{X}_{\mathcal{K}}$ is homeomorphic to a natural quotient of $C(S,X)^{V}$ in the extension by $\mathbb{B}$. Recall that we endow $C(S,X)^{V}$ with the compact-open topology. Now suppose that $G$ is a generic ultrafilter on $\mathbb{B}$, and, in $V[G]$, define an equivalence relation $\sim_{G}$ on $C(S,X)^{V}$ by letting $g\sim_{G}h$ if and only if there is $b\in G$ such that $g\restriction N_{b}=h\restriction N_{b}$. Let $C(S,X)^{V}/G$ denote the quotient space with respect to this equivalence relation. For each $U\in\tau$, define a set $N_{U}\subseteq C(S,X)^{V}/G$ as follows: for all $g\in C(S,X)^{V}$, put $[g]\in N_{U}$ if and only if there is $b\in G$ such that $g[N_{b}]\subseteq U$; note that this is well-defined.
Proposition 8.8.
$\{N_{U}\mid U\in\tau\}$ is a base for $C(S,X)^{V}/G$.
Proof.
It is immediate from the definition that each $N_{U}$ is open. Now let $\hat{O}\subseteq C(S,X)^{V}/G$ be open, and fix $g\in C(S,X)^{V}$ such that $[g]\in\hat{O}$. Let $O=\{h\in C(S,X)^{V}\mid[h]\in\hat{O}\}$. We can then find a compact $K\subseteq S$ and an open set $U\in\tau$ such that $g[K]\subseteq U$ and $\{h\in C(S,X)^{V}\mid h[K]\subseteq U\}\subseteq O$. Since $K$ is compact and $g$ is continuous, we can find a finite collection $b_{0},\ldots,b_{n-1}$ from $\mathbb{B}$ such that $K\subseteq\bigcup\{N_{b_{i}}\mid i<n\}$ and, for all $i<n$, we have $g[N_{b_{i}}]\subseteq U$. Letting $b:=\bigvee\{b_{i}\mid i<n\}$, we get $K\subseteq N_{b}$ and $g[N_{b}]\subseteq U$.
Suppose first that $b\in G$. In this case, we claim that $N_{U}\subseteq\hat{O}$. To see this, fix $h\in C(S,X)^{V}$ such that $[h]\in N_{U}$. We can therefore find $c\in G$ such that $h[N_{c}]\subseteq U$. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{c}=h\restriction N_{c}$ and letting $h^{\prime}\restriction(S\setminus N_{c})$ be constant, taking an arbitrary value in $U$. Then $h^{\prime}[K]\subseteq U$, so $[h^{\prime}]\in\hat{O}$. Moreover, since $c\in G$, we have $h^{\prime}\sim_{G}h$, so $[h^{\prime}]=[h]$. Thus, $N_{U}\subseteq\hat{O}$.
Suppose next that $b\notin G$. In this case, we claim that $\hat{O}$ is actually all of $C(S,X)^{V}/G$. To see this, let $h\in C(S,X)^{V}$ be arbitrary. Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{b}=g\restriction N_{b}$ and $h^{\prime}\restriction(S\setminus N_{b})=h\restriction(S\setminus N_{b})$. Then $h^{\prime}[N_{b}]=g[N_{b}]\subseteq U$, so $[h^{\prime}]\in\hat{O}$. Since $b\notin G$, we have $[h^{\prime}]=[h]$, so $[h]\in\hat{O}$. ∎
Now, in $V[G]$, define a map $k:C(S,X)^{V}/G\rightarrow\hat{X}_{\mathcal{K}}$ as follows. Given $g\in C(S,X)^{V}$, let $K:=g[S]$. Then $K\in\mathcal{K}(X)$, so $g$ corresponds in $V$ to a $\mathbb{B}$-name $\dot{F}_{g}$ for an element of $\hat{K}$. Let $F_{g}$ be the evaluation of this name in $V[G]$, and set $k([g])=[(F_{g},K)]$.
Proposition 8.9.
$k$ is a bijection.
Proof.
To show that $k$ is injective, fix two functions $g,h\in C(S,X)^{V}$. Familiar arguments yield the fact that, in $V$, the set
$\{b\in\mathbb{B}\mid g\restriction N_{b}=h\restriction N_{b}\text{ {or} }g[N_{b}]\cap h[N_{b}]=\emptyset\}$
is dense in $\mathbb{B}$. The genericity of $G$ then yields the injectivity of $k$. To show that $k$ is surjective, fix an element $[(F,K)]\in\hat{X}_{\mathcal{K}}$, and let $\dot{F}$ be a $\mathbb{B}$-name for $F$ that is forced by $1_{\mathbb{B}}$ to be a maximal filter on $(\tau_{K})_{c}$. Then $f_{\dot{F}}$, as defined at the beginning of this section, is in $C(S,X)^{V}$, and it is routine to verify that $k([f_{\dot{F}}])=[(F,K)]$. ∎
The continuity of $k$ and $k^{-1}$ will now follow immediately from the following proposition.
Proposition 8.10.
For all $U\in\tau$, we have $k[N_{U}]=\hat{U}_{K}$.
Proof.
Suppose that $g\in C(S,X)^{V}$ and $[g]\in N_{U}$. Let $K:=g[S]$, and fix $b\in G$ such that $g[N_{b}]\subseteq U$. Then, by construction, we have $(K\setminus U)\notin F_{g}$, so $k([g])=[(F_{g},K)]\in\hat{U}_{\mathcal{K}}$. Conversely, suppose that $g\in C(S,X)^{V}$ and $[g]\notin N_{U}$. Again, let $K:=g[S]$. By familiar arguments, the following set is dense in $\mathbb{B}$ (in $V$):
$\{b\in\mathbb{B}\mid g[N_{b}]\subseteq U\text{ {or} }g[N_{b}]\subseteq K\setminus U\}.$
Since $[g]\notin N_{U}$, the genericity of $G$ implies that $K\setminus U\in F_{g}$, so $k([g])=[(F_{g},K))]\notin\hat{U}_{\mathcal{K}}$. ∎
Altogether, we have established the following theorem.
Theorem 8.11.
Suppose that $X$ is a Hausdorff space, $\mathbb{B}$ is a complete Boolean algebra, and $G\subseteq\mathbb{B}$ is a generic ultrafilter. Then, in $V[G]$, we have
$C(S(\mathbb{B}),X)^{V}/G\cong\hat{X}_{\mathcal{K}}.$
We end this section with a few remarks. First, as was mentioned above, if $X$ is a locally compact Hausdorff space, then $\hat{X}_{\mathcal{K}}$ is the canonical interpretation of $X$ in $V^{\mathbb{B}}$ in the sense of [29]. Recall also that the category $\mathsf{ED}$ is precisely the category of Stone spaces of complete Boolean algebras; recall as well that $\underline{X}(S)$ equals, by definition, $C(S,X)$ for any $S\in\mathsf{ED}$. Therefore, the results of this section show that one way to think of the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, for all such spaces $\underline{X}$, the condensed set $\underline{X}$ is precisely an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
Second, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products. Namely, if $X_{i}$ is a Hausdorff space for all $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$. This is because the collection
$\{\prod_{i\in I}K_{i}\mid\forall i\in I\ K_{i}\in\mathcal{K}(X_{i})\}$
is $\subseteq$-cofinal in $\mathcal{K}(\prod_{i\in I}X_{i})$.
Third, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact. In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11. If $X$ is compact, then, for every $\mathbb{B}$-name $\dot{x}$ for an element of $\hat{X}_{\mathcal{K}}(=\hat{X})$, there is a $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. In particular, in this case we can take $\dot{x}$ to be a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$, then then let $g=f_{\dot{x}}$, as defined at the outset of this section. However, in general this may not be the case, and there may be $\mathbb{B}$-names $\dot{x}$ for elements of $\hat{X}_{\mathcal{K}}$ for which there is no $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. For such $\dot{x}$, one may need to work below some condition $b\in\mathbb{B}$ that forces $\dot{x}$ to lie inside the interpretation of some particular $K\in\mathcal{K}(X)$; once one does so, one may then find $g\in C(S,X)$ such that $b\Vdash\dot{k}([g])=\dot{x}$. This issue will lead to some slight complications and asymmetries in the next section.
We end this section by noting that these observations are of wider significance than might initially be suspected, by reason of the variety of expressions within condensed mathematics in which sets of the form $C(S,X)$, sometimes covertly, partake. The reader is referred to Lemma 11.1 below, and its argument, for a representative instance.
9. Compact interpretations and continuous functions
Fix for now a complete Boolean algebra $\mathbb{B}$, and let $S=S(\mathbb{B})$. In this section, we establish an equivalence between the settings of $C(S,X)$ on the one hand and $\mathbb{B}$-names for elements of $\hat{X}_{\mathcal{K}}$ on the other, with respect to interpretations of certain sentences. Fix for now a positive integer $n$ and Hausdorff spaces $\langle X_{i}\mid i<n\rangle=\vec{X}$. We will specify a collection $\mathcal{L}(\vec{X})$ of formulas. We will do this by recursively defining $\mathcal{L}_{j}(\vec{X})$ for $j\leq n$ and then letting $\mathcal{L}(\vec{X}):=\bigcup_{j\leq n}\mathcal{L}_{j}(\vec{X})$. For each $j\leq n$ and $\varphi\in\mathcal{L}_{j}(\vec{X})$, the free variables in $\varphi$ will be precisely $\{v_{i}\mid j\leq i<n\}$.
First, let $\mathcal{L}_{0}(\vec{X})$ consist of all formulas of the form $(v_{0},v_{1},\ldots,v_{n-1})\in U$, where $U$ is an open subset of $\prod_{i<n}X_{i}$. Now suppose that $j<n$ and we have specified $\mathcal{L}_{j}(\vec{X})$. Then $\mathcal{L}_{j+1}(\vec{X})$ consists precisely of the formulas that can be formed in one of the following ways. First, fix a formula $\varphi(v_{j},v_{j+1},\ldots,v_{n-1})\in\mathcal{L}_{j}(\vec{X})$. Then
  1. (1)
    the formula $\forall x\varphi(x,v_{j+1},\ldots,v_{n-1})$ is in $\mathcal{L}_{j+1}(\vec{X})$;
  2. (2)
    the formula $\exists x\varphi(x,v_{j+1},\ldots,v_{n-1})$ is in $\mathcal{L}_{j+1}(\vec{X})$;
  3. (3)
    for every compact $K\subseteq X_{j}$, the formula $\exists x\in K\varphi(x,v_{j+1},\ldots,v_{n-1})$ is in $\mathcal{L}_{j+1}(\vec{X})$.
We call a formula $\varphi\in\mathcal{L}(\vec{X})$ bounded if all of its existential quantifiers are bounded by compact sets (i.e., they arise through option (3) above rather than option (2)). In their most basic setting, that of $\prod_{i<n}X_{i}$, the free variables $v_{i}$ occurring in formulas in $\mathcal{L}(\vec{X})$ should be understood as standing for points in $X_{i}$, and the truth value of sentences arising from the substitution of these points for the free variables is evaluated in the natural way. But we also want to interpret these formulas in other contexts, namely in the context of $C(S,\prod_{i<n}X_{i})$ and of $V^{\mathbb{B}}$.
Let us first deal with $C(S,\prod_{i<n}X_{i})$, where the free variables $v_{i}$ stand for elements of $C(S,X_{i})$. We describe now what it means for $C(S,\prod_{i<n}X_{i})$ to satisfy a sentence formed by substituting appropriate continuous functions for free variables in formulas in $\mathcal{L}(\vec{X})$. By induction on $j\leq n$, we will deal with $\mathcal{L}_{j}(\vec{X})$, in fact specifying what it means for a sentence to be satisfied below a condition $b\in\mathbb{B}$.
Suppose that $\varphi\in\mathcal{L}_{0}(\vec{X})$, $b\in\mathbb{B}$, and, for all $i<n$, $f_{i}\in C(S,X_{i})$. Then $\varphi$ is of the form $(v_{0},\ldots,v_{n-1})\in U$ for some open set $U\subseteq\prod_{i<n}X_{i}$. We say that $C(S,\prod_{i<n}X_{i})$ satisfies $\varphi(f_{0},\ldots,f_{n-1})$ below $b$, written
$C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{0},\ldots,f_{n-1}),$
if and only if the set $\{s\in N_{b}\mid(f_{0}(s),\ldots,f_{n-1}(s))\in U\}$ is dense in $N_{b}$. If $b=1_{\mathbb{B}}$, then it is omitted from the terminology and notation.
Now suppose that $j<n$, $\varphi\in\mathcal{L}_{j+1}$, $b\in\mathbb{B}$, and, for all $j<i<n$, $f_{i}\in C(S,X_{i})$.
  • If $\varphi$ is of the form $\forall x\psi(x,v_{j+1},\ldots,v_{n-1})$, then
    $C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{j+1},\ldots,f_{n-1})$
    if and only if, for all $g\in C(N_{b},X_{j})$, we have
    $C(S,\prod_{i<n}X_{i})\models_{b}\psi(g,f_{j+1},\ldots,f_{n-1}).$
  • If $\varphi$ is of the form $\exists x\psi(x,v_{j+1},\ldots,v_{n-1})$, then
    $C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{j+1},\ldots,f_{n-1})$
    if and only if there exists $g\in C(N_{b},X_{j})$ such that
    $C(S,\prod_{i<n}X_{i})\models_{b}\psi(g,f_{j+1},\ldots,f_{n-1}).$
  • If $K\subseteq X_{j}$ is compact and $\varphi$ is of the form $\exists x\in K\psi(x,v_{j+1},\ldots,v_{n-1})$, then
    $C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{j+1},\ldots,f_{n-1})$
    if and only if there exists $g\in C(N_{b},K)$ such that
    $C(S,\prod_{i<n}X_{i})\models_{b}\psi(g,f_{j+1},\ldots,f_{n-1}).$
In $V^{\mathbb{B}}$, the free variables $v_{i}$ stand for $\mathbb{B}$-names for elements of $(\hat{X}_{i})_{\mathcal{K}}$. We now define satisfaction in this context. Suppose that $\varphi\in\mathcal{L}_{0}(\vec{X})$ is of the form $(v_{0},\ldots,v_{n-1})\in U$, $b\in\mathbb{B}$, and, for all $i<n$, $\dot{x}_{i}$ is a $\mathbb{B}$-name for an element of $(\hat{X}_{i})_{\mathcal{K}}$. Then $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{0},\ldots,\dot{x}_{n-1})$ if and only if $b\Vdash_{\mathbb{B}}(\dot{x}_{0},\ldots,\dot{x}_{n-1})\in\hat{U}_{\mathcal{K}}$.
Now suppose that $j<n$, $\varphi\in\mathcal{L}_{j+1}$, $b\in\mathbb{B}$, and, for all $j<i<n$, $\dot{x}_{i}$ is a $\mathbb{B}$-name for an element of $(\hat{X}_{i})_{\mathcal{K}}$.
  • If $\varphi$ is of the form $\forall x\psi(x,v_{j+1},\ldots,v_{n-1})$, then $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$ if and only if, for every $\mathbb{B}$-name $\dot{y}$ for an element of $(\hat{X}_{j})_{\mathcal{K}}$, we have $V^{\mathbb{B}}\models_{b}\psi(\dot{y},\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$.
  • If $\varphi$ is of the form $\exists x\psi(x,v_{j+1},\ldots,v_{n-1})$, then $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$ if and only if there exists a $\mathbb{B}$-name $\dot{y}$ for an element of $(\hat{X}_{j})_{\mathcal{K}}$ such that $V^{\mathbb{B}}\models_{b}\psi(\dot{y},\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$.
  • If $K\subseteq X_{j}$ is compact and $\varphi$ is of the form $\exists x\in K\psi(x,v_{j+1},\ldots,v_{n-1})$, then $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$ if and only if there exists a $\mathbb{B}$-name $\dot{y}$ such that $b\Vdash_{\mathbb{B}}\dot{y}\in\hat{K}$ and $V^{\mathbb{B}}\models_{b}\psi(\dot{y},\dot{x}_{j+1},\ldots,\dot{x}_{n-1})$.
Note that there is some monotonicity in these definitions. Namely, if $c\leq_{\mathbb{B}}b$ and $C(S,\prod_{i<n}X_{i})\models_{b}\varphi$, then $C(S,\prod_{i<n}X_{i})\models_{c}\varphi$, and similarly for $V^{\mathbb{B}}$.
Given $i<n$ and $f\in C(S,X_{i})$, let $\dot{x}_{f}$ be a $\mathbb{B}$-name for $\dot{k}([f])$, where $\dot{k}$ is a name for the homeomorphism constructed to witness Theorem 8.11. Concretely, if $K=f[S]$, then $\dot{x}_{f}$ is a $\mathbb{B}$-name for $[(\dot{F}_{f},K)]\in\hat{X}_{\mathcal{K}}$.
Theorem 9.1.
Suppose that $j\leq n$, $\varphi\in\mathcal{L}_{j}(\vec{X})$, $b\in\mathbb{B}$ and, for all $j\leq i<n$, we have $f_{i}\in C(S,X_{i})$.
  1. (1)
    If $C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{j},\ldots,f_{n-1})$, then $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$.
  2. (2)
    If $\varphi$ is bounded and $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, then $C(S,\prod_{i<n}X_{i})\models_{b}\varphi(f_{j},\ldots,f_{n-1})$.
Proof.
The proof is by induction on $j$. Suppose first that $j=0$. For notational convenience, we can assume here that $n=1$ by considering the single space $X=\prod_{i<n}X_{i}$ (recalling that $\widehat{(\prod_{i<n}X_{i})}_{\mathcal{K}}=\prod_{i<n}(\hat{X}_{i})_{\mathcal{K}}$). Then there is an open set $U\subseteq X$ such that $\varphi$ is the formula $v_{1}\in U$, and we are given a function $f\in C(S,X)$. Let $K:=\mathrm{im}(f)$, so $K$ is a compact subset of $X$, and $\dot{x}_{f}$ is forced to be of the form $[(\dot{F}_{f},K)]$.
To verify (1), suppose that $C(S,X)\models_{b}\varphi(f)$, and hence $\{s\in N_{b}\mid f(s)\in U\}$ is dense in $N_{b}$. Suppose for sake of contradiction that $b\not\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$. Then, by the definition of $\dot{F}_{f}$, there is $c\leq b$ such that $(\check{K}\setminus\check{U},c)\in\dot{F}_{f}$, and therefore, again by the definition of $\dot{F}_{f}$, we have $N_{c}\subseteq f^{-1}[K\setminus U]$, which is a contradiction.
For (2), suppose that $b\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$, i.e., $b\Vdash_{\mathbb{B}}(K\setminus U)\notin\dot{F}_{f}$. Suppose for sake of contradiction that $\{s\in N_{b}\mid f(s)\in U\}$ is not dense in $N_{b}$. Then there is $c\leq b$ such that $N_{c}\subseteq f^{-1}[K\setminus U]$. But then $c\Vdash_{\mathbb{B}}(K\setminus U)\in\dot{F}_{f}$, which is a contradiction.
Now suppose that $j=j^{\prime}+1$ for some $j^{\prime}<n$. To verify (1), suppose that $C(S,X)\models_{b}\varphi(f_{j},\ldots f_{n-1})$ but, for sake of contradiction, assume that $V^{\mathbb{B}}\not\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. Suppose first that $\varphi$ is of the form $\forall x\psi$ for some $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. By assumption, there is a $\mathbb{B}$-name $\dot{y}$ for an element of $(\hat{X}_{j^{\prime}})_{\mathcal{K}}$ such that $V^{\mathbb{B}}\not\models_{b}\psi(\dot{y},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. We can then find a condition $c\leq_{\mathbb{B}}b$ and a compact set $K\subseteq X_{j}$ such that $V^{\mathbb{B}}\not\models_{c}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$ and $c\Vdash_{\mathbb{B}}\dot{y}\in\hat{K}$. We can therefore find a function $g\in C(S,K)$ such that $c\Vdash_{\mathbb{B}}\dot{y}=[(\dot{F}_{g},\hat{K})]=\dot{x}_{g}$. Then $V^{\mathbb{B}}\not\models_{c}\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so by the induction hypothesis we have $C(S,X)\not\models_{c}\psi(g,f_{j},\ldots,f_{n-1})$, contradicting the assumption that $C(S,X)\models_{b}\varphi(f_{j},\ldots,f_{n-1})$.
Suppose next that $\varphi$ is of the form $\exists x\psi$ for some $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. Then we can fix $g\in C(S,X_{j^{\prime}})$ such that $C(S,\vec{X})\models_{b}\psi(g,f_{j},\ldots,f_{n-1})$. By the induction hypothesis, it follows that $V^{\mathbb{B}}\models\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so $V^{\mathbb{B}}\models\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. The case in which $\varphi$ is of the form $\exists x\in K\psi$ is the same.
To verify (2), suppose that $\varphi$ is bounded and $V^{\mathbb{B}}\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. Assume first that $\varphi$ is of the form $\forall x\psi$ for some (bounded) $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. Then, for all $g\in C(S,X_{j^{\prime}})$, we have $V^{\mathbb{B}}\models_{b}\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so, by the induction hypothesis, we have $C(S,\vec{X})\models_{b}\psi(g,f_{j},\ldots,f_{n-1})$. Since $g$ was arbitrary, we have $C(S,\vec{X})\models_{b}\varphi(f_{j},\ldots,f_{n-1})$.
Finally, assume that $\varphi$ is of the form $\exists x\in K\psi$ for some compact $K\subseteq X_{j^{\prime}}$ and some (bounded) $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$. Then we can find a $\mathbb{B}$-name $\dot{y}$ for an element of $\hat{K}$ such that $V^{\mathbb{B}}\models_{b}\psi(\dot{y},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$. We can then find a function $g\in C(S,K)$ such that $b\Vdash_{\mathbb{B}}\dot{y}=\dot{x}_{g}$. Then $V^{\mathbb{B}}\models_{b}\psi(\dot{x}_{g},\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$, so, by the induction hypothesis, we have $C(S,\vec{X})\models_{b}\psi(g,f_{j},\ldots,f_{n-1})$. Since $g\in C(S,K)$, it follows that $C(S,\vec{X})\models_{b}\varphi(f_{j},\ldots,f_{n-1})$, as desired. ∎
We now finally return to Whitehead’s problem, where the work of this and the previous section allows us to establish the following general result.
Theorem 9.2.
Suppose that $A$ is an abelian group and $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$, and let $B_{K}$ be a basis for $K$, and let $e:B_{K}\rightarrow\omega$ be a function. Suppose that $\mathbb{B}$ is a complete Boolean algebra that forces the existence of a homomorphism $\tau:K\rightarrow\mathbb{Z}$ such that
  • in $V^{\mathbb{B}}$, $\tau$ does not lift to a homomorphism $\nu:F\rightarrow\mathbb{Z}$;
  • for all $t\in B_{K}$, we have $|\tau(t)|\leq e(t)$.
Then, in $V$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$.
Proof.
Let $B_{F}$ be a basis for $F$, and note that elements of $\mathrm{Hom}(K,\mathbb{Z})$ and $\mathrm{Hom}(F,\mathbb{Z})$ can be identified with elements of ${{}^{B_{K}}}\mathbb{Z}$ and ${{}^{B_{F}}}\mathbb{Z}$, respectively, and the restriction map from $\mathrm{Hom}(F,\mathbb{Z})$ to $\mathrm{Hom}(K,\mathbb{Z})$ induces a continuous map $\pi:{{}^{B_{F}}}\mathbb{Z}\rightarrow{{}^{B_{K}}}\mathbb{Z}$. Note that the assertion that $A$ is not Whitehead is precisely the assertion that $\pi$ is not surjective, i.e., it corresponds to the sentence $\exists x\in{{}^{B_{K}}}\mathbb{Z}~{}\forall y\in{{}^{B_{F}}}\mathbb{Z}~{}\pi(y)\neq x$. Noting that the set $U:=\{(y,x)\in{{}^{B_{F}}}\mathbb{Z}\times{{}^{B_{K}}}\mathbb{Z}\mid\pi(y)\neq x\}$ is an open subset of ${{}^{B_{F}}}\mathbb{Z}\times{{}^{B_{K}}}\mathbb{Z}$, we see that this sentence is an element of $\mathcal{L}(({{}^{B_{F}}}\mathbb{Z},{{}^{B_{K}}}\mathbb{Z}))$.
By assumption, $A$ is not Whitehead in $V^{\mathbb{B}}$, and in fact more is true. Let $\hat{\pi}$ be the extension of $\pi$ to $V^{\mathbb{B}}$ induced by the restriction map from $\mathrm{Hom}(F,\mathbb{Z})$ to $\mathrm{Hom}(K,\mathbb{Z})$. Then $\mathbb{B}$ forces that there is an element of $\prod_{t\in B_{K}}[-e(t),e(t)]$ that is not in the range of $\pi$. Let $X_{K}$ and $X_{F}$ denote ${{}^{B_{K}}}\mathbb{Z}$ and ${{}^{B_{F}}}\mathbb{Z}$, respectively, and note that $K^{*}:=\prod_{t\in B_{K}}[-e(t),e(t)]$ is a compact subset of $X_{K}$. Let $\varphi$ be the (bounded) sentence $\exists x\in K^{*}\forall y\in X_{F}~{}\pi(y)\neq x$. It is readily verified that the following facts hold in $V^{\mathbb{B}}$.
  • $(\hat{X}_{K})_{\mathcal{K}}$ is (homeomorphic to) the set of all $x\in({{}^{B_{K}}}\mathbb{Z})^{V^{\mathbb{B}}}$ such that there exists $e:B_{K}\rightarrow\omega$ in $V$ such that $|x(t)|\leq e(t)$ for all $t\in B_{K}$. A similar characterization holds for $(\hat{X}_{F})_{\mathcal{K}}$.
  • $\hat{U}_{\mathcal{K}}=\{(y,x)\in(\hat{X}_{K})_{\mathcal{K}}\times(\hat{X}_{F})_{\mathcal{K}}\mid\hat{\pi}(y)\neq x\}$.
By assumption, we then have $V^{\mathbb{B}}\models\varphi$. By Theorem 9.1, it follows that $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi$, and in fact $C(S(\mathbb{B}),X_{F}\times K^{*})\models\varphi$, i.e., there is a continuous function $f:S(\mathbb{B})\rightarrow K^{*}$ such that, for all continuous $g:S(\mathbb{B})\rightarrow X_{F}$, the set of $s\in S(\mathbb{B})$ such that $\pi(g(s))\neq f(s)$ is dense in $S(\mathbb{B})$. But this immediately implies that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$, completing the proof of the theorem. ∎
We now immediately obtain the following corollary, establishing clause (3) of Theorem 1.3.
Corollary 9.3.
Suppose that $A$ is a nonfree abelian group and $\kappa$ is the least cardinality of a nonfree subgroup of $A$. Let $S$ be the Stone space of the Boolean completion of the forcing to add $\kappa$-many Cohen reals. Then $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)\neq\emptyset$.
Proof.
Let $\mathbb{B}$ denote the Boolean completion of the forcing to add $\kappa$-many Cohen reals, let $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ be a free resolution of $A$, and let $B_{K}$ be a basis for $K$. By Theorem 6.3, $\mathbb{B}$ forces the existence of a homomorphism $\tau:K\rightarrow\mathbb{Z}$ such that
  • in $V^{\mathbb{B}}$, $\tau$ does not lift to a homomorphism $\nu:F\rightarrow\mathbb{Z}$; and
  • for all $t\in B_{K}$, $\tau(t)\in\{0,1\}$.
Therefore, Theorem 9.2 implies that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$, as desired. ∎
We end this section with some observations regarding forcing extensions in which Whitehead’s problem has a negative answer. We first note that the converse of Theorem 9.2 is not true in general; namely, if $A$ is an abelian group and $\mathbb{B}$ is a complete Boolean algebra, then it does not follow from $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$ that $A$ is not Whitehead in $V^{\mathbb{B}}$. To see this, first recall that an abelian group $A$ is $\aleph_{1}$-free if all of its countable subgroups are free, and a subgroup $B$ of an $\aleph_{1}$-free abelian group $A$ is said to be $\aleph_{1}$-pure if $A/B$ is $\aleph_{1}$-free. An abelian group $A$ satisfies Chase’s condition if $A$ is $\aleph_{1}$-free and every countable subgroup of $A$ is contained in a countable $\aleph_{1}$-pure subgroup of $A$. In [9], Griffith showed that one can construct in $\sf ZFC$ an abelian group of size $\aleph_{1}$ that is not free but satisfies Chase’s condition. One half of Shelah’s independence result can then be stated as follows.
Theorem 9.4 (Shelah, [21], cf. [5]).
Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition. Then $A$ is Whitehead.
Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free. For concreteness, suppose it is constructed as in [5, Theorem 7.3]. It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free. Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Since $V=L$, $A$ is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.
However, we can recover an informative partial converse to Theorem 9.2, in the following sense. Note that, in the situation described in the paragraph above, forcing with $\mathbb{B}$ necessarily adds a new element $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ that extends $\tau$. However, this new homomorphism $\sigma$ must be quite far from any ground model homomorphism; in particular, it cannot be bounded by any ground model function when restricted to any basis for $F$ that lies in $V$. More generally, we have the following result.
Theorem 9.5.
Suppose that $A$ is a non-Whitehead abelian group, $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ is a free resolution of $A$, and $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ is a homomorphism that does not extend to a homomorphism $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Suppose moreover that $B_{F}$ is a basis for $F$, $e:B_{F}\rightarrow\omega$ is a function, and $\mathbb{B}$ is a complete Boolean algebra. Then, in $V^{\mathbb{B}}$, there is no homomorphism $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ such that
  • $\sigma\restriction K=\tau$; and
  • for all $y\in B_{F}$, $|\sigma(y)|\leq e(y)$.
Proof.
Let $B_{K}$ be a basis for $K$, let $X_{K}:={{}^{B_{K}}}\mathbb{Z}$ and $X_{F}:={{}^{B_{F}}}\mathbb{Z}$, let $\pi:X_{F}\rightarrow X_{K}$ be the continuous map from the proof of Theorem 9.2, and again let $U:=\{(y,x)\in X_{F}\times X_{K}\mid\pi(y)\neq x\}$. Let $\varphi(x)$ be the formula $\forall y\in X_{F}(y,x)\in U$, where $x$ is a free variable. Let $f:S(\mathbb{B})\rightarrow X_{K}$ be the constant function taking value $\tau\restriction B_{K}$. Then, by choice of $\tau$, we have $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi(f)$. By Theorem 9.1, it follows that $V^{\mathbb{B}}\models\varphi(\dot{x}_{f})$. Since $f$ is a constant function, it is readily verified that $\dot{x}_{f}$ is forced to be equal to the value of $f$, i.e., to $\tau\restriction B_{K}$. Therefore, it follows that, in $V^{\mathbb{B}}$, for all $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$, if $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$, then $\sigma\restriction K\neq\tau$. But $\sigma\restriction B_{F}\in(\hat{X}_{F})_{\mathcal{K}}$ is equivalent to the existence of an $e:B_{F}\rightarrow\omega$ in $V$ such that $|\sigma(y)|\leq e(y)$ for all $y\in B_{F}$, so the theorem follows. ∎
10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
In this section, we show that the independence of Whitehead’s problem in the classical setting can persist if we restrict ourselves to $\mathsf{CondAb}_{\kappa}$ for certain cardinal $\kappa$. Both for concreteness and because the category $\mathsf{CondAb}_{\omega_{1}}$ of light condensed sets provides the setting for [4], we focus on the case $\kappa=\omega_{1}$. The arguments in this section will be almost entirely set theoretic and combinatorial. We present a more algebraic path to related results in Section 11.
For a given abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\aleph_{1}}$, then $A$ is Whitehead in the classical sense, since
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{1}}}(\underline{A},\underline{\mathbb{Z}})(\ast)=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}).$
Therefore, one side of the independence of Whitehead’s problem in $\mathsf{CondAb}_{\omega_{1}}$ is clear: if it is the case that, classically, every Whitehead abelian group is free, then it follows that, for every abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\omega_{1}}$, then $A$ is free. In this section, we establish the nontrivial side of this independence. Namely, we prove the consistency of the existence of a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\omega_{1}}$. In fact, we shall prove that this follows from $\sf MA_{\omega_{1}}$, the same hypothesis used by Shelah in his proof of Theorem 1.1.
In what follows, when we write, e.g., that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, it should be implicitly understood that $\Lambda$ is a directed partial order and each $S_{i}$ is a finite (discrete) space. For each $i\in\Lambda$, we get a projection map $\pi_{i}:S\rightarrow S_{i}$. The basic open subsets of $S$ are of the form $\pi_{i}^{-1}\{\bar{s}\}$, where $i\in\Lambda$ and $\bar{s}\in S_{i}$. If we write $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, then this means, moreover, that $\Lambda$ is countable.
Proposition 10.1.
Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$ and $\{O_{k}\mid k<n\}$ is a finite partition of $S$ into nonempty clopen sets. Then there is an $i\in\Lambda$ and a function $f:S_{i}\rightarrow n$ such that, for all $k<n$, we have $O_{k}=\{s\in S\mid f(\pi_{i}(s))=k\}$.
Proof.
This follows immediately from the fact that each $O_{k}$ is compact and open and can thus be written as a finite union of basic open subsets of $S$, together with the directedness of $\Lambda$. ∎
Proposition 10.2.
Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, $A$ and $B$ are (discrete) abelian groups, and $\varphi:S\rightarrow\mathrm{Hom}(A,B)$ is a continuous map. For every finitely generated subgroup $A_{0}$ of $A$, there is an $i\in\Lambda$ and a map $\varphi_{0}:S_{i}\rightarrow\mathrm{Hom}(A_{0},B)$ such that, for all $s\in S$, we have
$\varphi(s)\restriction A_{0}=\varphi_{0}(\pi_{i}(s)).$
Proof.
Let $X_{0}$ be a finite subset of $A$, and let $A_{0}$ be the subgroup of $A$ generated by $X_{0}$. For each $\psi:X_{0}\rightarrow B$, the set $S_{\psi}:=\{s\in S\mid\varphi(s)\restriction X_{0}=\psi\}$ is a clopen subset of $S$, by the continuity of $\varphi$. By the compactness of $S$, there are only finitely many $\psi$ for which $S_{\psi}$ is nonempty. Therefore, by Proposition 10.1, there is $i\in\Lambda$ and $f:S_{i}\rightarrow{{}^{X_{0}}}B$ such that, for all $s\in S$, we have $\varphi(s)\restriction X_{0}=f(\pi_{i}(s))$.
Since $A_{0}$ is generated by $X_{0}$, every element of $\mathrm{Hom}(A_{0},B)$ is determined by its restriction to $X_{0}$. Therefore, for each $\bar{s}\in S_{i}$, there is a unique extension $\varphi_{0}(\bar{s})$ of $f(\bar{s})$ to an element of $\mathrm{Hom}(A_{0},B)$. [^4] Such an extension must exist because, if $s\in S$ is such that $\pi_{i}(s)=\bar{s}$, then $\varphi(s)\restriction A_{0}$ is such an extension; the uniqueness then follows from the previous sentence. It then follows that, for all $s\in S$, we have $\varphi(s)\restriction A_{0}=\varphi_{0}(\pi_{i}(s))$, as desired. ∎
Recall the definition of Chase’s condition from the end of Section 9. We now show that Theorem 9.4 transfers to the richer setting of $\mathsf{CondAb}_{\omega_{1}}$. By clause (1) of Theorem 1.3, this is sharp, since for any nonfree abelian group of size $\aleph_{1}$, the profinite set $2^{\omega_{1}}$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{2}}}(\underline{A},\underline{\mathbb{Z}})\neq 0$.
Theorem 10.3.
Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of cardinality $\omega_{1}$ satisfying Chase’s condition. Then
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{1}}}(\underline{A},\underline{\mathbb{Z}})=0.$
Proof.
Let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be the canonical free resolution of $A$, let $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, and let $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be continuous. We must produce a continuous map $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ such that $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$. We fix a basis $\{\bar{a}\mid a\in A\}$ for $F$; the surjection from $F$ to $A$ in the free resolution above is then generated by the map that sends $\bar{a}$ to $a$ for all $a\in A$. For each subgroup $A^{\prime}\subseteq A$, let $F_{A^{\prime}}$ be the free group generated by $\{\bar{a}\mid a\in A^{\prime}\}$, and let $K_{A^{\prime}}$ be the kernel of the canonical surjection from $F_{A^{\prime}}$ to $A^{\prime}$.
Claim 10.4.
Suppose that $A^{\prime}$ is a free subgroup of $A$, with a basis $B^{\prime}$, and let $f:S\rightarrow{{}^{B^{\prime}}}\mathbb{Z}$ be a continuous function (where $\mathbb{Z}$ is discrete and ${{}^{B^{\prime}}}\mathbb{Z}$ is given the product topology). Then there is a unique continuous map $\psi^{\prime}:S\rightarrow\mathrm{Hom}(F_{A^{\prime}},\mathbb{Z})$ such that
  1. (1)
    for all $s\in S$ and $b\in B^{\prime}$, $\psi^{\prime}(s)(\bar{b})=f(s)(b)$;
  2. (2)
    for all $s\in S$, $\psi^{\prime}(s)\restriction K_{A^{\prime}}=\varphi(s)\restriction K_{A^{\prime}}$.
Proof.
Fix $s\in S$. To fulfill requirement (1), we are obliged to let $\psi^{\prime}(s)(\bar{b}):=f(s)(b)$ for all $b\in B^{\prime}$. Now fix $a\in A^{\prime}\setminus B^{\prime}$. Then there is a unique way to express $a$ in the form $\sum_{b\in B^{*}}c_{b}b$, where $B^{*}\subseteq B^{\prime}$ is finite and, for all $b\in B^{*}$, $c_{b}$ is a nonzero integer. To satisfy requirement (2), we are obliged to let
$\psi^{\prime}(s)(\bar{a}):=\varphi(s)\left(\bar{a}-\sum_{b\in B^{*}}c_{b}\bar{b}\right)+\sum_{b\in B^{*}}c_{b}f(s)(b).$
We must then extend $\psi^{\prime}(s)$ linearly to the rest of $F_{A^{\prime}}$. It is routine to verify that $\psi^{\prime}(s)$ thusly defined is in $\mathrm{Hom}(F_{A^{\prime}},\mathbb{Z})$ satisfying (1) and (2), and the continuity of $\psi$ follows immediately from the continuity of $\varphi$ and $f$. The uniqueness of $\psi^{\prime}$ is evident from the fact that, as the above construction makes clear, all values of $\psi^{\prime}(s)$ were entirely determined by requirements (1) and (2) and the necessity for $\psi^{\prime}(s)$ to be a homomorphism. ∎
We define a forcing notion $\mathbb{P}$ to which we will apply $\sf MA_{\omega_{1}}$. Conditions in $\mathbb{P}$ are all pairs of the form $p=(A_{p},\psi_{p})$ such that
  • $A_{p}$ is a finitely-generated pure subgroup of $A$;
  • $\psi_{p}:S\rightarrow\mathrm{Hom}(F_{A_{p}},\mathbb{Z})$ is continuous;
  • for all $s\in S$, $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$.
If $p,q\in\mathbb{P}$, then $q\leq p$ if and only if $A_{q}\supseteq A_{p}$ and, for all $s\in S$, $\psi_{q}(s)\restriction F_{A_{p}}=\psi_{q}(s)$.
Claim 10.5.
For all $a\in A$, the set $D_{a}:=\{p\in\mathbb{P}\mid a\in A_{p}\}$ is dense in $\mathbb{P}$.
Proof.
Fix $a\in A$ and $p\in\mathbb{P}$. We will find $q\leq p$ with $q\in D_{a}$. Since $A$ is $\aleph_{1}$-free, we can find a finitely generated pure subgroup $A^{\prime}$ of $A$ such that $A_{p}\cup\{a\}\subseteq A^{\prime}$. Then $A^{\prime}/A_{p}$ is free, so we can fix a basis for $A^{\prime}$ of the form $B_{0}\cup B_{1}$, where $B_{0}$ is a basis for $A_{p}$. Define a function $f:S\rightarrow{{}^{B_{0}\cup B_{1}}}\mathbb{Z}$ as follows: for all $s\in S$ and $b\in b_{0}\cup B_{1}$, let
$f(s)(b):=\begin{cases}\psi_{p}(s)(b)&\text{if }b\in B_{0}\\ 0&\text{if }b\in B_{1}.\end{cases}$
Since $\psi_{p}$ is continuous, it follows that $f$ is continuous. Let $\psi^{\prime}:S\rightarrow\mathrm{Hom}(F_{A^{\prime}},\mathbb{Z})$ be the unique continuous map given by Claim 10.4 applied to $A^{\prime}$, $B_{0}\cup B_{1}$, and $f$. Then $q:=(A^{\prime},\psi^{\prime})\in D_{a}$. Moreover, the uniqueness clause of Claim 10.4 together with the definition of $f$ implies that $\psi^{\prime}(s)\restriction F_{A_{p}}=\psi_{p}(s)$ for all $s\in S$, so we have $q\leq p$. ∎
Claim 10.6.
$\mathbb{P}$ has the ccc.
Proof.
Let $\vec{p}=\langle p_{\eta}\mid\eta<\omega_{1}\rangle$ be a sequence of conditions in $\mathbb{P}$. By (the proof of) [5, Lemma 7.5], by thinning out $\vec{p}$ to some cofinal subsequence if necessary, we can assume that there is a free subgroup $A^{\prime}$ of $A$ that is pure in $A$ and such that $A_{p_{\eta}}\subseteq A^{\prime}$ for all $\eta<\omega_{1}$ (this is the only use of the full power of Chase’s condition in the proof). Let $B^{\prime}$ be a basis for $A^{\prime}$. Using Claim 10.5 and extending each $p_{\eta}$ if necessary, we can assume that, for all $\eta<\omega_{1}$, $A_{p_{\eta}}$ is freely generated by some finite subset $B^{\prime}_{\eta}$ of $B^{\prime}$. By thinning out $\vec{p}$ again if necessary, we can assume that the sets $\{B^{\prime}_{\eta}\mid\eta<\omega_{1}\}$ form a $\Delta$-system, with root $R$.
For all $\eta<\omega_{1}$, define a map $f_{\eta}:S\rightarrow{{}^{B^{\prime}_{\eta}}}\mathbb{Z}$ by letting $f_{\eta}(s)(b):=\psi_{p_{\eta}}(s)(\bar{b})$ for all $s\in S$ and $b\in B^{\prime}_{\eta}$. By Proposition 10.2, we can fix $i_{\eta}\in\Lambda$ and a map $g_{\eta}:S_{i_{\eta}}\rightarrow{{}^{R}}\mathbb{Z}$ such that, for all $s\in S$ and $b\in R$, we have $f_{\eta}(s)(b)=g_{\eta}(\pi_{i}(s))(b)$. Since there are only countably many choices for $i_{\eta}$ and $g_{\eta}$, we can find a fixed $i$ and $g$ and ordinals $\eta<\xi<\omega_{1}$ such that $i_{\eta}=i_{\xi}=i$ and $g_{\eta}=g_{\xi}=g$. Note that, for all $s\in S$ and $b\in R$, this implies that $f_{\eta}(s)(b)=f_{\xi}(s)(b)$.
Let $B^{*}:=B_{\eta}\cup B_{\xi}$, and let $A^{*}$ be the free group generated by $B^{*}$. Then $A^{*}$ is a pure subgroup of $A^{\prime}$; since $A^{\prime}$ is pure in $A$, this implies that $A^{*}$ is also pure in $A$. Define a function $f:S\rightarrow{{}^{B^{*}}}\mathbb{Z}$ by letting $f(s)=f_{\eta}(s)\cup f_{\xi}(s)$ for all $s\in S$. Since $\psi_{p_{\eta}}$ and $\psi_{p_{\xi}}$ are continuous (and hence $f_{\eta}$ and $f_{\xi}$ are continuous), it follows that $f$ is continuous. Let $\psi^{*}:S\rightarrow\mathrm{Hom}(F_{A^{*}},\mathbb{Z})$ be the unique continuous map obtained by applying Claim 10.4 to $A^{*}$, $B^{*}$, and $f$. Then $q=(A^{*},\psi^{*})\in\mathbb{P}$ and, by the uniqueness clause of Claim 10.4, it extends both $p_{\eta}$ and $p_{\xi}$. Therefore, $\vec{p}$ does not enumerate an antichain, and hence $\mathbb{P}$ has the ccc. ∎
Now apply $\sf MA_{\omega_{1}}$ to obtain a filter $G\subseteq\mathbb{P}$ such that, for all $a\in a$, we have $G\cap D_{a}\neq\emptyset$. Define $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ by letting $\psi(s)=\bigcup\{\psi_{p}(s)\mid p\in G\}$. The fact that each $\psi_{p}$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$ implies that $\psi$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K=\varphi(s)$, as desired. ∎
We thus obtain the following corollary, establishing Theorem 1.4(1).
Corollary 10.7.
If $\sf MA_{\omega_{1}}$ holds, then there is a nonfree abelian group $A$ of size $\aleph_{1}$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\omega_{1}}$.
11. A different vantage point
In this section, we describe and apply a lemma providing an alternative, more purely algebraic path to several of our results, as well as to a few new ones. Recall from Section 8 that if $\mathbb{B}$ is the complete Boolean algebra of regular open subsets of an extremally disconnected compact Hausdorff space $S$, then $C(S,\mathbb{Z})$ is precisely the group of $\mathbb{B}$-names for integers.
Lemma 11.1.
For all abelian groups $A$ and extremally disconnected profinite sets $S$,
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z})).$
Proof.
The chain of equalities is as follows:
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ $=\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A}\otimes\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}})$
$=\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}}))$
$=\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathrm{Hom}}_{\mathsf{Cond}}(\underline{S},\underline{\mathbb{Z}}))$
$=\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{C(S,\mathbb{Z}}))$
$=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))$
To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [13, Prop. 3.5 and Prop. 3.2]. The first equality follows from [20, p. 13 (iii) (cf. p. 25)]. This leaves us only with the second. Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [20, p. 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$. More particularly, letting $\underline{\mathcal{P}}$, much as in the proof of Lemma 3.2, denote the condensation of a free resolution of $\underline{A}$, we have:
$\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A}\otimes\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}})$ $=\mathrm{H}^{1}(\mathrm{Hom}_{\mathsf{CondAb}}(\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}}))$
$=\mathrm{H}^{1}(\mathrm{Hom}_{\mathsf{CondAb}}(\underline{\mathcal{P}},\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}}))$
$=\mathrm{Ext}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\mathbb{Z}[\underline{S}],\underline{\mathbb{Z}})),$
and this chain of equalities, if sound, will complete the argument. Let us check the details. Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [13, Prop. 2.18] and Proposition 12.3 below; this shows the third equality. The second is the aforementioned adjunction. For the first, we have noted already that $\mathbb{Z}[\underline{S}]$ is flat, so the question reduces to that of whether the terms of $\underline{\mathcal{P}}\otimes\mathbb{Z}[\underline{S}]$ are projective. For this observe simply that $\big{(}\underline{\bigoplus_{I}\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\big{(}\bigoplus_{I}\underline{\mathbb{Z}}\big{)}\otimes\mathbb{Z}[\underline{S}]=\bigoplus_{I}\big{(}\underline{\mathbb{Z}}\otimes\mathbb{Z}[\underline{S}]\big{)}=\bigoplus_{I}\mathbb{Z}[\underline{S}]$ for any set $I$; that such sums are projective follows from Proposition 12.3 below. ∎
Note that the lemma allows us to recast this paper’s essential background in uniformly classical terms: Shelah’s result was that the implication
$\big{[}\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for }S=*\,\big{]}\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
is independent of the $\mathsf{ZFC}$ axioms. Clausen and Scholze’s result, in contrast, can now be seen to follow from the fact that the implication
$\big{[}\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for all }S\in\mathsf{ED}\,\big{]}\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
is indeed a $\mathsf{ZFC}$ theorem. This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([16], see also [6, Proposition XI.4.4]), generalizing work of Specker [23], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
In fact, we may be more precise. Observe that, if $A$ is a nonfree abelian group of cardinality $\kappa$, then the canonical free resolution of $A$ yields $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\neq 0$, where $\mathbb{Z}^{(\kappa)}$ denotes the free abelian group on $\kappa$-many generators. Moreover, if $A$ is an abelian group and $A_{0}$ is a nonfree subgroup of $A$ of cardinality $\kappa$, then an application of $\mathrm{Hom}(\cdot,\mathbb{Z}^{(\kappa)})$ to the short exact sequence $0\rightarrow A_{0}\rightarrow A\rightarrow A/A_{0}\rightarrow 0$ yields a long exact sequence, a portion of which is
$\ldots\rightarrow\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\rightarrow\mathrm{Ext}^{1}_{\mathsf{Ab}}(A_{0},\mathbb{Z}^{(\kappa)})\rightarrow 0.$
It follows that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\neq 0$. Therefore, Lemma 11.1 immediately yields the following corollary, providing an alternative proof of clauses (1) and (3) of Theorem 1.3.
Corollary 11.2.
Suppose that $A$ is a nonfree abelian group and $\kappa$ is the least cardinality of a nonfree subgroup of $A$. Then
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$
for every $S\in\mathsf{ED}$ of weight at least $\kappa$. ∎
Lemma 11.1 also sheds additional light on the results of Section 10, both providing an alternative proof of a variation on the main result of that section, as well as showing that the existence of large cardinals provides a bound on the extent to which this result can be generalized.
We first note that, by results in [7], it is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$. Thus, Lemma 11.1 yields the following corollary, establishing clause (2) of Theorem 1.4 and indicating that, consistently, the positive answer to the interpretation of Whitehead’s problem in $\mathsf{CondAb}$ established by Theorem 1.2 fails in $\mathsf{CondAb}_{\kappa}$ for every uncountable cardinal $\kappa$.
Corollary 11.3.
It is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for all $S\in\mathsf{ED}_{\kappa}$. In particular, it is consistent that, for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$.
On the other hand, the existence of sufficiently large cardinals provides a limit on the extent to which the phenomena of Corollary 11.3 can occur. In particular, if $\kappa$ is a strongly compact cardinal, then $\mathsf{CondAb}_{\kappa}$ is sufficiently rich to guarantee a positive answer to the appropriate interpretation of Whitehead’s problem. The following corollary yields clause (3) of Theorem 1.4.
Corollary 11.4.
Suppose that $\kappa$ is a strongly compact cardinal and $A$ is an abelian group. If $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$, then $A$ is free.
Proof.
Suppose that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$. In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$. By Lemma 11.1, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$. By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free. By [6, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired. [^5] As [6, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
Remark 11.5.
In the proof of Corollary 11.4, the crucial property of the cardinal $\kappa$ was the fact that every $\kappa$-free group is free. This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$. In particular, it is shown in [12, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
12. Locally compact abelian groups
In this section, we show that Theorem 1.2 extends from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones. In particular, we prove the following theorem.
Theorem 12.1.
Let $A$ be an object of the category $\mathsf{LCA}$ of locally compact abelian groups. Then
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$
if and only if $A$ is projective in $\mathsf{LCA}$.
Proof.
By the structure theorem for locally compact abelian groups (see [20, Thm. 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some $n\in\mathbb{N}$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
(1) $0\to C\to B\to D\to 0$
with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well. [^6] See [13, Props. 4.9, 2.18]. Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [20, p. 25], we have
$(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0)\Leftrightarrow(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0).$
Consider the following portion of the long exact sequence deriving from equation (1):
$\cdots\to 0\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{D},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})\to 0\to\cdots$
(The rightmost 00 follows from [20, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [20, Prop. 4.2].) Hence
$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0$
if and only if
  1. (1)
    $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{D},\underline{\mathbb{Z}})=0$ and
  2. (2)
    $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=0$.
By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$. Also, the argument of [13, Prop. 4.11] together with [20, Prop. 4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([20, Prop. 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism. Again applying [20, Prop. 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$. Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$. As noted, by [15, Prop. 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$. ∎
This theorem carries as corollary a negative answer to one other condensed version of Whitehead’s Problem:
Corollary 12.2.
As the condensed abelian group $\underline{\mathbb{R}}$ witnesses, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$ does not imply that $A$ is projective in $\mathsf{CondAb}$.
To see this, recall that, as noted, the compact projective condensed abelian groups of the form $\mathbb{Z}[\underline{S}]$ with $S\in\mathsf{ED}$ generate $\mathsf{CondAb}$ ([3, p. 16–17 and Proposition 3.5]); in particular, for any condensed abelian group $X$ there exists a set $I$, a collection $\{S_{i}\mid i\in I\}$ of spaces in $\mathsf{ED}$, and an epimorphism $r:\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]\to X$. If $X$ is projective then $r$ admits a right inverse $s$, hence:
Proposition 12.3.
The projective objects of $\mathsf{CondAb}$ are exactly the retracts of direct sums of groups of the form $\mathbb{Z}[\underline{S}]$ with $S$ extremally disconnected.
Hence it will suffice to show that $\mathbb{R}$ is not a retract of such a sum. This is immediate from the following lemma.
Lemma 12.4.
The underlying group of $\mathbb{Z}[\underline{S}]$ is the free abelian group $\bigoplus_{S}\mathbb{Z}$.
By the lemma, any section $s:\underline{\mathbb{R}}\to\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]$ of an $r$ as above would evaluate over a point $\ast$ as an injection $s(*):\mathbb{R}\to\bigoplus_{I}\bigoplus_{S_{i}}\mathbb{Z}$, but any nonzero element in the image of $s(*)$ would need to be divisible by every positive $n$, a contradiction. The following, therefore, will complete the proof of Corollary 12.2.
Proof of Lemma 12.4.
As noted, $\mathbb{Z}[\underline{S}]$ is the sheafification of the presheaf
(2) $T\mapsto\bigoplus_{\mathrm{Cont}(T,S)}\mathbb{Z}$
Recall also that the sheafification $\mathcal{F}^{\sharp}$ of a presheaf $\mathcal{F}$ on a site is given by the twofold iteration of the $+$ operation, where
$\mathcal{F}^{+}(U)=\mathrm{colim}_{\mathcal{U}\in\mathrm{cov}(U)}\,\,\mathrm{H}^{0}(\mathcal{U};\mathcal{F}).$
Here
$\mathrm{H}^{0}(\mathcal{U};\mathcal{F})=\{(s_{i})\in\prod_{i\in I}\mathcal{F}(U_{i})\mid s_{i}|_{U_{i}\times_{U}U_{j}}=s_{j}|_{U_{i}\times_{U}U_{j}}\textnormal{ for all }i,j\in I\}$
[…]
13. Conclusion
Though the project leading to this article began with a desire simply to understand the combinatorial phenomena underlying considerations of Whitehead’s problem in the setting of condensed abelian groups, it quickly expanded into the beginnings of a more general investigation into connections between condensed mathematics and forcing, an investigation that we hope to see further developed in the future. These connections have manifested herein in a number of ways, notably in
  1. (1)
    the recognition that condensed objects can often be viewed as organized presentations of forcing names; and
  2. (2)
    the observation that interesting phenomena arise via the evaluation of condensed objects at the Stone spaces of complete Boolean algebras significant in set theoretic forcing (cf. clause (3) of Theorem 1.3).
In the context of Whitehead’s problem, these investigations culminated in our Theorem 9.2 that, if $A$ is an abelian group and $\mathbb{B}$ is a complete Boolean algebra which forces a certain technical strengthening of the failure of $A$ to be Whitehead, then $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. We conjecture that this technical strengthening can be foregone:
Conjecture 13.1.
Suppose that $A$ is an abelian group, $\mathbb{B}$ is a complete Boolean algebra, and
$\Vdash_{\mathbb{B}}``A\text{ is not Whitehead}".$
Then $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$.
Note that the converse of this conjecture does not hold, as shown in the discussion following Corollary 9.3. We feel that a resolution of this conjecture will shed light not just on the relationship between forcing and condensed mathematics, but also on considerations of Whitehead’s problem in the classical setting.
Finally, there remain questions about alternative interpretations of Whitehead’s problem in the condensed setting. In Section 12, we initiated some further exploration of two main parameters in play in condensed interpretations of Whitehead’s problem, namely the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$. After all, in retrospect, it’s not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless). A natural next question, then, is how much further this $d$ may be pushed. By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$. Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction. Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [20, §5–7] and [19, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2. Within it, in addition, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [20, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
Question 13.2.
Let $A$ be a solid abelian group. Does $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0$ (or equivalently $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$) imply that $A$ is projective?
Part of the question is whether the answer will follow from the $\mathsf{ZFC}$ axioms alone. The solid setting, being “very much a ‘nonarchimedean’ notion” [20, footnote 10 on p. 32], carries the additional virtue of circumventing the aforementioned issue of $\underline{\mathbb{R}}$. More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [20, Corollary 6.1(iii)]; in particular, $\underline{\mathbb{R}}$ is not solid. This leads to the following natural variation on Question 13.2: Suppose that $A$ is a condensed abelian group and $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$. Must $A^{\blacksquare}$ be projective in the category of solid abelian groups?

### v3

@@
 03E35, 03E40, 03E05, 13C10, 20J05
-The first author was supported by a María Zambrano Fellowship at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S.
+The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S.
 A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023.
@@
 The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023.
-We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks.
+We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks, and to thank Peter Scholze for valuable remarks on the paper as well.
 1. Introduction
@@
 Progress on the general question was slow, though, and the latter soon assumed the status of a major open problem in homological algebra.
-Shelah’s solution to the question 20 years later was completely unexpected, for Whitehead’s problem thereby became one of the first originating outside of logic or set theory to be proven independent of the $\sf ZFC$ axioms.
+Shelah’s solution to the question 20 years later was completely unexpected, for Whitehead’s problem thereby became one of the first originating unequivocally outside of logic or set theory to be proven independent of the $\sf ZFC$ axioms.
 Theorem 1.1 (Shelah, [27]).
@@
 We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
-We then exhibit a correspondence between satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$.
+We then exhibit a correspondence between the satisfaction of certain simple formulas concerning a topological space $X$ in two settings: first, in a Boolean-valued model $V^{\mathbb{B}}$, and second in the context of the space $C(S,X)$ of continuous functions from the Stone space $S$ of $\mathbb{B}$ to $X$.
 This correspondence, stated in Theorem 9.1, allows us, e.g., to immediately deduce clause (3) of Theorem 1.3 from clause (2) of the same theorem.
@@
 Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups.
+We record as well several observations by Peter Scholze showing that Theorem 1.2 does not extend to the category of solid abelian groups.
 The structure of the paper is as follows.
@@
 Section 9 then establishes the aforementioned equivalence between the settings of $C(S,X)$ and $V^{\mathbb{B}}$ with respect to interpretations of certain formulas.
-We expect the results of this section to be applicable in a wide variety of settings, and derive clause (3) of Theorem 1.3 as an immediate corollary.
+We expect the results of this section to be applicable in a wide variety of contexts, and derive clause (3) of Theorem 1.3 as an immediate corollary.
 In Section 10, we prove Theorem 1.4(1).
@@
 This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4.
-In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups.
-We close out our work in Section 13 with a few concluding remarks and main open questions.
+In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups, and record impediments to its extension to the class of solid abelian groups.
+We close out our work in Section 13 with a few concluding remarks and one main open question.
 We have pursued wherever possible an account of our results accessible to readers with no more than a basic knowledge of either set theory or homological algebra.
@@
 When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf.
-[20, Proposition 1.7]).
+[26, Proposition 1.7]).
 Similarly, given a topological abelian group $A$, one can define a condensed abelian group $\underline{A}$ by letting $\underline{A}(S)=C(S,A)$ (considered as an abelian group with pointwise addition) for all $S\in\mathsf{Prof}$.
@@
 For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$.
-Concretely, as noted in [20, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$.
+Concretely, as noted in [26, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$.
 Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$.
@@
 If $A$ and $B$ are sufficiently nice topological groups, then $\underline{\mathrm{Hom}}(\underline{A},\underline{B})$ has a particularly nice description:
-Proposition 3.1 (Clausen–Scholze, [20, Proposition 4.2]).
+Proposition 3.1 (Clausen–Scholze, [26, Proposition 4.2]).
 Suppose that $A$ and $B$ are Hausdorff topological abelian groups and $A$ is compactly generated. Then there is a natural isomorphism of condensed abelian groups
@@
 Proof.
-As noted in [20, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
+As noted in [26, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
 $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(T,M)(S)=R\mathrm{Hom}_{\mathsf{CondAb}}(T\otimes\mathbb{Z}[S],M).$
@@
 When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$.
-To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [13, Props.
+To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [18, Props.
 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$.
@@
 Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$.
-To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [13, Proposition 2.19].
-By [13, Remark 3.10 and Lemma 4.2], we then have
+To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [18, Proposition 2.19].
+By [18, Remark 3.10 and Lemma 4.2], we then have
 $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=\prod_{\kappa}R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})=\prod_{\kappa}\underline{\mathbb{Z}}[0],$
@@
 is a surjection.
-Recall that $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)$ consists of all continuous functions $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ (and similarly for $K$), where $\mathrm{Hom}(F,\mathbb{Z})$ is given the compact-open topology (which, since $F$ and $\mathbb{Z}$ are discrete, is simply the topology $\mathrm{Hom}(F,\mathbb{Z})$ inherits as a subspace of $\prod_{F}\mathbb{Z}$ endowed with the product topology).
+Recall that $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)$ consists of all continuous functions $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ (and similarly for $K$), where $\mathrm{Hom}(F,\mathbb{Z})$ is given the compact-open topology (which, since $F$ and $\mathbb{Z}$ are discrete, is simply the topology $\mathrm{Hom}(F,\mathbb{Z})$ inherits as a closed subspace of $\prod_{F}\mathbb{Z}$ endowed with the product topology).
 Moreover, the map $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$ sends a continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ to the map $\phi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ defined by letting $\phi(s)=\psi(s)\restriction K$ for all $s\in S$.
 Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$).
-The construction of such $S$ and $\varphi$ will be the subject of the next sections.
+The construction of such an $S$ and $\varphi$ will be the subject of the next sections.
 4. Countable groups
@@
 The following theorem gives a useful characterization of countable free abelian groups (recall that a finitely generated abelian group is free if and only if it is torsion-free).
-Theorem 4.1 ([17] (cf. [5, Theorem 4.2])).
+Theorem 4.1 ([23] (cf. [9, Theorem 4.2])).
 Suppose that $A$ is a countable torsion-free abelian group. Then $A$ is free if and only if every finitely generated subgroup of $A$ is contained in a finitely generated pure subgroup of $A$.
@@
 Being free, $K$ is isomorphic to the direct sum of copies of $\mathbb{Z}$ indexed by $B_{K}$.
-We consider the canonical embedding $\nu:K\to\mathbb{Z}^{B_{K}}$, i.e.
-for every $t,s\in B_{K}$, $\nu(t)(t)=1$ and $\nu(t)(s)=0$ if $s\neq t$.
-Then $\nu$ is pure, i.e.
-$\nu[K]$ is pure in $\mathbb{Z}^{B_{K}}$ (in fact, $\nu$ is even an elementary embedding).
+We consider the canonical embedding $\nu:K\to\mathbb{Z}^{B_{K}}$, i.e., for every $t,s\in B_{K}$, $\nu(t)(t)=1$ and $\nu(t)(s)=0$ if $s\neq t$.
+Then $\nu$ is pure, i.e., $\nu[K]$ is pure in $\mathbb{Z}^{B_{K}}$ (in fact, $\nu$ is even an elementary embedding).
 Since $K$ is not pure in $F$, it follows that there does not exist any homomorphism $\mu:F\to\mathbb{Z}^{B_{K}}$ such that $\mu\restriction K=\nu$.
 For each $t\in B_{K}$, let us denote by $p_{t}:\mathbb{Z}^{B_{K}}\to\mathbb{Z}$ the $t$-th canonical projection.
-Since any homomorphism into direct product is uniquely determined by specifying its compositions with all the canonical projections, there has to be some $t^{*}\in B_{K}$ such that $\varphi:=p_{t^{*}}\nu$ cannot be extended to a homomorphism from $F$ into $\mathbb{Z}$.
+Since any homomorphism into a direct product is uniquely determined by specifying its compositions with all the canonical projections, there has to be some $t^{*}\in B_{K}$ such that $\varphi:=p_{t^{*}}\nu$ cannot be extended to a homomorphism from $F$ into $\mathbb{Z}$.
 One readily checks that $t^{*}$ and $\varphi$ possess the desired properties.
@@
 The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf.
-[27, Proposition 3.1] or [18, Theorem 3.3.2].
+[34, Proposition 3.1] or [24, Theorem 3.3.2].
 Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@
 […]
-It follows that $F$ is countable and thus free since all countable subgroups of $\mathbb{Z}^{\omega}$ are such, cf.
-[6, Theorems IV.2.3 and IV.2.8].
+It follows that $F$ is countable and thus free since all countable subgroups of $\mathbb{Z}^{\omega}$ are; see [10, Theorems IV.2.3 and IV.2.8].
 Since $\mu$ extends the map $\nu$, we see, in particular, that $p\mu$ extends the map $p\nu$ for any $p\in\mathrm{Hom}(\mathbb{Z}^{\omega},\mathbb{Z})$.
@@
 It remains to notice that any $\varphi\in\mathrm{Hom}(\mathbb{Z}^{(\omega)},\mathbb{Z})$ which is nonzero only on finitely many elements from $B$ is of the form $p\nu$ for $p$ as above.
-On the other hand, since $\mathbb{Z}$ is a slender group ([6, Corollary III.2.4]), no homomorphism $\varphi:\mathbb{Z}^{(\omega)}\to\mathbb{Z}$ which is nonzero on infinitely many elements from $B$ can be extended to a homomorphism from $\mathbb{Z}^{\omega}$ into $\mathbb{Z}$.
+On the other hand, since $\mathbb{Z}$ is a slender group ([10, Corollary III.2.4]), no homomorphism $\varphi:\mathbb{Z}^{(\omega)}\to\mathbb{Z}$ which is nonzero on infinitely many elements from $B$ can be extended to a homomorphism from $\mathbb{Z}^{\omega}$ into $\mathbb{Z}$.
 Of course, it can happen that a particular such $\varphi$ can be extended to a homomorphism from $F$ into $\mathbb{Z}$.
@@
 Note that $2^{\kappa}$ is a profinite set.
+Our next result establishes item (1) of Theorem 1.3.
 Theorem 5.1.
@@
 We can therefore assume without loss of generality that $A$ is almost free.
-Thus, for the rest of the proof we may assume that $A$ is almost free and $|F|=\kappa$.
+Thus, for the rest of the proof we will assume that $A$ is almost free and $|F|=\kappa$.
 The case of $\kappa=\aleph_{0}$ is covered by Theorem 4.2, so we may assume that $\kappa>\aleph_{0}$.
@@
 Since $A$, $F$, and $K$ each have size at most $\kappa$ and $\vec{M}$ is $\in$-increasing, we have $A=\bigcup_{\alpha<\kappa}A_{\alpha}$, $F=\bigcup_{\alpha<\kappa}F_{\alpha}$, and $K=\bigcup_{\alpha<\kappa}K_{\alpha}$.
-Let $E$ be the set of $\alpha<\kappa$ such that there exists $\beta$ such that $\alpha<\beta<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree.
+Let $E$ be the set of $\alpha<\kappa$ such that there exists a $\beta$ such that $\alpha<\beta<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree.
 Note that, if $\alpha<\beta<\beta^{\prime}<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree, then also $A_{\beta^{\prime}}/A_{\alpha}$ is nonfree, since it contains $A_{\beta}/A_{\alpha}$ as a subgroup.
@@
 Therefore, it is straightforward to build a basis for $A$ by recursion along $D$ (cf.
-[5, Theorem 5.3]), contradicting the assumption that $A$ is not free.
+[9, Theorem 5.3]), contradicting the assumption that $A$ is not free.
 For all $\alpha\in E$, let $\alpha<\beta_{\alpha}<\kappa$ be such that $A_{\beta_{\alpha}}/A_{\alpha}$ is not free.
@@
 Similarly, let $B_{F,\alpha}^{*}:=B_{F}\cap(M_{\alpha+1}\setminus M_{\alpha})$ and $B_{K,\alpha}^{*}:=B_{K}\cap(M_{\alpha+1}\setminus M_{\alpha})$.
-Note that each of $F_{\alpha}^{*}$ and $K_{\alpha}^{*}$ is free, with bases given by $\{z+F_{\alpha}\mid z\in B_{F,\alpha}^{*}\}$ and $\{t+K_{\alpha}\mid t\in B_{K,\alpha}^{*}\}$, respectively.
+Note that each $F_{\alpha}^{*}$ and $K_{\alpha}^{*}$ is free, with bases given by $\{z+F_{\alpha}\mid z\in B_{F,\alpha}^{*}\}$ and $\{t+K_{\alpha}\mid t\in B_{K,\alpha}^{*}\}$, respectively.
 Also, our free resolution of $A$ induces a free resolution
@@
 Suppose that $A$ is a nonfree abelian group and $\kappa=|A|$. Then, in $V^{\mathrm{Add}(\omega,\kappa)}$, $A$ is not Whitehead.
-Before proving Theorem 6.1, we make the following observation.
+Note that this establishes item (2) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
 Lemma 6.2.
@@
 We claim that $\dot{\varphi}$ is forced to be as desired.
-To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
+To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
 […]
@@
 A more indirect route, involving a broader investigation into more general connections between condensed mathematics and forcing, occupies Sections 8 and 9, which can be read independently of this one.
-Given an infinite cardinal $\kappa$, let $\mathbb{B}_{\kappa}$ denote the Boolean completion of $\mathrm{Add}(\omega,\kappa)$, and let $S_{\kappa}$ denote the Stone space $S(\mathbb{B}_{\kappa})$.
+Given an infinite cardinal $\kappa$, let $\mathbb{B}_{\kappa}$ denote the Boolean completion of $\mathrm{Add}(\omega,\kappa)$, and let $S_{\kappa}$ denote the Stone space of $\mathbb{B}_{\kappa}$.
 We will sometimes slightly abuse notation and think of $\mathrm{Add}(\omega,\kappa)$ as a subset of $\mathbb{B}_{\kappa}$.
@@
     Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$.
-Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$.
+Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(U)(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$.
 Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
@@
 We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$.
-Suppose for sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise.
+Suppose for the sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise.
 Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
@@
 Proof.
-Suppose for sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$.
+Suppose for the sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$.
 By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$.
@@
 Proof.
-Suppose for sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$.
+Suppose for the sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$.
 By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$.
@@
 Proposition 8.10.
-For all $U\in\tau$, we have $k[N_{U}]=\hat{U}_{K}$.
+For all $U\in\tau$, we have $k[N_{U}]=\hat{U}_{\mathcal{K}}$.
 Proof.
@@
 First, as was mentioned above, if $X$ is a locally compact Hausdorff space, then $\hat{X}_{\mathcal{K}}$ is the canonical interpretation of $X$ in $V^{\mathbb{B}}$ in the sense of [36].
-Recall also that the category $\mathsf{ED}$ is precisely the category of Stone spaces of complete Boolean algebras; recall as well that $\underline{X}(S)$ equals, by definition, $C(S,X)$ for any $S\in\mathsf{ED}$.
-Therefore, the results of this section show that one way to think of the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, for all such spaces $\underline{X}$, the condensed set $\underline{X}$ is precisely an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
-Second, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products.
-Namely, if $X_{i}$ is a Hausdorff space for all $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$.
+Recall also that the category $\mathsf{ED}$ is precisely the category of Stone spaces of complete Boolean algebras, and that $\underline{X}(S)$ equals, by definition, $C(S,X)$ for any $S\in\mathsf{ED}$.
+Therefore, the results of this section show that one way to think about the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, as noted, for all such spaces $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
+Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [26, pp.
+8-9 and Appendix to Lecture II]).
+Observe next that in the $\mathbb{B}=\{0,1\}$ (and hence $S(\mathbb{B})=*$) instance of Theorem 8.11, $\hat{X}_{\mathcal{K}}$ is homeomorphic to $X$.
+Had we endowed it instead with the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology considered above, it would be homeomorphic to $cg(X)$, and the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology corresponds more generally to that given by post-composing the conversion recorded in Theorem 8.11 with the functor $cg$.
+Third, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products.
+Namely, if $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$.
 This is because the collection
@@
 is $\subseteq$-cofinal in $\mathcal{K}(\prod_{i\in I}X_{i})$.
-Third, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact.
+Fourth, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact.
 In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11.
@@
 Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free.
-For concreteness, suppose it is constructed as in [5, Theorem 7.3].
+For concreteness, suppose it is constructed as in [9, Theorem 7.3].
 It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free.
@@
 We present a more algebraic path to related results in Section 11.
-For a given abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\aleph_{1}}$, then $A$ is Whitehead in the classical sense, since
+For a given abelian group $A$, if $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\omega_{1}}$, then $A$ is Whitehead in the classical sense, since
 $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{1}}}(\underline{A},\underline{\mathbb{Z}})(\ast)=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}).$
@@
 Let $\vec{p}=\langle p_{\eta}\mid\eta<\omega_{1}\rangle$ be a sequence of conditions in $\mathbb{P}$.
-By (the proof of) [5, Lemma 7.5], by thinning out $\vec{p}$ to some cofinal subsequence if necessary, we can assume that there is a free subgroup $A^{\prime}$ of $A$ that is pure in $A$ and such that $A_{p_{\eta}}\subseteq A^{\prime}$ for all $\eta<\omega_{1}$ (this is the only use of the full power of Chase’s condition in the proof).
+By (the proof of) [9, Lemma 7.5], by thinning out $\vec{p}$ to some cofinal subsequence if necessary, we can assume that there is a free subgroup $A^{\prime}$ of $A$ that is pure in $A$ and such that $A_{p_{\eta}}\subseteq A^{\prime}$ for all $\eta<\omega_{1}$ (this is the only use of the full power of Chase’s condition in the proof).
 Let $B^{\prime}$ be a basis for $A^{\prime}$.
@@
 $=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))$
-To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [13, Prop.
+To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [18, Prop.
 3.5 and Prop.
 3.2].
-The first equality follows from [20, p.
+The first equality follows from [26, p.
 13 (iii) (cf.
@@
 This leaves us only with the second.
-Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [20, p.
+Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [26, p.
 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$.
@@
 Let us check the details.
-Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [13, Prop.
+Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [18, Prop.
 2.18] and Proposition 12.3 below; this shows the third equality.
@@
 ∎
-Note that the lemma allows us to recast this paper’s essential background in uniformly classical terms: Shelah’s result was that the implication
+Note that the lemma allows us to recast this paper’s shaping background in uniformly classical terms: Shelah’s result was that the implication
 $\big{[}\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for }S=*\,\big{]}\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
@@
 is indeed a $\mathsf{ZFC}$ theorem.
-This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([16], see also [6, Proposition XI.4.4]), generalizing work of Specker [23], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
+This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([22], see also [10, Proposition XI.4.4]), generalizing work of Specker [29], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
 In fact, we may be more precise.
@@
 for every $S\in\mathsf{ED}$ of weight at least $\kappa$. ∎
-Lemma 11.1 also sheds additional light on the results of Section 10, both providing an alternative proof of a variation on the main result of that section, as well as showing that the existence of large cardinals provides a bound on the extent to which this result can be generalized.
+Lemma 11.1 also sheds additional light on the results of Section 10, entailing both an alternative proof of a variation of that section’s main result and limits, in the presence of sufficiently large cardinals, on the extent to which this result can be generalized.
 We first note that, by results in [11], it is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$.
@@
 In particular, it is consistent that, for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$.
-On the other hand, the existence of sufficiently large cardinals provides a limit on the extent to which the phenomena of Corollary 11.3 can occur.
+On the other hand, the existence of sufficiently large cardinals will bound the extent to which the phenomena of Corollary 11.3 can occur.
 In particular, if $\kappa$ is a strongly compact cardinal, then $\mathsf{CondAb}_{\kappa}$ is sufficiently rich to guarantee a positive answer to the appropriate interpretation of Whitehead’s problem.
@@
 By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free.
-By [6, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired.
-[^5] As [6, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
+By [10, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired.
+[^5] As [10, Corollary II.3.11] makes clear, we could have merely assumed that $\kappa$ is only $L_{\omega_{1}\omega}$-compact, a weaker but lesser-known hypothesis than the strong compactness of $\kappa$.
 Remark 11.5.
@@
 This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$.
-In particular, it is shown in [12, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
-12. Locally compact abelian groups
+In particular, it is shown in [16, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
+12. Locally compact and solid abelian groups
 In this section, we show that Theorem 1.2 extends from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones.
@@
 Proof.
-By the structure theorem for locally compact abelian groups (see [20, Thm.
-4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some $n\in\mathbb{N}$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
+By the structure theorem for locally compact abelian groups (see [26, Thm.
+4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some nonnegative integer $n$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
 (1) $0\to C\to B\to D\to 0$
-with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well.
-[^6] See [13, Props.
-4.9, 2.18].
-Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [20, p.
+with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [18, Props.
+4.9, 2.18]).
+Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [26, p.
 25], we have
-$(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0)\Leftrightarrow(\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0).$
-Consider the following portion of the long exact sequence deriving from equation (1):
+$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0\;\Leftrightarrow\;\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0.$
+Consider next the following portion of the long exact sequence deriving from equation (1):
 $\cdots\to 0\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{D},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})\to 0\to\cdots$
-(The rightmost 00 follows from [20, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [20, Prop. 4.2].) Hence
+(The rightmost 00 follows from [26, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [26, Prop. 4.2].) Hence
 $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0$
@@
 By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$.
-Also, the argument of [13, Prop.
-4.11] together with [20, Prop.
-4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([20, Prop.
+Also, the argument of [18, Prop.
+4.11] together with [26, Prop.
+4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([26, Prop.
 4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism.
-Again applying [20, Prop.
+Again applying [26, Prop.
 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$.
 Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$.
-As noted, by [15, Prop.
+As noted, by [21, Prop.
 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$.
@@
 As the condensed abelian group $\underline{\mathbb{R}}$ witnesses, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$ does not imply that $A$ is projective in $\mathsf{CondAb}$.
-To see this, recall that, as noted, the compact projective condensed abelian groups of the form $\mathbb{Z}[\underline{S}]$ with $S\in\mathsf{ED}$ generate $\mathsf{CondAb}$ ([3, p.
+To see this, recall that, as noted, the compact projective condensed abelian groups of the form $\mathbb{Z}[\underline{S}]$ with $S\in\mathsf{ED}$ generate $\mathsf{CondAb}$ ([7, p.
 16–17 and Proposition 3.5]); in particular, for any condensed abelian group $X$ there exists a set $I$, a collection $\{S_{i}\mid i\in I\}$ of spaces in $\mathsf{ED}$, and an epimorphism $r:\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]\to X$.
@@
 […]
+Theorem 12.1 opens onto broader questions about alternative interpretations of Whitehead’s problem in the condensed setting.
+These may be thought of as defined by two main parameters: the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$.
+After all, in retrospect, it is not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless).
+A natural next question, then, is how much further this $d$ may be pushed.
+By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$.
+Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction.
+Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [26, §5–7] and [25, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2.
+The solid setting, being “very much a ‘nonarchimedean’ notion” [26, footnote 10 on p.
+32], is also one in which $\underline{\mathbb{R}}$ does not appear.
+[^6] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [26, Corollary 6.1(iii)].
+Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [26, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
+Question 12.5.
+Let $A$ be a solid abelian group.
+Does $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0$ (or equivalently $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$) imply that $A$ is projective?
+We thank Peter Scholze for answering with a multifaceted no; the following observations are essentially his.
+First, for the solid abelian group $A=\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\underline{\mathbb{Z}}$ and any $n\geq 0$,
+$\underline{\mathrm{Ext}}^{n}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})(S)\cong\mathrm{lim}^{n}\,\mathbf{A}[C(S,\mathbb{Z})],$
+where $\mathbf{A}[H]$ denotes for any abelian group $H$ the inverse system indexed by functions $f:\mathbb{N}\to\mathbb{N}$, with terms $A_{f}:=\bigoplus_{i\in\mathbb{N}}H^{(f(i))}$ and natural projection maps $A_{g}\to A_{f}$ for any $f$ which is pointwise dominated by $g$.
+See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 33]), from which the consistency with the ZFC axioms of the statement that
+$\mathrm{lim}^{1}\,\mathbf{A}[H]=0\;\text{ but }\mathrm{lim}^{2}\,\mathbf{A}[H]\neq 0$
+for every nontrivial abelian group $H$, for example, is known.
+In fact this holds in any model of the Proper Forcing Axiom; whenever it does, $A$ will clearly witness a negative answer to Question 12.5.
+Left open in the preceding analysis is the possibility of a consistently affirmative answer to Question 12.5, since
+$\mathrm{lim}^{n}\,\mathbf{A}[H]=0\text{ for all }n>0\text{ and abelian groups }H$
+and hence
+$\underline{\mathrm{Ext}}^{n}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0\text{ for all }n>0$
+is also consistent with the ZFC axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
+Consider the surjection
+$t:\bigoplus_{\mathbb{N}^{\mathbb{N}}}\prod_{\mathbb{N}}\mathbb{Z}\to\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$
+defined as follows: fix $g:\mathbb{N}\to\mathbb{N}$ and an element $a$ of the $g^{\mathrm{th}}$ summand of the domain of $t$; clearly it will suffice to describe the effect of $t$ on such $a$.
+Here it is convenient to adopt the (elsewhere innocuous) convention that 00 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on.
+Equipping $\mathbb{Z}$ with the discrete topology and letting products and sums induce the product and coproduct topologies, respectively, we may view $t$ as a continuous mapping of topological abelian groups.
+Consider the condensed image $\underline{t}$ of $t$: observe that its codomain is $A$, its domain is solid, and that it remains an epimorphism (cf.
+[18, Prop.
+2.18]).
+Thus, to show that $A$ is not projective in $\mathsf{Solid}$, it will suffice to show that $\underline{t}$ does not split, and for this it will suffice to show that the mapping $\underline{t}(*)=t$ of discrete abelian groups described above possesses no right-inverse $s$.
+By [10, Cor.
+III.1.9], though, for any
+$s:\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}\to\bigoplus_{\mathbb{N}^{\mathbb{N}}}\prod_{\mathbb{N}}\mathbb{Z}$
+there exists an $m\in\mathbb{N}$ and finite $F\subseteq\mathbb{N}^{\mathbb{N}}$ such that the image of the restriction of $s$ to $\prod_{n\geq m}\bigoplus_{\mathbb{N}}\mathbb{Z}$ is contained within $\bigoplus_{f\in F}\prod_{\mathbb{N}}\mathbb{Z}$.
+Letting $k=\max\{f(m)\mid f\in F\}+1$ and writing $x$ for any element of $\prod_{n\geq m}\bigoplus_{\mathbb{N}}\mathbb{Z}$ which is nonzero at the $k^{\mathrm{th}}$ summand of the $m^{\mathrm{th}}$ factor, we must then have $ts(x)\neq x$.
+This shows that $t$ possesses no right-inverse, [^7] One can alternatively rule out such a right-inverse via the $\alpha=2$ instance of [10, Thm.
+X.2.3].
+and hence that $A$ is not projective, completing the argument.
+Even this doesn’t altogether settle the question, though, since it is also consistent with the ZFC axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [17, Thm.
+2].
+For a model-independent negative answer to Question 12.5, take the solid abelian group $B=\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}$ for any distinct prime numbers $p$ and $q$.
+As noted in Clausen’s Lecture 11 of [8],
+$\mathrm{Ext}^{2}_{\mathsf{CondAb}}(B,\underline{\mathbb{Z}})=\mathbb{Z}_{q}/\mathbb{Z}[1/p],$
+hence $B$ is not projective; it remains only to check that it is Whitehead. In fact, for the numerator
+[…]
+of $B$, we have
+[…]
+[…]
+$=(\underline{\mathbb{Z}[1/p]/\mathbb{Z}})[-1]$
+(see [19, Rmk. 2, p. 2] for the first equality), and for its denominator
+[…]
+we have
+$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}[1/q]},\underline{\mathbb{Z}})$ $=R\mathrm{lim}\;(R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})\xleftarrow{\times q}R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})\leftarrow\cdots)$
+$=R\mathrm{lim}\;(\underline{\mathbb{Z}}\xleftarrow{\times q}\underline{\mathbb{Z}}\xleftarrow{\times q}\cdots)$
+$=(\underline{\mathbb{Z}_{q}/\mathbb{Z}})[-1].$
+Hence, applying $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\,\cdot\,,\underline{\mathbb{Z}})$ to the short exact sequence
+$0\to\underline{\mathbb{Z}[1/q]}\to\underline{\mathbb{Z}_{p}}\to\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}\to 0$
+we find that
+$R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]},\underline{\mathbb{Z}})=(\underline{\mathbb{Z}_{q}}/\underline{\mathbb{Z}[1/p]})[-2]$
+and that $\underline{\mathrm{Ext}}^{1}(B,\underline{\mathbb{Z}})$, in particular, equals zero.
 13. Conclusion
-Though the project leading to this article began with a desire simply to understand the combinatorial phenomena underlying considerations of Whitehead’s problem in the setting of condensed abelian groups, it quickly expanded into the beginnings of a more general investigation into connections between condensed mathematics and forcing, an investigation that we hope to see further developed in the future.
+The project leading to this article began with a desire simply to understand the combinatorial phenomena underlying considerations of Whitehead’s problem in the setting of condensed abelian groups, but quickly expanded into the beginnings of a more general investigation into connections between condensed mathematics and forcing, an investigation that we hope to see further developed in the future.
 These connections have manifested herein in a number of ways, notably in
   1. (1)
-    the recognition that condensed objects can often be viewed as organized presentations of forcing names; and
+    the recognition that condensed objects can often be viewed as organized presentations of forcing names, and
   2. (2)
@@
 We feel that a resolution of this conjecture will shed light not just on the relationship between forcing and condensed mathematics, but also on considerations of Whitehead’s problem in the classical setting.
-Finally, there remain questions about alternative interpretations of Whitehead’s problem in the condensed setting.
-In Section 12, we initiated some further exploration of two main parameters in play in condensed interpretations of Whitehead’s problem, namely the subscript $s$ of the $\underline{\mathrm{Ext}}^{1}$ term, and the domain $d$ of $\underline{\mathrm{Ext}}^{1}_{s}(\cdot,\underline{\mathbb{Z}})$.
-After all, in retrospect, it’s not altogether surprising that ($s=\mathsf{CondAb}$)-valued invariants $\underline{\mathrm{Ext}}^{1}_{s}$ should be strong enough to settle the problem over the comparatively weaker domains of $d=\mathsf{Ab}$ or even $d=\mathsf{LCA}$ (we hope to have shown that it is altogether interesting nevertheless).
-A natural next question, then, is how much further this $d$ may be pushed.
-By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$.
-Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction.
-Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [20, §5–7] and [19, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2.
-Within it, in addition, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [20, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
-Question 13.2.
-Let $A$ be a solid abelian group.
-Does $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0$ (or equivalently $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$) imply that $A$ is projective?
-Part of the question is whether the answer will follow from the $\mathsf{ZFC}$ axioms alone.
-The solid setting, being “very much a ‘nonarchimedean’ notion” [20, footnote 10 on p.
-32], carries the additional virtue of circumventing the aforementioned issue of $\underline{\mathbb{R}}$.
-More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [20, Corollary 6.1(iii)]; in particular, $\underline{\mathbb{R}}$ is not solid.
-This leads to the following natural variation on Question 13.2: Suppose that $A$ is a condensed abelian group and $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(A,\underline{\mathbb{Z}})=0$.
-Must $A^{\blacksquare}$ be projective in the category of solid abelian groups?

### v5

@@
 One of the better-known independence results in general mathematics is Shelah’s solution to Whitehead’s problem of whether $\mathrm{Ext}^{1}(A,\mathbb{Z})=0$ implies that an abelian group $A$ is free.
-The point of departure for the present work is Clausen and Scholze’s proof that, in contrast, one natural interpretation of Whitehead’s problem within their recently-developed framework of condensed mathematics has an affirmative answer in $\mathsf{ZFC}$.
+The point of departure for the present work is Clausen and Scholze’s proof that, in contrast, one natural interpretation of Whitehead’s problem within their recently-developed framework of condensed mathematics has an affirmative answer in $\sf ZFC$.
 We record two alternative proofs of this result, as well as several original variations on it, both for their intrinsic interest and as a springboard for a broader study of the relations between condensed mathematics and set theoretic forcing.
 We show more particularly how the condensation $\underline{X}$ of any locally compact Hausdorff space $X$ may be viewed as an organized presentation of the forcing names for the points of canonical interpretations of $X$ in all possible set-forcing extensions of the universe, and we argue our main result by way of this fact.
-We also show that when interpreted within the category of light condensed abelian groups, Whitehead’s problem is again independent of the $\mathsf{ZFC}$ axioms.
+We also show that when interpreted within the category of light condensed abelian groups, Whitehead’s problem is again independent of the $\sf ZFC$ axioms.
 In fact we show that it is consistent that Whitehead’s problem has a negative solution within the category of $\kappa$-condensed abelian groups for every uncountable cardinal $\kappa$, but that this scenario, in turn, is inconsistent with the existence of a strongly compact cardinal.
@@
 03E35, 03E40, 03E05, 13C10, 20J05
-The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S.
-A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023.
-We thank the Fields Institute for their support and hospitality.
-The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023.
-We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks, and to thank Peter Scholze for valuable remarks on the paper as well.
+The first author was supported by María Zambrano and Marie Skłodowska Curie (project 101110452: CatT) Fellowships at the University of Barcelona; the second author was supported by GAČR project 23-04683S and the Academy of Sciences of the Czech Republic (RVO 67985840); the third author was supported by GAČR project 23-05148S
 1. Introduction
@@
 Our motivation for doing this is twofold.
-First, we seek to excavate some of the combinatorial and set theoretic ideas underlying Theorem 1.2 and thereby to improve our understanding of the mathematics surrounding Whitehead’s problem, in both general and condensed contexts.
-To this end, we prove the following theorem.
+First, we seek to excavate some of the combinatorial and set theoretic ideas underlying Theorem 1.2 and thereby to improve our understanding of the mathematics surrounding Whitehead’s problem, in both classical and condensed contexts.
+To this end, we prove the following theorem, item (2) of which strengthens Theorem 1.2.
 Theorem 1.3.
@@
   1. (1)
-    $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(2^{\kappa})\neq 0$;
-  2. (2)
-    in $V^{\mathbb{B}}$, $A$ is not Whitehead;
+    in $V^{\mathbb{B}}$, $A$ is not Whitehead; and
   2. (2)
     $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$.
-Second, using Theorem 1.3, particularly clauses (2) and (3) thereof, as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing.
+Second, using Theorem 1.3 as a starting point, we begin to investigate a more general connection between condensed mathematics and set theoretic forcing.
 We isolate a precise sense in which, given a locally compact Hausdorff space $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
@@
   2. (2)
-    It is consistent with the axioms of $\mathsf{ZFC}$ that for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0$.
+    It is consistent with the axioms of $\sf ZFC$ that for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\kappa}}(\underline{A},\underline{\mathbb{Z}})=0$.
   3. (3)
@@
     then $A$ is free.
-Second, we extend Theorem 1.2 from the class of discrete abelian groups to the class of locally compact abelian groups and show that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})=0$ for a locally compact abelian group $A$ if and only if $A$ is projective in the category of locally compact abelian groups.
+Second, we extend Theorem 1.2 from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones.
+In particular, we prove the following theorem:
+Theorem 1.5.
+Let $A$ be an object of the category $\mathsf{LCA}$ of locally compact abelian groups. Then
+$\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$
+if and only if $A$ is projective in $\mathsf{LCA}$.
 We record as well several observations by Peter Scholze showing that Theorem 1.2 does not extend to the category of solid abelian groups.
@@
 In Section 4, we prove a slight strengthening of Stein’s aforementioned classical result [33] that all countable Whitehead groups are free.
-In Section 5, we prove clause (1) of Theorem 1.3, and, in Section 6, we adapt that proof to establish clause (2) of Theorem 1.3.
+In Section 5, we prove a precursor to Theorem 1.3 concerning the space $2^{\kappa}$, and, in Section 6, we adapt that proof to establish clause (1) of Theorem 1.3.
 From here we present two paths to clause (2) of Theorem 1.3.
@@
 In Section 11, we record an algebraic lemma converting the groups $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)$ into a more tractable, and suggestive, form.
-This lemma yields an alternative proof of clauses (1) and (3) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4.
-In Section 12 we prove an extension of Theorem 1.2 to the class of locally compact abelian groups, and record impediments to its extension to the class of solid abelian groups.
+This lemma yields an alternative proof of clause (2) of Theorem 1.3 (and hence of Theorem 1.2), as well as clauses (2) and (3) of Theorem 1.4.
+In Section 12 we prove Theorem 1.5 and record impediments to the extension of Theorem 1.2 to the class of solid abelian groups.
 We close out our work in Section 13 with a few concluding remarks and one main open question.
@@
 If $f$ is a function and $X\subseteq\mathrm{dom}(f)$, then $f[X]$ denotes the image of $X$ under $f$, i.e., $f[X]=\{f(x)\mid x\in X\}$.
-Given a Boolean algebra $\mathbb{B}$, we let $S(\mathbb{B})$ denote the Stone space of $\mathbb{B}$.
+Given a Boolean algebra $\mathbb{B}$, we let $\mathbb{B}^{+}$ denote the set of all nonzero elements of $\mathbb{B}$ and $S(\mathbb{B})$ denote the Stone space of $\mathbb{B}$.
 Concretely, $S(\mathbb{B})$ is the space of ultrafilters on $\mathbb{B}$, with basic clopen sets given by
-$N_{b}:=\{U\in S(\mathbb{B})\mid b\in U\}$
-for $b\in\mathbb{B}$.
+$N_{b}:=\{\mathcal{U}\in S(\mathbb{B})\mid b\in\mathcal{U}\}$
+for $b\in\mathbb{B}^{+}$.
 2. Condensed abelian groups
@@
     $T(S_{0}\sqcup S_{1})\rightarrow T(S_{0})\times T(S_{1})$
-    is a bijection.
+    induced by the inclusion maps of $S_{0}$ and $S_{1}$ into $S_{0}\sqcup S_{1}$ is a bijection.
   3. (3)
@@
     $T(S)\rightarrow\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$
-    is a bijection.
+    induced by $f$ is a bijection.
+[^2] In slightly more detail, $f$ induces a map from $T(S)$ to $T(S^{\prime})$ whose range is readily seen to be contained in the set $\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$.
+This condition then asserts that this map is a bijection between $T(S)$ and $\{x\in T(S^{\prime})\mid p_{0}^{*}(x)=p_{1}^{*}(x)\}$.
 […]
@@
 When restricted to the full subcategory of $\mathsf{Haus}$ spanned by the compactly generated Hausdorff spaces, it is a fully faithful embedding (cf.
-[26, Proposition 1.7]).
+[28, Proposition 1.7]).
 Similarly, given a topological abelian group $A$, one can define a condensed abelian group $\underline{A}$ by letting $\underline{A}(S)=C(S,A)$ (considered as an abelian group with pointwise addition) for all $S\in\mathsf{Prof}$.
@@
 For each uncountable cardinal $\kappa$, there is an obvious forgetful functor from $\mathsf{CondAb}_{\kappa}$ to $\mathsf{Cond}_{\kappa}$, and this functor admits a left adjoint that sends a $\kappa$-condensed set $T$ to a $\kappa$-condensed abelian group denoted $\mathbb{Z}[T]$.
-Concretely, as noted in [26, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{ED}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{ED}_{\kappa}$ to the free group on $T(S)$.
+Concretely, as noted in [28, §2], $\mathbb{Z}[T]$ is the sheafification of the functor from $\mathsf{Prof}_{\kappa}$ to $\mathsf{Ab}$ that sends each $S\in\mathsf{Prof}_{\kappa}$ to the free group on $T(S)$.
 Moreover, $\{\mathbb{Z}[\underline{S}]\mid S\in\mathsf{ED}_{\kappa}\}$ is, for any strong limit cardinal $\kappa$, a set of compact projective generators for $\mathsf{CondAb}_{\kappa}$.
@@
 Given any abelian group $A$, a free resolution of $A$ is a short exact sequence of the form
-[…]
+$0\rightarrow K\xrightarrow{\iota}F\xrightarrow{\pi}A\rightarrow 0$
 in which both $F$ and $K$ are free abelian groups.
@@
 We note that every infinite abelian group $A$ has a canonical free resolution, in which $F$ is the free abelian group generated by the underlying set of $A$ and $K$ is the kernel of the induced surjection from $F$ onto $A$.
-The elements of $K$ are therefore the formal linear combinations of elements of $F$ that evaluate to 00 in $A$.
+The elements of $K$ are therefore the formal linear combinations of elements of $F$ that evaluate to 0 in $A$.
 By applying the functor $\mathrm{Hom}(\cdot,\mathbb{Z})$ to a free resolution of $A$ and using the fact that $F$ and $K$ are free, and therefore Whitehead, we obtain an exact sequence of the form
@@
 If $A$ and $B$ are sufficiently nice topological groups, then $\underline{\mathrm{Hom}}(\underline{A},\underline{B})$ has a particularly nice description:
-Proposition 3.1 (Clausen–Scholze, [26, Proposition 4.2]).
+Proposition 3.1 (Clausen–Scholze, [28, Proposition 4.2]).
 Suppose that $A$ and $B$ are Hausdorff topological abelian groups and $A$ is compactly generated. Then there is a natural isomorphism of condensed abelian groups
@@
 Proof.
-As noted in [26, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
+As noted in [28, proof of Corollary 4.8] (and as follows from items (ii) and (iii) of its page 13),
 $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(T,M)(S)=R\mathrm{Hom}_{\mathsf{CondAb}}(T\otimes\mathbb{Z}[S],M).$
@@
 When $S=\ast$, our general claim follows from the facts that $\mathbb{Z}[*]=\underline{\mathbb{Z}}$ and that $\underline{\mathbb{Z}}$ is the unit object for the condensed tensor product $\otimes$.
-To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [18, Props.
+To see the assertion for discrete abelian groups, recall that their category fully and faithfully embeds into $\mathsf{CondAb}$, and observe by [19, Props.
 2.18 and 2.19] that if $\cdots\to P^{1}\to P^{0}\to A$ is a projective (hence free) resolution of $A$ then $\cdots\to\underline{P}^{1}\to\underline{P}^{0}\to\underline{A}$ is a projective resolution of $\underline{A}$.
@@
 Observe next that, in close analogy with the classical situation, if $F$ is a free (discrete) abelian group, then $\underline{F}$ is Whitehead in $\mathsf{CondAb}$.
-To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows then that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$, from [18, Proposition 2.19].
-By [18, Remark 3.10 and Lemma 4.2], we then have
+To see this, let $\kappa$ be a cardinal, and let $F=\bigoplus_{\kappa}\mathbb{Z}$ be the free group on $\kappa$ generators; it follows from [19, Proposition 2.19] that $\underline{F}\cong\bigoplus_{\kappa}\underline{\mathbb{Z}}$, where the direct sum on the right is computed in $\mathsf{CondAb}$.
+By [19, Remark 3.10 and Lemma 4.2], we then have
 $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=\prod_{\kappa}R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\underline{\mathbb{Z}},\underline{\mathbb{Z}})=\prod_{\kappa}\underline{\mathbb{Z}}[0],$
-where ¯$\underline{\mathbb{Z}}[0]$ denotes the chain complex that is $\underline{\mathbb{Z}}$ in degree 00 and 00 elsewhere.
+where $\underline{\mathbb{Z}}[0]$ denotes the chain complex that is $\underline{\mathbb{Z}}$ in degree 0 and 0 elsewhere.
 In particular, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{F},\underline{\mathbb{Z}})=0$.
@@
 In the coming sections, we will give a more concrete, combinatorial proof of Theorem 1.2.
-In particular, given a nonfree discrete abelian group $A$, we will identify a natural profinite set $S$ such that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$.
+In particular, given a nonfree discrete abelian group $A$, we will identify a natural $S\in\mathsf{ED}$ such that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\neq 0$.
 The process by which this will be done is analogous to that described in the classical setting at the beginning of this section, so in the interest of symmetry, we end this section by giving an outline of the structure of the argument.
 Suppose that $A$ is a nonfree discrete abelian group, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$.
-Applying $\underline{\mathrm{Hom}}(\cdot,\underline{\mathbb{Z}})$ to the image of this short exact sequence in $\mathsf{CondAb}$ yields the exact sequence
-$0\rightarrow\underline{\mathrm{Hom}}(\underline{A},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{F},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{K},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})\rightarrow 0$
-Now, evaluating this sequence at a profinite set $S$ and recalling Proposition 3.1, we obtain
-$0\rightarrow\underline{\mathrm{Hom}(A,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\rightarrow 0$
+Using the discreteness of the groups under consideration, it is straightforward to observe that the induced sequence $0\rightarrow\underline{K}\rightarrow\underline{F}\rightarrow\underline{A}\rightarrow 0$ is exact in $\mathsf{CondAb}$.
+Applying $\underline{\mathrm{Hom}}(\cdot,\underline{\mathbb{Z}})$ to this short exact sequence yields the exact sequence
+$0\rightarrow\underline{\mathrm{Hom}}(\underline{A},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{F},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Hom}}(\underline{K},\underline{\mathbb{Z}})\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})\rightarrow 0.$
+Given an $S\in\mathsf{ED}$, the functor from $\mathsf{CondAb}$ to $\mathsf{Ab}$ sending $T\in\mathsf{CondAb}$ to $T(S)\in\mathsf{Ab}$ is exact (cf.
+[16, Lemma 6.1.3]).
+Thus, evaluating the above sequence at a space $S\in\mathsf{ED}$ and recalling Proposition 3.1, we obtain
+$0\rightarrow\underline{\mathrm{Hom}(A,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)\rightarrow 0.$
 It follows that $\underline{\mathrm{Ext}}^{1}(\underline{A},\underline{\mathbb{Z}})(S)=0$ if and only if the map
@@
 Moreover, the map $\underline{\mathrm{Hom}(F,\mathbb{Z})}(S)\rightarrow\underline{\mathrm{Hom}(K,\mathbb{Z})}(S)$ sends a continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ to the map $\phi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ defined by letting $\phi(s)=\psi(s)\restriction K$ for all $s\in S$.
-Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a profinite set $S$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$ (i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$).
+Therefore, to show that $\underline{A}$ is not Whitehead, it will suffice to find a space $S\in\mathsf{ED}$ and a continuous $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ that does not “lift pointwise” to a map from $S$ to $\mathrm{Hom}(F,\mathbb{Z})$, i.e., for which there is no continuous $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ for which $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$.
 The construction of such an $S$ and $\varphi$ will be the subject of the next sections.
+For expository reasons, and because it more closely derives from classical arguments around Whitehead’s problem, we begin by finding $S$ and $\varphi$ as above but for which $S$ is merely profinite and not extremally disconnected.
+We then show how to improve these arguments to obtain $S\in\mathsf{ED}$, as desired.
 4. Countable groups
@@
 Suppose that $A$ is a countable, torsion-free, nonfree abelian group,
-[…]
+$0\rightarrow K\xrightarrow{\subseteq}F\xrightarrow{\pi}A\rightarrow 0$
 is a free resolution of $A$, and $B_{K}$ is a basis for $K$. Then there is a homomorphism $\varphi:K\rightarrow\mathbb{Z}$ such that
@@
 We will do this by constructing a strictly increasing sequence $\langle n_{i}\mid i<\omega\rangle$ of natural numbers and, for each $i<\omega$, a homomorphism $\varphi_{i}:K_{n_{i}}\rightarrow\mathbb{Z}$ in such a way that $\varphi_{i}[\{t_{k}\mid k<n_{i}\}]\subseteq\{0,1\}$ and, for all $i<j<\omega$, $\varphi_{j}\restriction K_{n_{i}}=\varphi_{i}$.
+In the process, we will also define a $\subseteq$-increasing sequence $\langle Y_{n}\mid n<\omega\rangle$ of finite subsets of $B_{F}$ such that, for all $i<\omega$, letting $F_{i}$ denote the subgroup of $F$ generated by $Y_{i}$, we have $K_{n_{i}}\subseteq F_{i}$.
 By Theorem 4.1, since $A$ is countable, torsion-free, and nonfree, there is a finitely generated subgroup $X_{0}$ of $A$ such that its pure closure $X$ is not finitely generated.
@@
 Moreover, since all of the coefficients $b_{k}$ for $k\in[n_{i},n_{i+1})$ are divisible by $j_{i}$, it follows that $s_{i}^{*}$ is divisible by $j_{i}$ (in $K$).
-Since $r=j_{i}z_{i}-s_{i}^{*}$, this means that there exists $r^{*}\in F_{i}$ such that $r=j_{i}r^{*}$.
+Since $r=j_{i}z_{i}-s_{i}^{*}$ is in $F_{i}$ and $F_{i}$ is generated by a subset of $B_{F}$, this means that there exists $r^{*}\in F_{i}$ such that $r=j_{i}r^{*}$.
 We get $s_{i}^{*}=j_{i}(z_{i}-r^{*})\in K$ which implies $z_{i}-r^{*}\in K$ using the fact that $A$ is torsion-free.
@@
 Let
-[…]
+$0\rightarrow K\xrightarrow{\subseteq}F\xrightarrow{\pi}A\rightarrow 0$
 be a short exact sequence of groups where $F$ is torsion-free and $K$ is free with a basis $B_{K}$.
@@
 The reason is that any given countable torsion-free group $A$, being the union of a countable chain of finitely generated free subgroups, is isomorphic to a (pure) subgroup of $\mathbb{Z}^{\omega}/\mathbb{Z}^{(\omega)}$, cf.
-[34, Proposition 3.1] or [24, Theorem 3.3.2].
+[26, Proposition 3.1] or [25, Theorem 3.3.2].
 Here, $\mathbb{Z}^{(\omega)}$ denotes the direct sum of countably many copies of $\mathbb{Z}$, i.e., the subgroup of $\mathbb{Z}^{\omega}$ consisting of elements with finite support.
@@
 Note that $2^{\kappa}$ is a profinite set.
-Our next result establishes item (1) of Theorem 1.3.
 Theorem 5.1.
@@
 Moreover, $\varphi$ does not lift pointwise to a continuous $\psi:2^{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$, as pointwise restriction of such a $\psi$ to $F_{0}$ would induce a continuous pointwise extension $\psi_{0}:2^{\kappa}\rightarrow\mathrm{Hom}(F_{0},\mathbb{Z})$ of $\varphi_{0}$.
-We can therefore assume without loss of generality that $A$ is almost free.
 Thus, for the rest of the proof we will assume that $A$ is almost free and $|F|=\kappa$.
@@
   2. (2)
-    for all $\eta<\kappa$ for which $\gamma_{\eta}\in E$, the set $A_{\gamma_{\eta+1}}/A_{\gamma}$ is not free.
+    for all $\eta<\kappa$ for which $\gamma_{\eta}\in E$, the set $A_{\gamma_{\eta+1}}/A_{\gamma_{\eta}}$ is not free.
 Therefore, by thinning out $\vec{M}$ to only include those models indexed by elements of $D$ and then reindexing, we may assume that the set
@@
 Let
-$t^{-}:=\sum_{z\in I^{-}}a_{z}z\text{\ \ and \ }t^{+}:=\sum_{z\in I^{+}}a_{z}z.$
+[…]
 The following observations are immediate, but crucial:
@@
 Suppose that $A$ is a nonfree abelian group and $\kappa=|A|$. Then, in $V^{\mathrm{Add}(\omega,\kappa)}$, $A$ is not Whitehead.
-Note that this establishes item (2) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
+Note that this establishes clause (1) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
 Lemma 6.2.
@@
 Indeed, if $A_{0}$ is a nonfree subgroup of $A$ and $|A_{0}|<\kappa$, then we can apply the inductive hypothesis to conclude that $A_{0}$, and hence also $A$, remains nonfree in $V^{\mathbb{P}}$.
-Thus, assume that $A$ is almost free, and hence $\kappa$ is regular.
+Thus, assume that $A$ is almost free, and hence, by Shelah’s aforementioned singular compactness theorem, $\kappa$ is regular.
 Since $A$ is nonfree, we can fix a $\subseteq$-increasing, continuous sequence $\vec{A}:=\langle A_{\alpha}\mid\alpha<\kappa\rangle$ such that
@@
   2. (2)
-    in every ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$, there is no $\psi\in\mathrm{Hom}(F,\mathbb{Z})$ extending $\varphi$.
+    if $\mathbb{Q}\in V$ is a ccc forcing notion, then, in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$, there is no $\psi\in\mathrm{Hom}(F,\mathbb{Z})$ extending $\varphi$.
 Proof.
@@
 To see this, let $N$, $A_{0}$, $F_{0}$, $K_{0}$, $B_{F,0}$, and $B_{K,0}$ be as in the analogous argument in the earlier proof.
-Suppose that we are able to obtain, in $V^{\mathrm{Add}(\omega,\kappa)}$, a $\varphi_{0}\in\mathrm{Hom}(K_{0},\mathbb{Z})$ such that $\varphi_{0}[B_{K,0}]\subseteq\{0,1\}$ and $\varphi_{0}$ cannot be extended to a map $\psi_{0}\in\mathrm{Hom}(F_{0},\mathbb{Z})$ in any ccc forcing extension of $V^{\mathrm{Add}(\omega,\kappa)}$.
+Suppose that we are able to obtain, in $V^{\mathrm{Add}(\omega,\kappa)}$, a $\varphi_{0}\in\mathrm{Hom}(K_{0},\mathbb{Z})$ such that $\varphi_{0}[B_{K,0}]\subseteq\{0,1\}$ and, for every ccc $\mathbb{Q}\in V$, $\varphi_{0}$ cannot be extended to a map $\psi_{0}\in\mathrm{Hom}(F_{0},\mathbb{Z})$ in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$.
 Extend $\varphi_{0}$ to a map $\varphi\in\mathrm{Hom}(K,\mathbb{Z})$ by letting $\varphi(t)=0$ for all $t\in B_{K}\setminus B_{K,0}$.
 Then $\varphi\in V^{\mathrm{Add}(\omega,\kappa)}$.
-Moreover, if $\dot{\mathbb{P}}$ is an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, then $\varphi$ does not extend in $V^{\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}}$ to an element of $\mathrm{Hom}(F,\mathbb{Z})$, as the restriction of such an extension to $F_{0}$ would extend $\varphi_{0}$, contradicting our choice of $\varphi_{0}$.
+Moreover, if $\mathbb{Q}\in V$ is a ccc forcing notion, then $\varphi$ does not extend in $V^{\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}}$ to an element of $\mathrm{Hom}(F,\mathbb{Z})$, as the restriction of such an extension to $F_{0}$ would extend $\varphi_{0}$, contradicting our choice of $\varphi_{0}$.
 Assume now that $A$ is almost free and $|F|=\kappa$. For $\alpha<\kappa$, let $M_{\alpha}$, $\delta_{\alpha}$, $A_{\alpha}$, $F_{\alpha}$, and $K_{\alpha}$ be as in the proof of Theorem 5.1. Let
@@
 For each $\alpha\in S$, let $\kappa_{\alpha}$, $A_{\alpha}^{*}$, $F_{\alpha}^{*}$, $K_{\alpha}^{*}$, $B^{*}_{F,\alpha}$, and $B^{*}_{K,\alpha}$ be as in the proof of Theorem 5.1.
-Note that $\mathrm{Add}(\omega,\kappa)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,\kappa)$, and hence we can assume that we are in fact working in an extension $V_{1}$ of the ground model $V$ by (the first copy of) $\mathrm{Add}(\omega,\kappa)$.
+Note that $\mathrm{Add}(\omega,\kappa)\cong\mathrm{Add}(\omega,\kappa)\times\mathrm{Add}(\omega,\kappa)$; move to an extension $V_{1}$ of the ground model $V$ by (the first copy of) $\mathrm{Add}(\omega,\kappa)$.
 In particular, by the inductive hypothesis, we can assume that, for every $\alpha\in S$, we have a $\varphi_{\alpha}\in\mathrm{Hom}(K_{\alpha}^{*},\mathbb{Z})$ such that
   • $\varphi_{\alpha}[\{t+K_{\alpha}\mid t\in B^{*}_{K,\alpha}\}]\subseteq\{0,1\}$;
-  • in every ccc forcing extension of $V_{1}$, there is no $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$.
+  • for every ccc $\mathbb{Q}\in V$, in $V_{1}^{\mathbb{Q}}$ there is no $\psi_{\alpha}\in\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$.
 We now describe a method for producing elements of $\mathrm{Hom}(K,\mathbb{Z})$.
@@
 We claim that $\dot{\varphi}$ is forced to be as desired.
-To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
+To see this, let $\mathbb{Q}\in V$ be a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}\in V_{1}$ is an $\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
 […]
   • $\delta_{\alpha}=\alpha$;
-  • for all $\gamma<\alpha$, we have $\mathrm{dom}(p_{\gamma},q_{\gamma})\subseteq\alpha$.
+  • for all $\gamma<\alpha$, we have $\mathrm{dom}(p_{\gamma})\subseteq\alpha$.
 For each $p\in\mathrm{Add}(\omega,\kappa)$, define $\hat{p}\in\mathrm{Add}(\omega,\kappa)$ by letting $\mathrm{dom}(\hat{p})=\mathrm{dom}(p)$ and, for all $\beta\in\mathrm{dom}(p)$, letting
 $\hat{p}(\beta)=\begin{cases}p(\beta)&\text{if }\beta\neq\alpha\\ 1-p(\beta)&\text{if }\beta=\alpha.\end{cases}$
-Since the map $p\mapsto\hat{p}$ is an automorphism of $\mathrm{Add}(\omega,\kappa)$, the filter $\hat{G}=\{\hat{p}\mid p\in G\}$ is generic over $V_{1}$, and the function $\hat{g}=\bigcup\hat{G}$ agrees with $g$ everywhere except $\alpha$, where it takes the opposite value.
-Without loss of generality, assume that $g(\alpha)=0$ and $\hat{g}(\alpha)=1$.
-For ease of comprehension, we then denote $(G,g)$ by $(G_{0},g_{0})$ and $(\hat{G},\hat{g})$ by $(G_{1},g_{1})$.
-Note also that $V[G_{0}]=V[G_{1}]$, and hence $H$ is $\mathbb{P}$-generic over both.
-For $i<2$, let $\varphi_{i}:=\varphi(g_{i})=\dot{\varphi}^{G_{i}}$, and let $\psi_{i}:=\dot{\psi}^{G_{i}\ast H}$.
+[…]
 By our assumptions about $\dot{\psi}$, we know that, for each $i<2$, $\psi_{i}\in\mathrm{Hom}(F,\mathbb{Z})$ extends $\varphi_{i}$.
@@
 The proof of this is exactly the same as that of the analogous fact in the proof of Theorem 5.1, so we leave it to the reader.
-But then, in $V_{1}[G\ast H]$, we have an element of $\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$, contradicting our choice of $\varphi_{\alpha}$ and the fact that $V_{1}[G\ast H]$ is a ccc extension of $V_{1}$.
+But then, in $V_{1}[G\times H]$, we have an element of $\mathrm{Hom}(F_{\alpha}^{*},\mathbb{Z})$ extending $\varphi_{\alpha}$, contradicting our choice of $\varphi_{\alpha}$ and the fact that $V_{1}[G\times H]$ is an extension of $V_{1}$ by $\mathrm{Add}(\omega,\kappa)\times\mathbb{Q}$, which is a ccc forcing notion in $V$.
 ∎
@@
 Recall that a clopen basis for $S_{\kappa}$ is given by all sets of the form
-$N_{b}:=\{U\in S_{\kappa}\mid b\in U\}$
-for $b\in\mathbb{B}_{\kappa}$.
+$N_{b}:=\{\mathcal{U}\in S_{\kappa}\mid b\in\mathcal{U}\}$
+for $b\in\mathbb{B}^{+}_{\kappa}$.
 Definition 7.1.
@@
     Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$.
-Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(U)(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$.
-Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
-It is routine to verify that $\dot{f}_{\varphi}$ as defined in Definition 7.1(1) is indeed a $\mathbb{B}_{\kappa}$ name for a function from $X$ to $Y$ and that $\varphi_{\dot{f}}$ as defined in Definition 7.1(2) is a continuous function from $S_{\kappa}$ to $Y^{X}$.
+Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $\mathcal{U}\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(\mathcal{U})(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in\mathcal{U}$.
+Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $\mathcal{U}$.
+It is routine to verify that $\dot{f}_{\varphi}$ as defined in Definition 7.1(1) is indeed a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$ and that $\varphi_{\dot{f}}$ as defined in Definition 7.1(2) is a continuous function from $S_{\kappa}$ to $Y^{X}$.
 (These facts will also follow from more general arguments appearing below in Section 8.) This translation, together with Theorem 6.3, allows us to replace the space $2^{\kappa}$ in Theorem 5.1 with the Stone space $S_{\kappa}$.
@@
 Let $\dot{f}$ be the $\mathbb{B}_{\kappa}$-name for the restriction of $\tau$ to $B_{K}$; $\dot{f}$ is thus a name for a function from $B_{K}$ to $2$.
-Let $\varphi_{\dot{f}}$ be the continuous function from $S_{\kappa}$ to $2^{B_{K}}$ given Definition 7.1(2), and let $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be defined by letting $\varphi(U)$ be the linear extension of $\varphi_{\dot{f}}(U)$ to all of $K$.
+Let $\varphi_{\dot{f}}$ be the continuous function from $S_{\kappa}$ to $2^{B_{K}}$ given by Definition 7.1(2), and let $\varphi:S_{\kappa}\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be defined by letting $\varphi(\mathcal{U})$ be the linear extension of $\varphi_{\dot{f}}(\mathcal{U})$ to all of $K$.
 Then $\varphi$ is also a continuous function.
 We claim that $\varphi$ is as desired.
-We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$.
+We clearly have $\varphi(\mathcal{U})[B_{K}]\subseteq\{0,1\}$ for all $\mathcal{U}\in S_{\kappa}$.
 Suppose for the sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise.
@@
 By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$.
-Let $U\in S_{\kappa}$ be such that $b\in U$.
-Then by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(y)=k_{y}$, $\psi(U)(z)=k_{z}$, and $\psi(U)(y+z)=k_{yz}$, contradicting the fact that $\psi(U)\in\mathrm{Hom}(F,\mathbb{Z})$.
+Let $\mathcal{U}\in S_{\kappa}$ be such that $b\in\mathcal{U}$.
+Then by the definition of $\dot{f}_{\psi}$, we have $\psi(\mathcal{U})(y)=k_{y}$, $\psi(\mathcal{U})(z)=k_{z}$, and $\psi(\mathcal{U})(y+z)=k_{yz}$, contradicting the fact that $\psi(\mathcal{U})\in\mathrm{Hom}(F,\mathbb{Z})$.
 ∎
@@
 By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$.
-Let $U\in S_{\kappa}$ be such that $b\in U$.
-Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(U)(t)=\varphi(U)(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise.
+Let $\mathcal{U}\in S_{\kappa}$ be such that $b\in\mathcal{U}$.
+Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(\mathcal{U})(t)=\varphi(\mathcal{U})(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(\mathcal{U})(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise.
 ∎
@@
 In this section and the next, we begin a general investigation into connections between condensed mathematics and forcing, the results of which will yield clause (2) of Theorem 1.3 as a corollary.
-Suppose that $(X,\tau)$ is a compact Hausdorff space and $W$ is an outer model of $V$.
-Then there is a canonical way to interpret $(X,\tau)$ as a compact Hausdorff space $(\hat{X},\hat{\tau})$ in $W$.
-First, let $\tau_{c}$ denote the collection of all closed subsets of $X$ (in $V$).
-In $W$, let $\hat{X}$ denote the collection of all maximal filters on $\tau_{c}\setminus\{\emptyset\}$.
-In other words, $\hat{X}$ is the collection of all maximal collections of elements of $\tau_{c}$ with the finite intersection property.
-There is a natural map $\pi:X\rightarrow\hat{X}$ defined by letting $\pi(x):=\{D\in\tau_{c}\mid x\in D\}$ for all $x\in X$.
-We also define $\pi:\tau\rightarrow\mathscr{P}(\hat{X})$ by letting $\pi(U):=\{F\in\hat{X}\mid(X\setminus U)\notin F\}$.
-Then let $\hat{\tau}$ be the topology on $\hat{X}$ generated by $\pi[\tau]$.
+[…]
+[…]
 Remark 8.1.
-The space $(\hat{X},\hat{\tau})$ identified above, together with the associated maps $\pi:X\rightarrow\hat{X}$ and $\pi:\tau\rightarrow\hat{\tau}$ is precisely the interpretation of $(X,\tau)$ in $W$, in the sense of [36].
-In particular, this implies the following properties, which can also readily be directly verified:
-  • $(\hat{X},\hat{\tau})$ is compact in $W$;
+[…]
+[…]
   • $\pi[\tau]$ is a basis for $\hat{\tau}$;
   • for all $x\in X$ and $U\in\tau$, we have $x\in U$ if and only if $\pi(x)\in\pi(U)$.
-Now suppose that $\mathbb{B}$ is a complete Boolean algebra and $S=S(\mathbb{B})$ is its Stone space.
-We now show how to identify the space $(\hat{X},\hat{\tau})$, as computed in the forcing extension by $\mathbb{B}$, with an appropriate quotient of the space $C(S,X)^{V}$, equipped with the compact-open topology.
-In fact, we will prove something more general.
+[…]
 We first establish a correspondence between elements of $C(S,X)$ and $\mathbb{B}$-names for points in $\hat{X}$.
@@
 In the forward direction, fix a function $f\in C(S,X)$, and define a $\mathbb{B}$-name $\dot{F}_{f}$ for a subset of $\tau_{c}$ as follows.
-For every $D\in\tau_{c}$ and every $b\in\mathbb{B}$, put $(\check{D},b)$ in $\dot{F}_{f}$ if and only if $N_{b}\subseteq f^{-1}[D]$.
+For every $D\in\tau_{c}$ and every $b\in\mathbb{B}^{+}$, put $(\check{D},b)$ in $\dot{F}_{f}$ if and only if $N_{b}\subseteq f^{-1}[D]$.
 Lemma 8.2.
@@
 To show that it is forced to be maximal, fix an arbitrary $C\in\tau_{c}$.
-It suffices to show that the following set is dense in $\mathbb{B}$:
-$E_{C}:=\{b\in\mathbb{B}\mid N_{b}\subseteq f^{-1}[C]\text{ {or} }\exists D\in\tau_{c}\,[D\cap C=\emptyset\wedge N_{b}\subseteq f^{-1}[D]]\}.$
-To this end, fix an arbitrary $b_{0}\in\mathbb{B}$, and assume that $N_{b_{0}}\not\subseteq f^{-1}[C]$.
+It suffices to show that the following set is dense in $\mathbb{B}^{+}$:
+$E_{C}:=\{b\in\mathbb{B}^{+}\mid N_{b}\subseteq f^{-1}[C]\text{ {or} }\exists D\in\tau_{c}\,[D\cap C=\emptyset\wedge N_{b}\subseteq f^{-1}[D]]\}.$
+To this end, fix an arbitrary $b_{0}\in\mathbb{B}^{+}$, and assume that $N_{b_{0}}\not\subseteq f^{-1}[C]$.
 Fix $\mathcal{U}\in N_{b_{0}}$ such that $f(\mathcal{U})\in X\setminus C$.
@@
 Thus, $\dot{F}_{\mathcal{U}}$ has the finite intersection property; since $X$ is compact, it follows that $\bigcap\dot{F}_{\mathcal{U}}$ is nonempty.
-Now suppose for sake of contradiction that $x_{0}$ and $x_{1}$ are distinct elements of $\dot{F}_{\mathcal{U}}$.
+Now suppose for the sake of contradiction that $x_{0}$ and $x_{1}$ are distinct elements of $\bigcap\dot{F}_{\mathcal{U}}$.
 Since $X$ is Hausdorff, we can find disjoint open neighborhoods $U_{0}$ and $U_{1}$ of $x_{0}$ and $x_{1}$, respectively.
@@
 Therefore, since $\mathcal{U}$ is an ultrafilter, there must be $i<2$ such that $\llbracket D_{i}\in\dot{F}\rrbracket_{\mathbb{B}}\in\mathcal{U}$, and hence $D_{i}\in\dot{F}_{\mathcal{U}}$.
-But then $x_{1-i}\notin D_{i}$ and $\bigcap\dot{F}_{\mathcal{U}}\subseteq D_{i}$, so $x_{1-i}\notin\bigcap\dot{F}_{\mathcal{U}}$, which is a contradiction.
+But then $x_{i}\notin D_{i}$ and $\bigcap\dot{F}_{\mathcal{U}}\subseteq D_{i}$, so $x_{i}\notin\bigcap\dot{F}_{\mathcal{U}}$, which is a contradiction.
 It follows that $\bigcap\dot{F}_{\mathcal{U}}$ contains precisely one point.
@@
 It is routine to verify that this is a reflexive duality: Given $g\in C(S,X)$, we have $f_{\dot{F}_{g}}=g$ and, given a $\mathbb{B}$-name $\dot{E}$ for a maximal filter on $\tau_{c}$, we have $\Vdash_{\mathbb{B}}\dot{E}=\dot{F}_{f_{\dot{E}}}$.
-Also, this story can clearly be relativized below a given condition $b\in\mathbb{B}$: functions in $C(N_{b},X)$ correspond to $\mathbb{B}$-names $\dot{F}$ such that $b\Vdash_{\mathbb{B}}``\dot{F}\text{ is a maximal filter on }\tau_{c}"$.
+Also, this story can clearly be relativized below a given condition $b\in\mathbb{B}^{+}$: functions in $C(N_{b},X)$ correspond to $\mathbb{B}$-names $\dot{F}$ such that $b\Vdash_{\mathbb{B}}``\dot{F}\text{ is a maximal filter on }\tau_{c}$".
 Now suppose that $X$ is an arbitrary Hausdorff space, not necessarily compact.
@@
 Given an open set $U\in\tau$ and $K\in\mathcal{K}(X)$, let $\hat{U}_{K}$ denote its interpretation in $\hat{\tau}_{K}$.
-We then define $\hat{U}_{\mathcal{K}}\subseteq\hat{X}_{\mathcal{K}}$ as follows: given $K\in\mathcal{K}(X)$ and $x\in\hat{K}$, put $[(x,K)]\in\hat{U}_{\mathcal{K}}$ if and only if $x\in\hat{U}_{K}$.
+We then define $\hat{U}_{\mathcal{K}}\subseteq\hat{X}_{\mathcal{K}}$ as follows: given $K\in\mathcal{K}(X)$ and $x\in\hat{K}$, put $(x,K)]\in\hat{U}_{\mathcal{K}}$ if and only if $x\in\hat{U}_{K}$.
 It is routine to check that this is well-defined.
@@
 Fix such an $x$, and let $A:=\{n<\omega\mid d(n)<x(n)\}$.
-By choice of $d$, $A$ is finite.
+By the choice of $d$, AA is finite.
 Then, in $V$, the set $U=\{y\in{{}^{\omega}}\omega\mid\forall n\in A~y(n)\neq d(n)\}$ is in $\tau$, and we have $O\cap\hat{K}_{x}=\hat{U}_{K_{x}}$.
@@
 $y(n)=\begin{cases}\sigma(n)&\text{if }n\in A\\ d(n^{*})&\text{if }n=n^{*}\\ 0&\text{otherwise.}\end{cases}$
-Then, in $W$, we have $y\in O\setminus\hat{U}_{\mathcal{K}}$, so $\hat{U}_{\mathcal{K}}\not\subseteq O$. ∎
+Then, in $W$, we have $y\in\hat{U}_{\mathcal{K}}\setminus O$, so $\hat{U}_{\mathcal{K}}\not\subseteq O$. ∎
 Since $\{\hat{U}_{\mathcal{K}}\mid U\in\tau\}$ is a base for $\hat{\tau}_{\mathcal{K}}$, this completes the proof. ∎
@@
 Let $C(S,X)^{V}/G$ denote the quotient space with respect to this equivalence relation.
-For each $U\in\tau$, define a set $N_{U}\subseteq C(S,X)^{V}/G$ as follows: for all $g\in C(S,X)^{V}$, put $[g]\in N_{U}$ if and only if there is $b\in G$ such that $g[N_{b}]\subseteq U$; note that this is well-defined.
+For each $U\in\tau$, define a set $N_{U}\subseteq C(S,X)^{V}/G$ as follows: for all $g\in C(S,X)^{V}$, put $g]\in N_{U}$ if and only if there is $b\in G$ such that $g[N_{b}]\subseteq U$; note that this is well-defined.
 Proposition 8.8.
@@
 It is immediate from the definition that each $N_{U}$ is open.
-Now let $\hat{O}\subseteq C(S,X)^{V}/G$ be open, and fix $g\in C(S,X)^{V}$ such that $[g]\in\hat{O}$.
+Now let $\hat{O}\subseteq C(S,X)^{V}/G$ be open, and fix $g\in C(S,X)^{V}$ such that $g]\in\hat{O}$.
 Let $O=\{h\in C(S,X)^{V}\mid[h]\in\hat{O}\}$.
 We can then find a compact $K\subseteq S$ and an open set $U\in\tau$ such that $g[K]\subseteq U$ and $\{h\in C(S,X)^{V}\mid h[K]\subseteq U\}\subseteq O$.
-Since $K$ is compact and $g$ is continuous, we can find a finite collection $b_{0},\ldots,b_{n-1}$ from $\mathbb{B}$ such that $K\subseteq\bigcup\{N_{b_{i}}\mid i<n\}$ and, for all $i<n$, we have $g[N_{b_{i}}]\subseteq U$.
+Since $K$ is compact and $g$ is continuous, we can find a finite collection $b_{0},\ldots,b_{n-1}$ from $\mathbb{B}^{+}$ such that $K\subseteq\bigcup\{N_{b_{i}}\mid i<n\}$ and, for all $i<n$, we have $g[N_{b_{i}}]\subseteq U$.
 Letting $b:=\bigvee\{b_{i}\mid i<n\}$, we get $K\subseteq N_{b}$ and $g[N_{b}]\subseteq U$.
@@
 In this case, we claim that $N_{U}\subseteq\hat{O}$.
-To see this, fix $h\in C(S,X)^{V}$ such that $[h]\in N_{U}$.
+To see this, fix $h\in C(S,X)^{V}$ such that $h]\in N_{U}$.
 We can therefore find $c\in G$ such that $h[N_{c}]\subseteq U$.
 Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{c}=h\restriction N_{c}$ and letting $h^{\prime}\restriction(S\setminus N_{c})$ be constant, taking an arbitrary value in $U$.
-Then $h^{\prime}[K]\subseteq U$, so $[h^{\prime}]\in\hat{O}$.
-Moreover, since $c\in G$, we have $h^{\prime}\sim_{G}h$, so $[h^{\prime}]=[h]$.
+Then $h^{\prime}[K]\subseteq U$, so $h^{\prime}]\in\hat{O}$.
+Moreover, since $c\in G$, we have $h^{\prime}\sim_{G}h$, so $h^{\prime}]=[h]$.
 Thus, $N_{U}\subseteq\hat{O}$.
@@
 Define $h^{\prime}\in C(S,X)^{V}$ by letting $h^{\prime}\restriction N_{b}=g\restriction N_{b}$ and $h^{\prime}\restriction(S\setminus N_{b})=h\restriction(S\setminus N_{b})$.
-Then $h^{\prime}[N_{b}]=g[N_{b}]\subseteq U$, so $[h^{\prime}]\in\hat{O}$.
-Since $b\notin G$, we have $[h^{\prime}]=[h]$, so $[h]\in\hat{O}$.
+Then $h^{\prime}[N_{b}]=g[N_{b}]\subseteq U$, so $h^{\prime}]\in\hat{O}$.
+Since $b\notin G$, we have $h^{\prime}]=[h]$, so $h]\in\hat{O}$.
 ∎
@@
 To show that $k$ is injective, fix two functions $g,h\in C(S,X)^{V}$. Familiar arguments yield the fact that, in $V$, the set
-$\{b\in\mathbb{B}\mid g\restriction N_{b}=h\restriction N_{b}\text{ {or} }g[N_{b}]\cap h[N_{b}]=\emptyset\}$
-is dense in $\mathbb{B}$.
+$\{b\in\mathbb{B}^{+}\mid g\restriction N_{b}=h\restriction N_{b}\text{ {or} }g[N_{b}]\cap h[N_{b}]=\emptyset\}$
+is dense in $\mathbb{B}^{+}$.
 The genericity of $G$ then yields the injectivity of $k$.
-To show that $k$ is surjective, fix an element $[(F,K)]\in\hat{X}_{\mathcal{K}}$, and let $\dot{F}$ be a $\mathbb{B}$-name for $F$ that is forced by $1_{\mathbb{B}}$ to be a maximal filter on $(\tau_{K})_{c}$.
+To show that $k$ is surjective, fix an element $(F,K)]\in\hat{X}_{\mathcal{K}}$, and let $\dot{F}$ be a $\mathbb{B}$-name for $F$ that is forced by $1_{\mathbb{B}}$ to be a maximal filter on $(\tau_{K})_{c}$.
 Then $f_{\dot{F}}$, as defined at the beginning of this section, is in $C(S,X)^{V}$, and it is routine to verify that $k([f_{\dot{F}}])=[(F,K)]$.
@@
 Proof.
-Suppose that $g\in C(S,X)^{V}$ and $[g]\in N_{U}$.
+Suppose that $g\in C(S,X)^{V}$ and $g]\in N_{U}$.
 Let $K:=g[S]$, and fix $b\in G$ such that $g[N_{b}]\subseteq U$.
 Then, by construction, we have $(K\setminus U)\notin F_{g}$, so $k([g])=[(F_{g},K)]\in\hat{U}_{\mathcal{K}}$.
-Conversely, suppose that $g\in C(S,X)^{V}$ and $[g]\notin N_{U}$.
+Conversely, suppose that $g\in C(S,X)^{V}$ and $g]\notin N_{U}$.
 Again, let $K:=g[S]$.
-By familiar arguments, the following set is dense in $\mathbb{B}$ (in $V$):
-$\{b\in\mathbb{B}\mid g[N_{b}]\subseteq U\text{ {or} }g[N_{b}]\subseteq K\setminus U\}.$
-Since $[g]\notin N_{U}$, the genericity of $G$ implies that $K\setminus U\in F_{g}$, so $k([g])=[(F_{g},K))]\notin\hat{U}_{\mathcal{K}}$. ∎
+By familiar arguments, the following set is dense in $\mathbb{B}^{+}$ (in $V$):
+$\{b\in\mathbb{B}^{+}\mid g[N_{b}]\subseteq U\text{ {or} }g[N_{b}]\subseteq K\setminus U\}.$
+Since $g]\notin N_{U}$, the genericity of $G$ implies that $K\setminus U\in F_{g}$, so $k([g])=[(F_{g},K))]\notin\hat{U}_{\mathcal{K}}$. ∎
 Altogether, we have established the following theorem.
@@
 Therefore, the results of this section show that one way to think about the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, as noted, for all such spaces $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
-Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [26, pp.
+Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [28, pp.
 8-9 and Appendix to Lecture II]).
@@
 Third, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products.
-Namely, if $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$.
+Namely, if $I$ is an index set and $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$.
 This is because the collection
@@
 is $\subseteq$-cofinal in $\mathcal{K}(\prod_{i\in I}X_{i})$.
-Fourth, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact.
-In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11.
-If $X$ is compact, then, for every $\mathbb{B}$-name $\dot{x}$ for an element of $\hat{X}_{\mathcal{K}}(=\hat{X})$, there is a $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$.
-In particular, in this case we can take $\dot{x}$ to be a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$, then then let $g=f_{\dot{x}}$, as defined at the outset of this section.
-However, in general this may not be the case, and there may be $\mathbb{B}$-names $\dot{x}$ for elements of $\hat{X}_{\mathcal{K}}$ for which there is no $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$.
-For such $\dot{x}$, one may need to work below some condition $b\in\mathbb{B}$ that forces $\dot{x}$ to lie inside the interpretation of some particular $K\in\mathcal{K}(X)$; once one does so, one may then find $g\in C(S,X)$ such that $b\Vdash\dot{k}([g])=\dot{x}$.
-This issue will lead to some slight complications and asymmetries in the next section.
+[…]
 We end this section by noting that these observations are of wider significance than might initially be suspected, by reason of the variety of expressions within condensed mathematics in which sets of the form $C(S,X)$, sometimes covertly, partake.
@@
 In this section, we establish an equivalence between the settings of $C(S,X)$ on the one hand and $\mathbb{B}$-names for elements of $\hat{X}_{\mathcal{K}}$ on the other, with respect to interpretations of certain sentences.
-Fix for now a positive integer $n$ and Hausdorff spaces $\langle X_{i}\mid i<n\rangle=\vec{X}$.
+Fix a positive integer $n$ and Hausdorff spaces $\langle X_{i}\mid i<n\rangle=\vec{X}$.
 We will specify a collection $\mathcal{L}(\vec{X})$ of formulas.
@@
 Given $i<n$ and $f\in C(S,X_{i})$, let $\dot{x}_{f}$ be a $\mathbb{B}$-name for $\dot{k}([f])$, where $\dot{k}$ is a name for the homeomorphism constructed to witness Theorem 8.11.
-Concretely, if $K=f[S]$, then $\dot{x}_{f}$ is a $\mathbb{B}$-name for $[(\dot{F}_{f},K)]\in\hat{X}_{\mathcal{K}}$.
+Concretely, if $K=f[S]$, then $\dot{x}_{f}$ is a $\mathbb{B}$-name for $(\dot{F}_{f},K)]\in\hat{X}_{\mathcal{K}}$.
 Theorem 9.1.
@@
 For notational convenience, we can assume here that $n=1$ by considering the single space $X=\prod_{i<n}X_{i}$ (recalling that $\widehat{(\prod_{i<n}X_{i})}_{\mathcal{K}}=\prod_{i<n}(\hat{X}_{i})_{\mathcal{K}}$).
-Then there is an open set $U\subseteq X$ such that $\varphi$ is the formula $v_{1}\in U$, and we are given a function $f\in C(S,X)$.
-Let $K:=\mathrm{im}(f)$, so $K$ is a compact subset of $X$, and $\dot{x}_{f}$ is forced to be of the form $[(\dot{F}_{f},K)]$.
+Then there is an open set $U\subseteq X$ such that $\varphi$ is the formula $v_{0}\in U$, and we are given a function $f\in C(S,X)$.
+Let $K:=\mathrm{im}(f)$, so $K$ is a compact subset of $X$, and $\dot{x}_{f}$ is forced to be of the form $(\dot{F}_{f},K)]$.
 To verify (1), suppose that $C(S,X)\models_{b}\varphi(f)$, and hence $\{s\in N_{b}\mid f(s)\in U\}$ is dense in $N_{b}$.
-Suppose for sake of contradiction that $b\not\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$.
+Suppose for the sake of contradiction that $b\not\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$.
 Then, by the definition of $\dot{F}_{f}$, there is $c\leq b$ such that $(\check{K}\setminus\check{U},c)\in\dot{F}_{f}$, and therefore, again by the definition of $\dot{F}_{f}$, we have $N_{c}\subseteq f^{-1}[K\setminus U]$, which is a contradiction.
 For (2), suppose that $b\Vdash_{\mathbb{B}}\dot{x}_{f}\in\hat{U}$, i.e., $b\Vdash_{\mathbb{B}}(K\setminus U)\notin\dot{F}_{f}$.
-Suppose for sake of contradiction that $\{s\in N_{b}\mid f(s)\in U\}$ is not dense in $N_{b}$.
+Suppose for the sake of contradiction that $\{s\in N_{b}\mid f(s)\in U\}$ is not dense in $N_{b}$.
 Then there is $c\leq b$ such that $N_{c}\subseteq f^{-1}[K\setminus U]$.
@@
 Now suppose that $j=j^{\prime}+1$ for some $j^{\prime}<n$.
-To verify (1), suppose that $C(S,X)\models_{b}\varphi(f_{j},\ldots f_{n-1})$ but, for sake of contradiction, assume that $V^{\mathbb{B}}\not\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$.
+To verify (1), suppose that $C(S,X)\models_{b}\varphi(f_{j},\ldots f_{n-1})$ but, for the sake of contradiction, assume that $V^{\mathbb{B}}\not\models_{b}\varphi(\dot{x}_{f_{j}},\ldots,\dot{x}_{f_{n-1}})$.
 Suppose first that $\varphi$ is of the form $\forall x\psi$ for some $\psi\in\mathcal{L}_{j^{\prime}}(\vec{X})$.
@@
 Theorem 9.2.
-Suppose that $A$ is an abelian group and $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$, and let $B_{K}$ be a basis for $K$, and let $e:B_{K}\rightarrow\omega$ be a function.
+Suppose that $A$ is an abelian group and $0\rightarrow K\xrightarrow{\subseteq}F\rightarrow A\rightarrow 0$ is a free resolution of $A$.
+Let $B_{K}$ be a basis for $K$, and let $e:B_{K}\rightarrow\omega$ be a function.
 Suppose that $\mathbb{B}$ is a complete Boolean algebra that forces the existence of a homomorphism $\tau:K\rightarrow\mathbb{Z}$ such that
@@
 Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$.
-Since $V=L$, $A$ is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$.
+Since $V=L$, AA is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$.
 Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$.
 However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.
-However, we can recover an informative partial converse to Theorem 9.2, in the following sense.
+We can, though, recover an informative partial converse to Theorem 9.2, in the following sense.
 Note that, in the situation described in the paragraph above, forcing with $\mathbb{B}$ necessarily adds a new element $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$ that extends $\tau$.
@@
 Let $f:S(\mathbb{B})\rightarrow X_{K}$ be the constant function taking value $\tau\restriction B_{K}$.
-Then, by choice of $\tau$, we have $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi(f)$.
+Then, by the choice of $\tau$, we have $C(S(\mathbb{B}),X_{F}\times X_{K})\models\varphi(f)$.
 By Theorem 9.1, it follows that $V^{\mathbb{B}}\models\varphi(\dot{x}_{f})$.
@@
 10. Martin’s Axiom and $\mathsf{CondAb}_{\omega_{1}}$
-In this section, we show that the independence of Whitehead’s problem in the classical setting can persist if we restrict ourselves to $\mathsf{CondAb}_{\kappa}$ for certain cardinal $\kappa$.
+In this section, we show that the independence of Whitehead’s problem in the classical setting can persist if we restrict ourselves to $\mathsf{CondAb}_{\kappa}$ for certain cardinals $\kappa$.
 Both for concreteness and because the category $\mathsf{CondAb}_{\omega_{1}}$ of light condensed sets provides the setting for [8], we focus on the case $\kappa=\omega_{1}$.
@@
 In fact, we shall prove that this follows from $\sf MA_{\omega_{1}}$, the same hypothesis used by Shelah in his proof of Theorem 1.1.
-In what follows, when we write, e.g., that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, it should be implicitly understood that $\Lambda$ is a directed partial order and each $S_{i}$ is a finite (discrete) space.
-For each $i\in\Lambda$, we get a projection map $\pi_{i}:S\rightarrow S_{i}$.
-The basic open subsets of $S$ are of the form $\pi_{i}^{-1}\{\bar{s}\}$, where $i\in\Lambda$ and $\bar{s}\in S_{i}$.
-If we write $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, then this means, moreover, that $\Lambda$ is countable.
+[…]
 Proposition 10.1.
-Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$ and $\{O_{k}\mid k<n\}$ is a finite partition of $S$ into nonempty clopen sets.
-Then there is an $i\in\Lambda$ and a function $f:S_{i}\rightarrow n$ such that, for all $k<n$, we have $O_{k}=\{s\in S\mid f(\pi_{i}(s))=k\}$.
+[…]
 Proof.
@@
 Proposition 10.2.
-Suppose that $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}$, $A$ and $B$ are (discrete) abelian groups, and $\varphi:S\rightarrow\mathrm{Hom}(A,B)$ is a continuous map.
-For every finitely generated subgroup $A_{0}$ of $A$, there is an $i\in\Lambda$ and a map $\varphi_{0}:S_{i}\rightarrow\mathrm{Hom}(A_{0},B)$ such that, for all $s\in S$, we have
+[…]
 $\varphi(s)\restriction A_{0}=\varphi_{0}(\pi_{i}(s)).$
@@
 ∎
-Recall the definition of Chase’s condition from the end of Section 9.
-We now show that Theorem 9.4 transfers to the richer setting of $\mathsf{CondAb}_{\omega_{1}}$.
-By clause (1) of Theorem 1.3, this is sharp, since for any nonfree abelian group of size $\aleph_{1}$, the profinite set $2^{\omega_{1}}$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{2}}}(\underline{A},\underline{\mathbb{Z}})\neq 0$.
+Recall the definition of Chase’s condition from the end of Section 9. We now show that Theorem 9.4 transfers to the richer setting of $\mathsf{CondAb}_{\omega_{1}}$.
 Theorem 10.3.
-Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of cardinality $\omega_{1}$ satisfying Chase’s condition. Then
+Suppose that $\sf MA_{\omega_{1}}$ holds and $A$ is an abelian group of cardinality $\aleph_{1}$ satisfying Chase’s condition. Then
 $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}_{\omega_{1}}}(\underline{A},\underline{\mathbb{Z}})=0.$
 Proof.
-Let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be the canonical free resolution of $A$, let $S=\lim_{i\in\Lambda}S_{i}\in\mathsf{Prof}_{\omega_{1}}$, and let $\varphi:S\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ be continuous.
-We must produce a continuous map $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ such that $\psi(s)\restriction K=\varphi(s)$ for all $s\in S$.
-We fix a basis $\{\bar{a}\mid a\in A\}$ for $F$; the surjection from $F$ to $A$ in the free resolution above is then generated by the map that sends $\bar{a}$ to $a$ for all $a\in A$.
+[…]
+Fix a basis $\{\bar{a}\mid a\in A\}$ for $F$; the surjection from $F$ to $A$ in the free resolution above is then generated by the map that sends $\bar{a}$ to $a$ for all $a\in A$.
 For each subgroup $A^{\prime}\subseteq A$, let $F_{A^{\prime}}$ be the free group generated by $\{\bar{a}\mid a\in A^{\prime}\}$, and let $K_{A^{\prime}}$ be the kernel of the canonical surjection from $F_{A^{\prime}}$ to $A^{\prime}$.
@@
 Now fix $a\in A^{\prime}\setminus B^{\prime}$.
-Then there is a unique way to express $a$ in the form $\sum_{b\in B^{*}}c_{b}b$, where $B^{*}\subseteq B^{\prime}$ is finite and, for all $b\in B^{*}$, $c_{b}$ is a nonzero integer.
+There is a unique way to express $a$ in the form $\sum_{b\in B^{*}}c_{b}b$, where $B^{*}\subseteq B^{\prime}$ is finite and, for all $b\in B^{*}$, $c_{b}$ is a nonzero integer.
 To satisfy requirement (2), we are obliged to let
@@
   • for all $s\in S$, $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$.
-If $p,q\in\mathbb{P}$, then $q\leq p$ if and only if $A_{q}\supseteq A_{p}$ and, for all $s\in S$, $\psi_{q}(s)\restriction F_{A_{p}}=\psi_{q}(s)$.
+If $p,q\in\mathbb{P}$, then $q\leq p$ if and only if $A_{q}\supseteq A_{p}$ and, for all $s\in S$, $\psi_{q}(s)\restriction F_{A_{p}}=\psi_{p}(s)$.
 Claim 10.5.
@@
 Define a function $f:S\rightarrow{{}^{B_{0}\cup B_{1}}}\mathbb{Z}$ as follows: for all $s\in S$ and $b\in b_{0}\cup B_{1}$, let
-$f(s)(b):=\begin{cases}\psi_{p}(s)(b)&\text{if }b\in B_{0}\\ 0&\text{if }b\in B_{1}.\end{cases}$
+$f(s)(b):=\begin{cases}\psi_{p}(s)(\bar{b})&\text{if }b\in B_{0}\\ 0&\text{if }b\in B_{1}.\end{cases}$
 Since $\psi_{p}$ is continuous, it follows that $f$ is continuous.
@@
 Define $\psi:S\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ by letting $\psi(s)=\bigcup\{\psi_{p}(s)\mid p\in G\}$.
-The fact that each $\psi_{p}$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$ implies that $\psi$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K=\varphi(s)$, as desired.
+The fact that each $\psi_{p}$ is a continuous map such that, for all $s\in S$, we have $\psi_{p}(s)\restriction K_{A_{p}}=\varphi(s)\restriction K_{A_{p}}$ implies that $\psi$ is a continuous map such that, for all $s\in S$, we have $\psi(s)\restriction K=\varphi(s)$, as desired.
 ∎
-We thus obtain the following corollary, establishing Theorem 1.4(1).
+Recalling Griffith’s result [13] stating that there is provably in $\sf ZFC$ a nonfree abelian group of size $\aleph_{1}$ satisfying Chase’s condition, we thus obtain the following corollary, establishing Theorem 1.4(1).
 Corollary 10.7.
@@
 $=\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))$
-To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [18, Prop.
+To see the fifth equality, apply Lemma 3.2; for the third and fourth, respectively, apply [19, Prop.
 3.5 and Prop.
 3.2].
-The first equality follows from [26, p.
+The first equality follows from [28, p.
 13 (iii) (cf.
@@
 This leaves us only with the second.
-Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [26, p.
+Here the idea, evidently, is to apply the Hom-tensor adjunction at the level of $\mathrm{Ext}^{1}_{\mathsf{CondAb}}$, and we may do so by the naturality of the adjunction at the $\mathrm{Hom}_{\mathsf{CondAb}}$-level, the flatness of $\mathbb{Z}[\underline{S}]$ [28, p.
 13], and the fact, noted already, that $\underline{\bigoplus_{I}\mathbb{Z}}=\bigoplus_{I}\underline{\mathbb{Z}}$ for any set $I$.
@@
 Let us check the details.
-Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [18, Prop.
+Observe first that $\underline{\mathcal{P}}$ is a projective resolution of $\underline{A}$, by [19, Prop.
 2.18] and Proposition 12.3 below; this shows the third equality.
@@
 $\big[\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for }S=*\,\big]\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
-is independent of the $\mathsf{ZFC}$ axioms. Clausen and Scholze’s result, in contrast, can now be seen to follow from the fact that the implication
-$\big{[}\,\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,C(S,\mathbb{Z}))=0\text{ for all }S\in\mathsf{ED}\,\big{]}\;\Rightarrow\;\text{the abelian group }A\text{ is free}$
-is indeed a $\mathsf{ZFC}$ theorem.
+is independent of the $\sf ZFC$ axioms. Clausen and Scholze’s result, in contrast, can now be seen to follow from the fact that the implication
+[…]
+is indeed a $\sf ZFC$ theorem.
 This implication is, moreover, straightforward to prove, since by a result of Nöbeling ([23], see also [10, Proposition XI.4.4]), generalizing work of Specker [31], the group $C(S,\mathbb{Z})$ is itself a free abelian group, one isomorphic to the free abelian group on the topological weight of $S$.
@@
 It follows that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\kappa)})\neq 0$.
-Therefore, Lemma 11.1 immediately yields the following corollary, providing an alternative proof of clauses (1) and (3) of Theorem 1.3.
+Therefore, Lemma 11.1 immediately yields the following corollary, providing an alternative proof of clause (2) of Theorem 1.3.
 Corollary 11.2.
@@
 Lemma 11.1 also sheds additional light on the results of Section 10, entailing both an alternative proof of a variation of that section’s main result and limits, in the presence of sufficiently large cardinals, on the extent to which this result can be generalized.
-We first note that, by results in [11], it is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$.
+We first note that, by results in [11], it is consistent with the axioms of $\sf ZFC$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,B)=0$ for every abelian group $B$ of cardinality ${\leq}\kappa$.
 Thus, Lemma 11.1 yields the following corollary, establishing clause (2) of Theorem 1.4 and indicating that, consistently, the positive answer to the interpretation of Whitehead’s problem in $\mathsf{CondAb}$ established by Theorem 1.2 fails in $\mathsf{CondAb}_{\kappa}$ for every uncountable cardinal $\kappa$.
 Corollary 11.3.
-It is consistent with the axioms of $\mathsf{ZFC}$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for all $S\in\mathsf{ED}_{\kappa}$.
+It is consistent with the axioms of $\sf ZFC$ that, for every cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$ for all $S\in\mathsf{ED}_{\kappa}$.
 In particular, it is consistent that, for every uncountable cardinal $\kappa$, there is a nonfree abelian group $A$ such that $\underline{A}$ is Whitehead in $\mathsf{CondAb}_{\kappa}$.
@@
 In particular, for for every $S\in\mathsf{ED}_{\kappa}$, we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S)=0$.
-By Lemma 11.1, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$.
-By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality ${<}\kappa$ is free.
+By Lemma 11.1 and the aforementioned result of Nöbeling indicating that, for every $S\in\mathsf{ED}$, $C(S,\mathbb{Z})$ is the free abelian group on the topological weight of $S$, it follows that, for every cardinal $\lambda<\kappa$, we have $\mathrm{Ext}^{1}_{\mathsf{Ab}}(A,\mathbb{Z}^{(\lambda)})=0$.
+By the observations in the paragraph immediately preceding Corollary 11.2, it follows that $A$ is $\kappa$-free, i.e., every subgroup of $A$ of cardinality $\kappa$ is free.
 By [10, Corollary II.3.11], since $\kappa$ is strongly compact, $A$ is free, as desired.
@@
 This is true if $\kappa$ is a strongly compact cardinal, but we remark that, consistently, this can also be true for smaller values of $\kappa$.
-In particular, it is shown in [16, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
+In particular, it is shown in [17, §4] that, assuming the consistency of the existence of infinitely many supercompact cardinals, there exists a model of $\sf ZFC$ in which this condition holds for the first cardinal fixed point of the $\aleph$ function, i.e., the least $\kappa$ such that $\kappa=\aleph_{\kappa}$.
 12. Locally compact and solid abelian groups
-In this section, we show that Theorem 1.2 extends from the category of (discrete) abelian groups to the category of locally compact abelian groups; here it is useful to recall that the projective objects of the former category are exactly the free ones.
-In particular, we prove the following theorem.
+In this section, we prove Theorem 1.5, which we restate here for the reader’s convenience.
 Theorem 12.1.
@@
 Proof.
-By the structure theorem for locally compact abelian groups (see [26, Thm.
+By the structure theorem for locally compact abelian groups (see [28, Thm.
 4.1(i)]), any locally compact abelian group $A$ is of the form $\mathbb{R}^{n}\times B$ for some nonnegative integer $n$, where $B$ is an extension of a compact group by a discrete one, or where, in other words, there exists a short exact sequence
 (1) $0\to C\to B\to D\to 0$
-with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [18, Props.
+with $C$ compact and $D$ discrete whose image in $\mathsf{CondAb}$ is exact as well (see [19, Props.
 4.9, 2.18]).
-Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [26, p.
+Since $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$ [28, p.
 25], we have
@@
 $\cdots\to 0\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{D},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})\to 0\to\cdots$
-(The rightmost 00 follows from [26, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 00 follows from [26, Prop. 4.2].) Hence
+(The rightmost 0 follows from [28, p. 27] (e.g., from the comment following Cor. 4.9); the leftmost 0 follows from [28, Prop. 4.2].) Hence
 $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{B},\underline{\mathbb{Z}})=0$
@@
 By Theorem 1.2, item (1) holds if and only if $D=\bigoplus_{I}\mathbb{Z}$ for some $I$.
-Also, the argument of [18, Prop.
-4.11] together with [26, Prop.
-4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([26, Prop.
-4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline{\mathbb{R}/\mathbb{Z}}$ is an isomorphism.
-Again applying [26, Prop.
+Also, the argument of [19, Prop.
+4.11] together with [28, Prop.
+4.3(ii)] shows that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{R}})=0$; this together with the fact that $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}})=\underline{\mathrm{Hom}(C,\mathbb{R})}=0$ ([28, Prop.
+4.2]) implies that the map $\underline{\mathrm{Hom}}(\underline{C},\underline{\mathbb{R}/\mathbb{Z}})\to\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})$ appearing in the long exact sequence induced by applying $\underline{\mathrm{Hom}}(\underline{C},-)$ to $\underline{\mathbb{Z}}\to\underline{\mathbb{R}}\to\underline$$\mathbb{R}$$\mathbb{Z}$} is an isomorphism.
+Again applying [28, Prop.
 4.2], we have $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{C},\underline{\mathbb{Z}})=\underline{\mathrm{Hom}(C,\mathbb{R}/\mathbb{Z})}=\underline{\mathbb{D}(C)}$, the condensed image of the Pontryagin dual group of $C$; in particular, item (2) holds if and only if $C=0$.
 Summing up, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})=0$ if and only if $A$ is of the form $\mathbb{R}^{n}\times D$ for some $D=\bigoplus_{I}\mathbb{Z}$.
-As noted, by [21, Prop.
+As noted, by [22, Prop.
 3.3], such $A$ are precisely the projective objects of $\mathsf{LCA}$.
@@
 The underlying group of $\mathbb{Z}[\underline{S}]$ is the free abelian group $\bigoplus_{S}\mathbb{Z}$.
-By the lemma, any section $s:\underline{\mathbb{R}}\to\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]$ of an $r$ as above would evaluate over a point $\ast$ as an injection $s(*):\mathbb{R}\to\bigoplus_{I}\bigoplus_{S_{i}}\mathbb{Z}$, but any nonzero element in the image of $s(*)$ would need to be divisible by every positive $n$, a contradiction.
+By the lemma, any section $s:\underline{\mathbb{R}}\to\bigoplus_{I}\mathbb{Z}[\underline{S_{i}}]$ of an $r$ as above would evaluate over a point $\ast$ as an injection $s(*):\mathbb{R}\to\bigoplus_{I}\bigoplus_{S_{i}}\mathbb{Z}$, but any nonzero element in the image of s()s(*) would need to be divisible by every positive $n$, a contradiction.
 The following, therefore, will complete the proof of Corollary 12.2.
@@
 (2) $T\mapsto\bigoplus_{\mathrm{Cont}(T,S)}\mathbb{Z}$
-Recall also that the sheafification $\mathcal{F}^{\sharp}$ of a presheaf $\mathcal{F}$ on a site is given by the twofold iteration of the $+$ operation, where
+Recall also that the sheafification $\mathcal{F}^{\sharp}$ of a presheaf $\mathcal{F}$ on a site is given by the twofold iteration of the ++ operation, where
 $\mathcal{F}^{+}(U)=\mathrm{colim}_{\mathcal{U}\in\mathrm{cov}(U)}\,\,\mathrm{H}^{0}(\mathcal{U};\mathcal{F}).$
@@
 A natural next question, then, is how much further this $d$ may be pushed.
-By Corollary 12.2, of course, Whitehead’s problem has a ZFC solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$.
+By Corollary 12.2, of course, Whitehead’s problem has a $\sf ZFC$ solution when $d=\mathsf{CondAb}$ as well, but it is an unsatisfying one, deriving as it does from the arguably spurious case of $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{\mathbb{R}},\underline{\mathbb{Z}})=0$.
 Against this background, particularly interesting is the problem when $d$ equals the subcategory $\mathsf{Solid}$ of $\mathsf{CondAb}$ mentioned briefly in our introduction.
-Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [26, §5–7] and [25, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2.
-The solid setting, being “very much a ‘nonarchimedean’ notion” [26, footnote 10 on p.
+Forming, as it does, the paradigmatic example of an analytic ring, this is one of $\mathsf{CondAb}$’s most critical subcategories (see [28, §5–7] and [27, §2]); moreover, as noted, it was this category’s features which Clausen and Scholze leveraged in their original proof of Theorem 1.2.
+The solid setting, being “very much a ‘nonarchimedean’ notion” [28, footnote 10 on p.
 32], is also one in which $\underline{\mathbb{R}}$ does not appear.
-[^6] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 00 [26, Corollary 6.1(iii)].
-Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [26, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
+[^7] More precisely, the inclusion $\mathsf{Solid}\to\mathsf{CondAb}$ possesses a left adjoint solidification functor notated $M\mapsto M^{\blacksquare}$, and the solidification $\underline{\mathbb{R}}^{\blacksquare}$ of $\underline{\mathbb{R}}$ is 0 [28, Corollary 6.1(iii)].
+Within it, lastly, $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}$ and $\underline{\mathrm{Ext}}^{1}_{\mathsf{Solid}}$ coincide [28, Lemma 5.9], so that this instance of the problem recovers the $s=d$ symmetry of the classical one:
 Question 12.5.
@@
 where $\mathbf{A}[H]$ denotes for any abelian group $H$ the inverse system indexed by functions $f:\mathbb{N}\to\mathbb{N}$, with terms $A_{f}:=\bigoplus_{i\in\mathbb{N}}H^{(f(i))}$ and natural projection maps $A_{g}\to A_{f}$ for any $f$ which is pointwise dominated by $g$.
-See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 33]), from which the consistency with the ZFC axioms of the statement that
+See [5] for details of this conversion; the immediate point, though, is that these derived limits have been the subject of much recent study (see [4, 3, 35]), from which the consistency with the $\sf ZFC$ axioms of the statement that
 $\mathrm{lim}^{1}\,\mathbf{A}[H]=0\;\text{ but }\mathrm{lim}^{2}\,\mathbf{A}[H]\neq 0$
@@
 $\underline{\mathrm{Ext}}^{n}_{\mathsf{Solid}}(A,\underline{\mathbb{Z}})=0\text{ for all }n>0$
-is also consistent with the ZFC axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
+is also consistent with the $\sf ZFC$ axioms [1, Thm. 3]. By the following argument, though, $A$ witnesses a negative answer to Question 12.5 in this scenario as well.
 Consider the surjection
@@
 defined as follows: fix $g:\mathbb{N}\to\mathbb{N}$ and an element $a$ of the $g^{\mathrm{th}}$ summand of the domain of $t$; clearly it will suffice to describe the effect of $t$ on such $a$.
-Here it is convenient to adopt the (elsewhere innocuous) convention that 00 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on.
+Here it is convenient to adopt the (elsewhere innocuous) convention that 0 is not a natural number: the first $g(1)$ many entries in $a$ are then mapped to the first $g(1)$ summands of the first factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, the next $g(2)$ many entries to the first $g(2)$ many summands of the second factor of $\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}\mathbb{Z}$, and so on.
 Equipping $\mathbb{Z}$ with the discrete topology and letting products and sums induce the product and coproduct topologies, respectively, we may view $t$ as a continuous mapping of topological abelian groups.
 Consider the condensed image $\underline{t}$ of $t$: observe that its codomain is $A$, its domain is solid, and that it remains an epimorphism (cf.
-[18, Prop.
+[19, Prop.
 2.18]).
@@
 and hence that $A$ is not projective, completing the argument.
-Even this doesn’t altogether settle the question, though, since it is also consistent with the ZFC axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [17, Thm.
+Even this doesn’t altogether settle the question, though, since it is also consistent with the $\sf ZFC$ axioms (and follows from the continuum hypothesis) that $\underline{\mathrm{Ext}}^{1}(A,\underline{\mathbb{Z}})\neq 0$ [18, Thm.
 2].
@@
 $=(\underline{\mathbb{Z}[1/p]/\mathbb{Z}})[-1]$
-(see [19, Rmk. 2, p. 2] for the first equality), and for its denominator
+(see [20, Rmk. 2, p. 2] for the first equality), and for its denominator
 […]
@@
 $=(\underline{\mathbb{Z}_{q}/\mathbb{Z}})[-1].$
-Hence, applying $R\underline{\mathrm{Hom}}_{\mathsf{CondAb}}(\,\cdot\,,\underline{\mathbb{Z}})$ to the short exact sequence
+[…]
 $0\to\underline{\mathbb{Z}[1/q]}\to\underline{\mathbb{Z}_{p}}\to\underline{\mathbb{Z}_{p}}/\underline{\mathbb{Z}[1/q]}\to 0$
@@
 We feel that a resolution of this conjecture will shed light not just on the relationship between forcing and condensed mathematics, but also on considerations of Whitehead’s problem in the classical setting.
+Acknowledgements
+A portion of this work was carried out while all three authors were participating in the Thematic Program on Set Theoretic Methods in Algebra, Dynamics and Geometry at the Fields Institute in spring of 2023.
+We thank the Fields Institute for their support and hospitality.
+The results of this paper were presented in detail at the Cornell Logic Seminar in Fall 2023.
+We would like to thank the participants of the seminar, particularly Mark Schachner and Justin Moore, for their engagement with the material and for a number of helpful remarks, and to thank Peter Scholze for valuable remarks on the paper as well.
+We thank Radek Honzík for pointing out a problem with an earlier formulation of Theorem 6.3 and Nick Rozenblyum for pointing out a gap in an earlier version of Theorem 1.3.
+Finally, we thank the anonymous referee for a number of helpful suggestions and corrections.